### v1

# SDP bounds on quantum codes

Abstract.
This paper provides a semidefinite programming hierarchy based on state polynomial optimization to determine the existence of quantum codes with given parameters. The hierarchy is complete, in the sense that if a $(\!(n,K,\delta)\!)_{2}$ code does not exist then a level of the hierarchy is infeasible. It is not limited to stabilizer codes and thus applicable generally. While it is formally dimension-free, we restrict it to qubit codes through quasi-Clifford algebras. We derive the quantum analog of a range of classical results: first, from an intermediate level a Lovász bound for self-dual quantum codes is recovered. Second, a symmetrization of a minor variation of this Lovász bound recovers the quantum Delsarte bound. Third, a symmetry reduction using the Terwilliger algebra leads to semidefinite programming bounds of size $O(n^{4})$. With this we give an alternative proof that there is no $(\!(7,1,4)\!)_{2}$ quantum code, and show that $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes do not exist.
GM and FH were supported by the Foundation for Polish Science through TEAM-NET (POIR.04.04.00-00-17C1/18-00). AN was supported by QLCI. We thank Christine Bachoc, Simeon Ball, Bram Bekker, Marcus Grassl, Carlos de Gois, Otfried Gühne, Kiara Hansenne, Moisés Bermejo Morán, Claudio Procesi, René Schwonnek, Andreas Winter, and Zhen-Peng Xu for fruitful discussions and helpful comments.
1. Introduction
Coding theory studies how to send messages error-free through a noisy channel. Typically, a classical message consists of a bit string of length $k$, with the noisy classical channel inducing errors that may flip some bits in the message. A error-correcting code allows one to encode the message with redundant information, so that it can be recovered after being sent through the channel. A paradigmatic example is the repetition code of block length $n$ where the message bit is repeated $n$ times, e.g., the bits 00 and $1$ are encoded as $000$ and $111$ respectively when $n=3$.
A quantum message is represented by a positive semidefinite matrix of trace one acting on the $K$-dimensional Hilbert space $\mathds{C}^{K}$. That is, $\varrho\in L(\mathds{C}^{K})$ with $\varrho\succeq 0$ and $\operatorname{tr}(\varrho)=1$. A noisy quantum channel, represented in its Kraus decomposition by $\mathcal{N}(\varrho)=\sum_{E\in\mathcal{N}}E\varrho E^{\dagger}$ with $\sum_{E\in\mathcal{N}}E^{\dagger}E\preceq\mathds{1}$, may not only induce flip errors but also errors in the phases of the state. The repetition code is no longer valid for quantum codes due to the no-cloning theorem [73], and therefore more involved strategies have to be used for the error-correction of quantum messages. A quantum error-correcting code encodes a $K$-dimensional Hilbert space to a subspace of the $n$-qubit Hilbert space, such that the action of a noisy quantum channel can be corrected. The Knill-Laflamme conditions provide necessary and sufficient conditions for this to be possible, and can be formulated in terms of the projector $\Pi$ onto the code subspace: if $\Pi E^{\dagger}F\Pi=c_{EF}\Pi$ for all errors $E,F$ appearing in the channel $\mathcal{N}$, then there exists a recovery map $\mathcal{R}$ such that $\mathcal{R}(\mathcal{N}(\varrho))\propto\varrho$ [35]. The size $K$ of the quantum code then equals the rank of $\Pi$.
One of the fundamental problems in coding theory is to identify the maximum code size $K$ for a given block length $n$ and minimum distance $\delta$. This problem is usually approached by defining weight enumerators [13, 67]: for classical linear codes $C$, these count the number of codewords of given Hamming weight, $A_{j}(C)=|\{c\in C\,:\,\operatorname{wt}(c)=j\}|$, where $\operatorname{wt}(c)$ is the number of non-zero positions in the bit string $c$. In turn, the quantum weight enumerators $A_{j}(\Pi)$ decompose the Hilbert-Schmidt norm $\operatorname{tr}(\Pi^{\dagger}\Pi)$ of the projector $\Pi$ into contributions according to the support of different quantum correlations [28]. One has $A_{j}=\sum_{E\in\mathcal{E}_{n}\,:\,\operatorname{wt}(E)=j}=\operatorname{tr}(\Pi E^{\dagger})\operatorname{tr}(\Pi E)$, where $\mathcal{E}_{n}$ is the $n$-qubit Pauli basis and $\operatorname{wt}(E)$ the number of coordinates in which $E$ acts nontrivially; whereas the dual enumerator is given by $B_{j}=\sum_{E\in\mathcal{E}_{n}\,:\,\operatorname{wt}(E)=j}=\operatorname{tr}(\Pi E^{\dagger}\Pi E)$. For stabilizer and codeword-stabilized quantum codes, the weight enumerators have combinatorial interpretations similar to their classical counterparts [47, 37].
The introduction of weight enumerators allows for efficiently computable linear and semidefinite programming bounds on the size of classical codes. To understand the origin of these classical code bounds, consider the Lovász number $\vartheta$ of the Hamming graph. A subset of vertices is termed independent, if no two of its vertices are adjacent. The Lovász number $\vartheta$ then bounds the independence number of a graph, that is the largest cardinality of an independent set, as $\alpha(G)\leq\vartheta(G)$. While the computation of $\alpha$ is NP-hard [32], $\vartheta$ can be expressed as semidefinite program [41]. A suitable symmetrization of the semidefinite program formulation of the Lovász theta number then yields the classical Delsarte bound, namely the condition that both the primal and dual enumerators of a code are non-negative. This result relies on the fact that given the indexing vector of a code $\chi$, the matrix $\chi\chi^{T}$ is positive semidefinite and has zero entries when the row and column codewords have a distance in $11,\dots,\delta-$ 1 [22].
We show that for self-dual quantum codes a quantum Delsarte bound arises in a similar fashion [Result B, Section 7], with the added constraint that the matrix also has zeros in rows and columns whose “codewords”, that is Pauli-strings, anticommute. As a consequence, it relates to a Lovász number for a graph with loops, and thus fits into a Lovász framework recently developed for hypergraphs [10]. Finally, following the strategy in Ref.  [23] we give efficiently computable semidefinite programming bounds on the size of quantum codes [Result C, Section 8]. Similar to the classical case, these scale as $O(n^{4})$, where $n$ is the block length, and build on the block-diagonalization of the Terwilliger algebra of the non-binary Hamming scheme. With this, we give an an alternative proof that there is no $(\!(7,1,4)\!)_{2}$ quantum code 111 Note that the original proof in Ref. [30] does not make use of mathematical optimization., as well as prove the non-existence of quantum codes with parameters $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ [Section 9]. Thus the method is strictly stronger than the “enhanced” quantum linear programming bounds that include the shadow bounds by Rains [58, 57], which do not recover this result.
Our second main result is that these bounds are part of a complete semidefinite programming (SDP) hierarchy [Result A, Section 4]. If a code with given parameters $(\!(n,K,\delta)\!)_{2}$ does not exist, then some level of the hierarchy will be infeasible. This hierarchy is based on the state polynomial optimization framework, which provides a sequence of outer approximations to the set of non-linear quantum correlations. Approximating $A_{j}$ one can also recover $B_{j}$ through the quantum MacWilliams identity, and impose the constraints $KB_{i}=A_{i}$ for $0\leq i<\delta$ based on the Knill-Laflamme conditions for quantum error correction. While this machinery is formally dimension-free, a characterization of quasi-Clifford algebras by Gastineau-Hills [21] pulls this result back to $n$-qubit systems.
Overall, our approach is applicable to both stabilizer and non-stabilizer codes. It is based on moment matrices and thus offers high flexibility on the constraints to be included, so that small relaxations can be constructed. Lastly, because our formalism is dimension-free, it can also target scenarios where the Hilbert space is unknown.
1.1. Related work
Lai et al. [37] have introduced an SDP relaxation in the context of entanglement-assisted codeword stabilized codes. When restricted to the case with no pre-shared entanglement, their bounds showed no advantage over the LP bounds. Berta et al. [6] presented a SDP hierarchy to determine the existence of a quantum code based on the symmetric extension hierarchy [15]. While this hierarchy is also complete, it does not easily allow for numerically tractable relaxations. In the classical case, a complete hierarchy for the existence of codes can be formulated using $0-1$ programming for maximum independent sets [42, 38].
2. Contributions
Our aim is to determine the existence of quantum codes with block-length $n$, code distance $\delta$, and the dimension of the code space $K$. A quantum code with parameters $(\!(n,K,\delta)\!)_{2}$ is then able to encode a $K$-dimensional system into $n$ qubits, such that any errors acting on $\lfloor\tfrac{\delta-1}{2}\rfloor$ physical qubits can be corrected. One of the fundamental problems of quantum coding theory is [26, 16]:
Question.
Does there exist a quantum code with given parameters $(\!(n,K,\delta)\!)_{2}$?
To date, partial answers to this questions are known in form of the quantum linear programming bounds and analytical upper bounds. Our main results are the following.
Result A.
A quantum code with parameters $(\!(n,K,\delta)\!)_{2}$ exists if and only if the semidefinite programming hierarchy of Eq. (4.5) is feasible at every level $\ell\in\mathds{N}$.
Result A is shown in Section 4, Theorem 6. The proof is based on the characterization of quantum error correction through the Knill-Laflamme conditions, the quantum MacWilliams identity, and the state polynomial optimization framework to optimize over valid quantum weight enumerators. From an intermediate level of the hierarchy we recover the following for self-dual quantum codes:
Result B.
The Lovász number bounds the existence of self-dual quantum codes. The symmetrization over distance preserving automorphisms of a minor variation of the Lovász bound recovers the quantum Delsarte bound.
Result B is shown in Section 5, Corollary 9; and Section 7, Theorem 14. The proof is based on a confusability graph constructed from the Pauli basis $\mathcal{E}_{n}$ in Definition 8, with additional edges arising from anti-commutation relations. Finally, we provide a symmetry-reduced semidefinite program for general quantum codes with $K\geq 1$:
Result C.
There is a symmetry-reduction of an intermediate level of the semidefinite programming hierarchy based on the Terwilliger algebra of size $O(n^{4})$. This recovers the non-existence of the $(\!(7,1,4)\!)_{2}$ code, and proves the non-existence of $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes.
Result C is shown in Section 8, Theorem 21; and Section 9, Observations 23 and 24. The proof is based on symmetrizing an intermediate level of the hierarchy over distance preserving automorphisms of the Pauli basis $\mathcal{E}_{n}$ that leave the identity invariant. The non-existence proof for the $(\!(7,1,4)\!)_{2}$ code is provided as an infeasibility certificate.
2.1. Proof sketch: SDP hierarchy
We now give a sketch of the proof of Result A. Consider a quantum channel $\mathcal{N}(\varrho)=\sum_{E\in\mathcal{N}}E\varrho E^{\dagger}$.
A quantum code requires that the Knill-Laflamme conditions [34] are met: A projector $\Pi$ corresponds to a quantum code that is able to correct errors induced by $\mathcal{N}$, if and only if
(1) $\Pi E^{\dagger}F\Pi=c_{EF}\Pi\quad\text{for all }\quad E,F\in\mathcal{N}\,,\quad c_{EF}\in\mathds{C}\,.$
In this paper we consider errors $E_{a}$ of a complete orthonormal $n$ qubit tensor-product basis ${\mathcal{E}_{n}}$, having the form $E_{a}=\sigma_{a_{1}}\otimes\dots\otimes\sigma_{a_{n}}$ where $\sigma_{0}=\mathds{1}$ and $\operatorname{tr}(\sigma_{i}^{\dagger}\sigma_{j})=\delta_{ij}$. The weight $\operatorname{wt}(E_{a})$ of an operator $E_{a}$ is the number of positions it acts non-trivially on. A code has distance $\delta$, if all errors of weight $\lfloor\tfrac{\delta-1}{2}\rfloor$ can be corrected. The size of the code space is $K=\operatorname{tr}(\Pi)$ with $\Pi$ being a projector acting on a $n$ qubit Hilbert space.
We now ask whether a code with parameters $(\!(n,K,\delta)\!)_{2}$ exists. This question can formulated as a feasibility problem:
find $\Pi$
subject to $\Pi^{2}=\Pi\,,$
$\operatorname{tr}(\Pi)=K\,,$
(2) $\Pi E_{a}^{\dagger}E_{b}\Pi=c_{ab}\Pi\quad\quad\text{for all}\quad E_{a},E_{b}\in\mathcal{E}_{n}\,:\,\operatorname{wt}(E_{a}^{\dagger}E_{b})<\delta\,,$
where $c_{ab}\in\mathds{C}$. A code with parameters $(\!(n,K,\delta)\!)_{2}$ exists, if and only if the program  2.1 is feasible.
In this work we show how problem (2.1) can be approximated by a semidefinite programming (SDP) hierarchy. The hierarchy is complete, in the sense that a set of parameters for which no code exists will be detected at some level of the hierarchy. Our proof is based on the following observations:
  • A.
    The Knill-Laflamme conditions for a projector $\Pi$ or rank $K$ to correspond to a quantum code are nonlinear in $\Pi$. That is, they can be formulated as the requirement that $KB_{j}(\Pi)=A_{j}(\Pi)$ for all $j=0,\dots,\delta-1$ [55], where
    (3) $A_{j}(\Pi)=\sum_{\begin{subarray}{c}E\in\mathcal{E}_{n}\\ \operatorname{wt}(E)=j\end{subarray}}\operatorname{tr}(E\Pi)\operatorname{tr}(E^{\dagger}\Pi)\,,\quad B_{j}(\Pi)=\sum_{\begin{subarray}{c}E\in\mathcal{E}_{n}\\ \operatorname{wt}(E)=j\end{subarray}}\operatorname{tr}(E\Pi E^{\dagger}\Pi)\,.$
  • B.
    The quantum MacWilliams identity allows to formulate this condition in terms of the $A_{j}(\Pi)$ alone [55]. Define $A(x,y)=\sum_{j=0}^{n}A_{j}(\Pi)x^{n-j}y^{j}$ and likewise for $B(x,y)$. The identity states that the $B_{j}(\Pi)$ can be expressed as linear combinations of the $A_{j}(\Pi)$,
    (4) $B(x,y)$ $=A\Big{(}\frac{x+3y}{2},\frac{x-y}{2}\Big{)}\,.$
We are now left with the task to determine the set of possible $A_{j}(\Pi)$ where $\Pi$ is a projector acting on $(\mathds{C}^{2})^{\otimes n}$.
  • C.
    First note that the conditions for a $n$-qubit state $\varrho$ to be proportional to a projector $\Pi$ of rank $K$ are nonlinear in its expectation values. To see this, note that swap operator exchanging the two tensor factors of $(\mathds{C}^{d})^{\otimes n}\otimes(\mathds{C}^{d})^{\otimes n}$ can be written as $(1,2)=2^{-n}\sum_{E\in\mathcal{E}_{n}}E^{\dagger}\otimes E$ where the sum is over the $n$-qubit Pauli basis $\mathcal{E}_{n}$  [18]. Decomposing cycles into elementary transpositions, a generalization of the swap trick then allows to evaluate traced powers of a state [27],
    (5) $\operatorname{tr}(\varrho^{m})=\operatorname{tr}\big{(}(1,2,\dots,m)\varrho^{\otimes m}\big{)}\,.$
    The fact that $\varrho=\frac{\Pi}{K}$ if and only if $\operatorname{tr}(\varrho^{m})=\frac{1}{K^{m-1}}$ for all $m\in\mathds{N}$ recovers the conditions for $\Pi$ to be a projector.
  • D.
    Second, note that the set of $n$-qubit Pauli operators can be defined algebraically: The characterization of algebras $\mathscr{C}$ over a complex field $\mathds{C}$ whose elements satisfy $A_{i}^{2}=1$ and $A_{i}A_{j}=(-1)^{\delta_{ij}}A_{j}A_{i}$ with $\delta_{ij}\in\{0,1\}$ was given by Gastineau-Hills [21]: These quasi-Clifford algebras $\mathscr{C}$ are isomorphic to the direct sum of Pauli operators acting on $m$-qubit systems,
    (6) $\mathscr{C}\simeq\bigoplus^{2^{r}}_{i=1}(\mathds{C}_{2})^{\otimes m}\,,$
    where $\mathds{C}_{2}$ is the space of complex $2\times 2$ matrices and $m,r\in\mathds{N}$.
    In combination with the constraints of the hierarchy, this restricts the Gelfand-Naimark-Segal construction of the state polynomial optimization framework to states and operators acting on the Hilbert space of $n$ qubits.
  • E.
    Third, the framework of state polynomial optimization [33] (also known as scalar extension [53, 54]) allows us to optimize over expressions that are nonlinear in expectation values of a state, such as $A_{j}(\varrho)=\sum_{E\in\mathcal{E}_{n},\,\operatorname{wt}(E)=j}\operatorname{tr}(E\varrho)\operatorname{tr}(E^{\dagger}\varrho)$. This is done by a variant of the Navascués-Pironio-Acín hierarchy [51] for non-commutative optimization. It consists of a hierarchy of semidefinite programs whose entries converge to monomials in the expectation values of operators on a state.
This way, the state polynomial optimization hierarchy allows us to optimize over the set of possible weight enumerators arising from $n$-qubit quantum codes. This gives a complete semidefinite programming hierarchy for the existence of qubit quantum codes with parameters $(\!(n,K,\delta)\!)_{2}$. That is, a code with given parameters exist if and only if every level of the hierarchy is feasible. We remark that semidefinite programming duality allows to find infeasibility certificates to prove that no code with the given parameters exists.
2.2. Proof sketch: Lovász and quantum Delsarte bounds
We give proof sketches for Results B and  C.
  • A.
    We consider a relaxation of the complete hierarchy of Result B that is indexed by $\langle E_{a}\rangle E_{a}$ with $E_{a}\in\mathcal{E}_{n}$. The key object is a matrix with entries $\Gamma_{ab}=\Re({\langle E_{a}^{\dagger}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E_{b}^{\dagger}\rangle})$. Such matrix is positive semidefinite by construction, and satisfies $\Gamma_{00}=1$, $\Gamma_{a0}=\Gamma_{aa}$, and $\Gamma_{ab}=0$ if $E_{a}E_{b}+E_{b}E_{a}=0$. For pure codes additionally $\Gamma_{ab}=0$ if $0<\operatorname{wt}(E_{a})<\delta$, $0<\operatorname{wt}(E_{b})<\delta$, or $0<\operatorname{wt}(E_{a}E^{\dagger}_{b})<\delta$.
  • B.
    Consider the case of self-dual codes ($K=1$). Then the set of matrices $\Gamma$ with this structure can be identified with the Lovász theta body for a graph with loops, that is the feasible region of Eq. (3.5),
    $\mathrm{TH}(G)=\Big{\{}\operatorname{diag}(M)\,\big{|}\,\begin{pmatrix}1&x^{T}\\ x&M\end{pmatrix}\succeq 0\,,x_{a}=M_{aa}\,\forall a\,,M_{ab}=0\text{ if }a\sim b\Big{\}}\,.$
    A projector onto a self-dual code additionally satisfies $\sum_{E_{a}\in\mathcal{E}_{n}}\Gamma_{aa}=2^{n}$. As a consequence, the Lovász theta number gives a bound on the existence of self-dual quantum codes.
  • C.
    Averaging $\mathrm{TH}(H)$ of the $4$-ary Hamming graph $H$ under distance preserving automorphisms $\operatorname{Aut}(4,n)$ yields, under an additional equality constraint, the quantum Delsarte bound [Eq. (3.4)].
  • D.
    For arbitrary quantum codes, the matrix $\Gamma$ satisfies a range of additional constraints. These arise from the (projective) group structure of the indexing set and from the fact that the code subspace corresponds to a projector $\Pi^{2}=\Pi$. In combination, this yields the main semidefinite programming relaxation in Eq. (5.7).
  • E.
    Using the Terwilliger algebra for $4$-ary codes, we average this main relaxation over distance-preserving automorphisms that keep the zero codeword invariant $\operatorname{Aut}_{0}(4,n)$. A block-diagonalization using the Terwilliger algebra [23] yields a symmetry-reduced semidefinite programming bound of size $O(n^{4})$.
A variation of the Lovász bound recovers the non-existence of the $(\!(7,1,4)\!)_{2}$ code; while a variation of the symmetry-reduced semidefinite programming bounds shows the non-existence of $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$.
3. Key concepts
3.1. Quantum error-correcting codes.
A quantum error-correcting code (QECC) with parameters $(\!(n,K,\delta)\!)_{2}$ encodes a Hilbert space of dimension $K$ in $n$ qubits. A QECC that can correct all errors acting on up to $\lfloor\tfrac{\delta-1}{2}\rfloor$ arbitrary qubit positions is said to have distance $\delta$.
Define a $n$-qubit error basis (or Pauli basis) ${\mathcal{E}_{n}}$ by considering all tensor products of the identity and the three Pauli matrices
(7) $\mathds{1}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\,,\quad X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\,,\quad Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}\,,\quad Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\,.$
The elements of $\mathcal{E}_{n}$ then generate the $n$-qubit Pauli group $P_{n}=\{i^{b}E_{a}\,|\,E_{a}\in\mathcal{E}_{n}\,,b=0,1,2,3\}$. Denote by $\operatorname{wt}(E)$ the weight of an error $E$, which is the number of qubit positions on which an error $E\in{\mathcal{E}_{n}}$ acts non-trivially.
Given a projector onto a QECC subspace $\Pi$, the Knill-Laflamme conditions [34] states that the code has distance $\delta$, if and only if
(8) $\Pi E_{a}^{\dagger}E_{b}\Pi=c_{ab}\Pi\quad\text{for all}\quad\operatorname{wt}(E_{a}^{\dagger}E_{b})<\delta\,,$
where $c_{ab}\in\mathds{C}$. A code is pure if $c_{ab}=\operatorname{tr}(E_{a}^{\dagger}E_{b})$ holds [9, 66]. Note that all codes with $K=1$ are defined to be pure. A code with $K=1$ is termed self-dual.
3.2. Stabilizer codes
Given $n-k$ commuting generators $g_{i}\in P_{n}$, let $S=\langle g_{1},\dots,g_{n-k}\rangle$ be a subgroup of the $n$-qubit Pauli group $P_{n}$ not containing $-\mathds{1}$. The stabilizer $S$ defines a code subspace, whose corresponding projector is given by
(9) $\Pi=\frac{1}{2^{n-k}}\prod^{n-k}_{i=0}(1+g_{i})=\frac{1}{2^{n-k}}\sum_{s_{i}\in S}s_{i}\,.$
In particular, the $s_{i}\in S$ stabilize $\Pi$, so that $s\Pi=\Pi s=\Pi$ for all $s\in S$.
A stabilizer code then encodes $k$ logical qubits into $n$ physical qubits with $n\geq k$ and therefore, $K=2^{k}$. The size of the stabilizer group is $|S|=\frac{2^{n}}{K}=2^{n-k}$ and when $k=0$ ($K=1$), we obtain a pure stabilizer state. For stabilizer codes, the coefficient $c_{ab}$ from Eq. (8) can only be ±$\pm 1$ or 00.
3.3. Quantum weight enumerators
Given a projector $\Pi$ onto a code space, we define two quantum weight enumerators [^2] We follow Rains [55] and omit the normalization factor from the original formulation by Shor and Laflamme [67]. as
(10) $A_{j}(\Pi)=$ $\sum_{\begin{subarray}{c}E\in{\mathcal{E}_{n}}\\ \operatorname{wt}(E)=j\end{subarray}}\operatorname{tr}(E^{\dagger}\Pi)\operatorname{tr}(E\Pi)\,,$
(11) $B_{j}(\Pi)=$ $\sum_{\begin{subarray}{c}E\in{\mathcal{E}_{n}}\\ \operatorname{wt}(E)=j\end{subarray}}\operatorname{tr}(\Pi E^{\dagger}\Pi E)\,,$
where the sum runs over all errors $E$ of weight $j$ in the $n$-qubit Pauli basis ${\mathcal{E}_{n}}$. Let $K=\operatorname{rank}(\Pi)$, then
$A_{j}(\Pi)$ $\geq 0\,,$
$B_{j}(\Pi)$ $\geq 0\,,$
(12) $KB_{j}(\Pi)$ $\geq A_{j}(\Pi)\,,$
for all $j=0,\dots,n$ and $A_{0}(\Pi)=K^{2}$. Self-dual codes ($K=1$) satisfy $A_{j}=B_{j}$ for all $j$. This is seen by expanding the definitions in Eq. (10) in terms of a projector onto a pure state $\Pi=|\psi\rangle\langle\psi|$.
Interestingly, the Knill-Laflamme conditions [see Eq. (8)] can be formulated in terms of the quantum weight enumerators.
Theorem 1.
[55, Theorem 17] A projector $\Pi$ of rank $K$ on $n$ qubits corresponds to a quantum code of distance $\delta$, if and only if
(13) $KB_{j}(\Pi)$ $=A_{j}(\Pi)\quad\text{for}\quad 0\leq j<\delta\,.$
For pure codes, the enumerators additionally fulfill $A_{j}(\Pi)=0$ for $0<j<\delta$. The quantum MacWilliams identity relates the two enumerators through a polynomial transform [55],
(14) $B(x,y)$ $=A\Big{(}\frac{x+3y}{2},\frac{x-y}{2}\Big{)}.$
where $A(x,y)$ is a polynomial defined as $A(x,y)=\sum^{n}_{j=0}A_{j}\,x^{n-j}y^{j}$ and likewise for $B(x,y)$. The coefficients of this polynomial transformation can be expanded as
(15) $B_{j}(\Pi)$ $=2^{-n}\sum^{n}_{i=0}K_{j}(i;n)A_{i}(\Pi)\,,$
with $K_{j}(i):=K_{j}(i;n,4)$ being the quaternary Krawtchouk polynomial defined as [44],
(16) $K_{j}(i)=\sum^{i}_{\alpha=0}(-1)^{\alpha}3^{j-\alpha}\binom{i}{\alpha}\binom{n-i}{j-\alpha}\,.$
Note that, $KB_{0}(\Pi)=A_{0}(\Pi)$ corresponds to $\sum^{n}_{i=0}A_{i}(\Pi)=2^{n}K$. This implies that the weight enumerator of a self-dual code is invariant under the quantum MacWilliams transform.
The quantum shadow enumerator $S_{j}$ can be obtained through the transform [57]
(17) $S(x,y)$ $=A\Big{(}\frac{x+3y}{2},\frac{y-x}{2}\Big{)}\,.$
Its coefficients expand as
(18) $S_{j}(\Pi)$ $=2^{-n}\sum^{n}_{i=0}(-1)^{i}K_{j}(i;n)A_{i}(\Pi)\,.$
It is known that $S_{j}\geq 0$ for all $j=0,\dots,n$, giving the so-called shadow bounds on quantum codes [56, 57]. If $K=1$ and $n-j$ odd then $S_{j}=0$ [57, Theorem 12].
In the case of qubit stabilizer codes, one has additionally the constraint that $\sum_{j\geq 0}A_{2j}=2^{n-\log_{2}(K)-1}$ or $\sum_{j\geq 0}A_{2j}=2^{n-\log_{2}(K)}$ for codes of type I and II respectively [9].
3.4. Linear programming bounds
The linear programming bounds are given by the conditions from Eqs. (3.3)–(18[67, 55],
find $A_{j},\,B_{j},\,S_{j}\geq 0\,,$
subject to $A_{0}=K^{2}\,,$
$KB_{j}\geq A_{j}\quad\quad\quad\text{for }j=0,\dots,n\quad\text{with equality for $j<\delta$}\,,$
[…]
For given parameters $(\!(n,K,\delta)\!)_{2}$, if the linear program in Eq. (3.4) is infeasible, then a corresponding quantum codes does not exist.
In analogy to the classical case [13, 71], we will call the following subset of conditions the quantum Delsarte linear programming bounds for self-dual codes ($K=1$),
$\eta=\text{max}\quad$ $\sum^{n}_{j=0}A_{j}\,,$
subject to $A_{0}=1\,,$
$A_{j}\geq 0\ \,\,\text{with equality for }\,\,1<j<\delta\,,$
(20) $\sum^{n}_{i=0}K_{j}(i)A_{i}\geq 0\ \quad\text{for }j=0,\dots,n\,.$
If one has $\eta<2^{n}$ in Eq. (3.4), a code with parameters $(\!(n,1,\delta)\!)_{2}$ does not exist. We will formally derive this bound from a Lovász bound for self-dual quantum codes in Section 7.
Remark 2.
Note that for $K>1$ and without adding extra constraints like the quantum MacWilliams identity or the Knill-Laflamme conditions (that is $B_{j}=\sum_{i}K_{j}(i;n)A_{i}$ for $j=0,\dots,n$ and $KB_{j}=A_{j}$ for $j=0,\dots,\delta-1$) the quantum Delsarte bounds becomes weaker.
3.5. Lovász theta number of a graph
Let $G$ be a simple graph defined by a set of $N$ vertices $V$ connected by a set of edges $E$. Then the Lovász theta number can take the following two semidefinite programming formulations [20]. We call the following SDP1 for the Lovász number:
$\vartheta(G)=\max_{M}\quad$ $\sum_{i,j=1}^{N}M_{ij}\,,$
subject to $\operatorname{tr}(M)=1\,,$
$M_{ij}=0\quad\,\,\text{for all}\quad(i,j)\in E\,,$
(21) $M\succeq 0\,.$
We call the following SDP2 for the Lovász number:
$\vartheta(G)=\max_{M}\quad$ $\sum^{N}_{i=1}M_{ii}\,,$
subject to $M_{ii}=a_{i}\quad\,\text{for all}\quad i\in V\,,$
$M_{ij}=0\quad\,\,\text{for all}\quad(i,j)\in E\,,$
(22) $\Delta=\begin{pmatrix}1&a^{T}\\ a&M\end{pmatrix}\succeq 0\,.$
The Lovász theta number upper bounds the independence number of a graph, that is, the maximum cardinality of a subset of vertices that do not share any edge, by $\alpha(G)\leq\vartheta(G)$. In turn, this also bounds the Shannon capacity of a graph [41]. Other mathematical formulations of the Lovász theta number exist [36, 52, 20].
A slight variation of SDP1 from Eq. (3.5) with the additional entry-wise non-negativity constraint $M_{ij}\geq 0$ for all $i,j\in V$, defines $\vartheta^{\prime}(G)$. In the case of the Hamming cube, the computation of $\vartheta^{\prime}(G)$ can be shown to reduce to the classical Delsarte bound [64, 71],
$\eta_{C}=\max\quad$ $2^{n}\sum_{i=0}^{n}x_{i}$
subject to $x_{0}=\frac{1}{2^{n}}\,,$
$x_{1}=\dots=x_{\delta-1}=0\,,$
$x_{\delta},\dots,x_{n}\geq 0\,,$
(23) $\sum_{i}K_{j}(i;n,2)x_{i}\geq 0\quad\text{for all }j=0,\dots,n\,.$
with $K_{j}(i;n,2)$ the binary Krawtchouk polynomial defined as [44],
(24) $K_{j}(i;n,2)=\sum^{i}_{\alpha=0}(-1)^{\alpha}\binom{i}{\alpha}\binom{n-i}{j-\alpha}\,.$
3.6. State polynomial optimization
Given a state $\varrho$ and a set of observables $\{X_{i}\}^{m}_{i=1}$, denote by ${\langle X_{i}\rangle}_{\varrho}=\operatorname{tr}(X_{i}\varrho)$ their expectation values on $\varrho$. Consider now polynomials in products of operators and their expectations e.g., ${\langle(X_{1})^{4}\rangle}_{\varrho}\,X_{2}\,+\,(X_{3})^{2}$. We term these state polynomials in the variables $X_{i}$.
An optimization problem over the set of quantum states $\varrho$ and operators $X_{i}$ on a Hilbert space $\mathcal{H}$ with an objective state polynomial function $f$ and constraints $g_{j}$ then reads
(25) $\alpha^{*}=\max_{\mathcal{H},\{X_{i}\},\varrho}$ $\langle\,f(X_{1},\dots,X_{m})\,\rangle_{\varrho}$
subject to $g_{j}(X_{1},\dots,X_{m})\succeq 0\quad\text{for}\quad j=0,\dots,r\,.$
Note that Eq. (25) requires an optimization over the Hilbert space $\mathcal{H}$ and all operators and quantum states with support on $\mathcal{H}$ satisfying the inequality constraints. The framework of state polynomial optimization [33] (also known as scalar extension [53, 54]) provides a hierarchy of semidefinite programming approximations to Eq. (25) that converge.
To see how this works, consider a set of (non-commutative) letters $\{x_{i}\}^{m}_{i=1}$ and form words $w=x_{j_{1}}\dots x_{j_{m}}$. We associate to each word $w$ a scalar with the symbol ${\langle w\rangle}={\langle x_{j_{1}}\dots x_{j_{k}}\rangle}$. We refer to ${\langle w\rangle}$ as the “expectation value” of a word. These satisfy $w{\langle v\rangle}={\langle v\rangle}w$ and $({\langle v\rangle}w)^{*}={\langle v^{*}\rangle}w^{*}$ for all words $v,w$. Furthermore, denote the empty word as $1$, satisfying $1w=w1=w$ for all words $w$. The degree of a word is defined as the number of letters in it. The involution of a word is given by $(x_{1}\cdots x_{m})^{*}=(x^{*}_{m}\cdots x^{*}_{1})$. Products of words and expectation values of words are known as state monomials, e.g. $w_{i_{1}}{\langle w_{i_{2}}\rangle}\dots\langle w_{i_{m}}\rangle$. The degree of a state monomial is the sum over the degree of all words that compose it, including the words within expectation values. State polynomials are formed by a linear combination of monomials. The degree of a state polynomial is the largest degree of the monomials that it is composed of.
Naturally, state polynomials behave in the same manner as operators. A word $x_{i}$ and its expectation value ${\langle x_{i}\rangle}$ correspond to an operator $X_{i}$ and its expectation value ${\langle X_{i}\rangle}_{\varrho}$ from a state $\varrho$. In particular, the empty word $\mathds{1}$ corresponds to the identity, since $\mathds{1}W_{i}=W_{i}\mathds{1}=W_{i}$. Similarly, a state polynomial like $w={\langle x_{i}\rangle}x_{j}$ corresponds to $W={\langle X_{i}\rangle}_{\varrho}X_{j}$. However, a set of values ${\langle x_{i}\rangle}$ does not necessarily have to be compatible with a quantum state, i.e., it can be that there does not exist a $\varrho$ such that ${\langle x_{i}\rangle}\neq{\langle x_{i}\rangle}_{\varrho}$.
Let $\vec{w}^{(\ell)}$ be a vector that contains all state monomials up to degree $\ell$. Then, we can define the matrix $M^{(\ell)}_{ij}={\langle{w_{i}^{*}}^{(\ell)}w_{j}^{(\ell)}\rangle}$ with $M^{(\ell)}_{00}=1$. It is important to remark that if $M^{(\ell)}_{ij}$ comes from a state, then it can be written as
(26) $M^{(\ell)}_{ij}=\operatorname{tr}({W_{i}^{\dagger}}^{(\ell)}W_{j}^{(\ell)}\varrho)\,.$
Then, $M^{(\ell)}$ is a positive semidefinite matrix with $M^{(\ell)}_{00}=1$. This can be seen by writing the spectral decomposition of the state $\varrho=\sum_{v}p_{v}|\phi_{v}\rangle\langle\phi_{v}|$ with $p_{v}\geq 0$; then for each $\mathinner{|\phi_{v}\rangle}$ form the vector
(27) $\phi_{i}=\big{(}W^{(\ell)}_{1}\mathinner{|\phi_{v}\rangle},W^{(\ell)}_{2}\mathinner{|\phi_{v}\rangle},\dots,W^{(\ell)}_{m}\mathinner{|\phi_{v}\rangle}\big{)}^{\dagger}\,.$
It is then clear that $M^{(l)}=\sum_{i}p_{i}\phi^{\dagger}_{i}\phi_{i}\succeq 0$. Likewise a matrix defined by
(28) $G^{(\ell-q_{k})}_{ij}=\operatorname{tr}({W_{i}^{\dagger}}^{(l)}g_{k}(X_{1},\dots,X_{m})W_{j}^{(l)}\varrho)\,,$
with $q_{k}:=\lceil\frac{\text{deg}(g_{k})}{2}\rceil$ is positive semidefinite when $g_{k}(X_{1},\dots,X_{m})$ is positive semidefinite. This can be seen from the fact that a matrix $g_{k}\succeq 0$ is decomposable as a product $g_{k}=b_{k}b_{k}^{\dagger}$ of a matrix $b_{k}$.
Note that $G^{(\ell-q_{k})}$ are sub-matrices of $M^{(\ell)}$.
The optimization problem from Eq. (25) is then approximated by the following semidefinite programming hierarchy indexed by $\ell\in\mathds{N}$ [33, 54, 53],
(29) $\alpha_{\ell}=\max_{M^{(\ell)}}$ $\langle\,f(x_{1},\dots,x_{m})\,\rangle\,,$
subject to $M^{(\ell)}_{00}=1\,,$
$M^{(\ell)}\succeq 0\,,$
$G^{(\ell-q_{k})}\succeq 0\,,$
$M^{(\ell)}_{ij}=M^{(\ell)}_{\alpha\beta}\quad\quad\text{if}\quad{\langle w_{i}^{*}w_{j}\rangle}={\langle w_{\alpha}^{*}w_{\beta}\rangle}\,,$
$G^{(\ell-q_{k})}_{ij}=M^{(\ell)}_{\alpha\beta}\quad\text{if}\quad{\langle w_{i}^{*}g_{k}w_{j}\rangle}={\langle w_{\alpha}^{*}w_{\beta}\rangle}\,,$
for $1\leq k\leq r$ and with $\ell$ greater than or equal to the degree of the polynomial $f$.
Generally $\alpha_{\ell}\neq\alpha^{*}$ since the constraints from Eq. (29) do not imply that $\alpha_{\ell}$ comes from a state $\varrho$ and operators $X_{i}$ satisfying the constraints $g_{i}\succeq 0$. However, when the Archimedeanity condition is met, i.e., there exists a $N\geq 0$ such that $N-\sum^{m}_{i=0}x^{*}_{i}x_{i}\geq 0$, the relaxation converges to the optimum $\alpha^{*}$ of Eq. (25[33, Proposition 6.7],
(30) $\lim_{\ell\rightarrow\infty}\alpha_{\ell}=\alpha^{*}\,.$
A variant of the Gelfand-Naimark-Segal (GNS) construction then allows us to construct a direct sum of states $\varrho$ and operators $\{X_{i}\}$ on a Hilbert space $\mathcal{H}$ from the sequence $\{M^{(\ell)}\}_{\ell=0}^{\infty}$ such that each block satisfies the constraints of the optimization problem [33, 68].
3.7. Quasi-Clifford algebra
Let $F$ be a field of characteristic not equal to $2$ and $m$ a positive integer. A quasi-Clifford (QC) algebra is defined as an associative algebra over $F$ containing the identity $1$ on $m$ generators $\alpha_{1},\dots,\alpha_{m}$, such that $\alpha_{i}^{2}=k_{i}$ and $\alpha_{j}\alpha_{i}=(-1)^{\delta_{ij}}\alpha_{i}\alpha_{j}$ with $k_{i}\in F$. Denote by $\mathbb{C}_{b}$ a QC algebra on one generator $\beta$ with $\beta^{2}=b$ and by $\mathbb{Q}_{cd}$ a QC algebra on two generators $\gamma$, $\delta$ with $\gamma^{2}=c$, $\delta^{2}=d$, and $\gamma\delta=-\delta\gamma$.
Theorem 3.
[21, Theorem 2.7] A QC algebra $\mathscr{C}$ with $m$ generators is isomorphic to
(31) $\mathscr{C}\cong\mathbb{C}_{b_{1}}\otimes\dots\otimes\mathbb{C}_{b_{r}}\otimes\mathbb{Q}_{c_{1}d_{1}}\otimes\dots\otimes\mathbb{Q}_{c_{s}d_{s}}\,,$
where $m=r+2s$ with $r$, $s>0$ and $b_{i}$, $c_{j}$, $d_{k}$ are plus or minus products of some $k_{i}$.
When $k_{i}=1$ for all $i$ and $F$ is the field of complex numbers, then the QC algebra $\mathscr{C}$ contains only ±$\mathbb{C}_{\pm 1}$ and ±$\mathbb{Q}_{\pm 1\pm 1}$ as tensor factors. Ref. [21, Eq. (3.1)] lists the irreducible representation of the generators of these algebras
(32) $\begin{matrix}[2]\text{Field}&\text{Generators}\\ \hline\cr\mathbb{C}_{1}&\beta\rightarrow 1\quad\text{or}\quad\beta\rightarrow-1\\ \mathbb{C}_{-1}&\beta\rightarrow i\quad\text{or}\quad\beta\rightarrow-i\\ \mathbb{Q}_{\pm 1,1}&\gamma\rightarrow\begin{pmatrix}[1]0&\pm 1\\ 1&0\end{pmatrix},\quad\delta\rightarrow\begin{pmatrix}[1]1&0\\ 0&-1\end{pmatrix}\\ \mathbb{Q}_{1,-1}&\gamma\rightarrow\begin{pmatrix}[1]0&\phantom{-}1\\ 1&\phantom{-}0\end{pmatrix},\quad\delta\rightarrow\begin{pmatrix}[1]0&-1\\ 1&0\end{pmatrix}\\ \mathbb{Q}_{-1,-1}&\gamma\rightarrow\begin{pmatrix}[1]0&-1\\ 1&0\end{pmatrix},\quad\delta\rightarrow\begin{pmatrix}[1]i&0\\ 0&-i\end{pmatrix}\\ \end{matrix}$
Furthermore, Ref. [21, Eq. (2.4)] also shows that ±$\mathbb{C}_{\pm 1}\cong 2F$ and ±$\mathbb{Q}_{\pm 1\pm 1}\cong F_{2}$, where $2F=F\oplus F$ and $F_{2}$ is the algebra of complex $2\times 2$ matrices. This allows us to state the following corollary.
Corollary 4.
Let $\{\alpha_{1},\dots,\alpha_{m}\}$ be a set of $m$ operators from a Hilbert space over $\mathbb{C}$ with $\alpha^{2}_{i}=1$ and $\alpha_{j}\alpha_{i}=(-1)^{\delta_{ij}}\alpha_{i}\alpha_{j}$ with $\delta_{ij}\in\{0,1\}$. Then, these operators are isomorphic, up to phases, to a direct sum of $2^{r}$ $s$-qubit error basis with $m=r+2s$.
Proof.
The set of operators $\{\alpha_{1},\dots,\alpha_{m}\}$ defines a QC algebra $\mathscr{C}$ with $k_{i}=1$ for all $i$. Therefore, according to Theorem 3,
(33) $\mathscr{C}\cong\mathbb{C}_{b_{1}}\otimes\dots\otimes\mathbb{C}_{b_{r}}\otimes\mathbb{Q}_{c_{1}d_{1}}\otimes\dots\otimes\mathbb{Q}_{c_{s}d_{s}}\,,$
with $b_{i}$, $c_{j}$, $d_{k}\in\{-1,1\}$. Using the fact that ±$\mathbb{C}_{\pm 1}\cong 2F$ and ±$\mathbb{Q}_{\pm 1\pm 1}\cong F_{2}$, one has
(34) $\mathscr{C}\cong 2F\otimes\cdots\otimes 2F\otimes F_{2}\otimes\cdots\otimes F_{2}=2^{r}F_{2^{s}}\,.$
Here we used that $\otimes$ distributes over $\oplus$ and $F_{m}\otimes F_{n}\cong F_{nm}$. Note that $F_{2^{r}}$ is the tensor product of $2^{s}$ complex matrices of size $2\times 2$. Therefore, these operators $\{\alpha_{1},\dots,\alpha_{m}\}$ are isomorphic to a direct sum of $s$-qubit operators. According to Eq. (32), these operators are given by a tensor product of Pauli matrices [see Eq. (7)], up to phases. This ends the proof. ∎
The direct sum in Eq. (34) comes from the fact that given a set of operators with some commutation/anti-commutation relations and squaring to $\mathds{1}$, the mapping to Pauli operators is not unique. For example, given $x_{1},x_{2},x_{3}$ with $x_{i}x_{j}=-x_{j}x_{i}$ and $x^{2}_{i}=\mathds{1}$, the elements $x_{1}$, $x_{2}$, $x_{3}$ can be mapped to the Pauli matrices $X$, $Z$, $Y$ but also to $X$, $Z$, $-Y$, respectively. Different blocks from the direct sum give thus different mappings to Pauli operators.
3.8. Semidefinite programming
A semidefinite program optimizes a linear function over the self-dual cone of positive semidefinite matrices under a set of linear matrix constraints. Denote by $S^{k}$ the set of real symmetric matrices of size $k$. We now follow the exposition from Boyd and Vandenberghe [7, Section 4.6.2, Example 5.12] for the cone of positive semidefinite matrices, with the Hilbert-Schmidt inner product $\langle A,B\rangle=\operatorname{tr}(A^{\dagger}B)$.
Consider a semidefinite program in standard form
minimize $\quad\operatorname{tr}(CX)$
subject to $\quad\operatorname{tr}(A_{i}X)=b_{i}\quad\text{for}\quad i=1,\dots,p\,,$
(35) $\quad X\succeq 0\,.$
where $A_{1},\dots,A_{n},C\in S^{k}$ and $b_{i}\in\mathds{R}$. Eq. (3.8) is called the primal program and $\operatorname{tr}(CX)$ is its objective function. To each primal program is associated a dual program which reads
maximize $\sum^{p}_{i=1}y_{i}b_{i}$
(36) subject to $C-\sum^{p}_{i=1}y_{i}A_{i}\succeq 0\quad\text{for}\quad i=1,\dots,p\,.$
A variable $X$ (or set of $y_{i}$) satisfying the constraints of the primal (or dual) problem is said to be primal (or dual) feasible. Any pair of feasible $X$ and $\{y_{i}\}$ satisfy the weak duality:
(37) $\operatorname{tr}(CX)-\sum^{p}_{i=1}y_{i}b_{i}\geq 0\,.$
An optimization problem is converted to a feasibility problem by setting $C=0$. The weak duality can then be used to prove primal infeasibility: if there exists a dual solution $\{y_{i}\}$ with $\sum^{p}_{i=1}y_{i}b_{i}\geq 0$, then Eq. (37) is violated and the primal problem is infeasible.
It is straightforward to embed complex semidefinite programs in the above formalism, by encoding a complex hermitian matrix $X$ as,
(38) $\tilde{X}=\begin{pmatrix}\operatorname{Re}(X)&-\operatorname{Im}(X)\\ \operatorname{Im}(X)&\operatorname{Re}(X)\end{pmatrix},$
where $\operatorname{Re}(X)$ and the $\operatorname{Im}(X)$ are the real and the imaginary part of $X$, respectively. Then $X\succeq 0$ if and only if $\tilde{X}\succeq 0$
4. Complete SDP hierarchy bounding quantum codes
Our aim is now to express the conditions for the existence of quantum codes as a state polynomial optimization problem. For this we formulate the Knill-Laflamme conditions, the quantum MacWilliams identity, and the constraints for the state to be proportional to a projector of rank $K$ as state polynomial constraints. This gives rise to a complete hierarchy on the existence of an $(\!(n,K,\delta)\!)_{2}$ code.
4.1. Framework
Consider the $n$-qubit error basis ${\mathcal{E}_{n}}$ formed by tensor products of Pauli matrices [see Eq. (7)]. Define a set of letters $\mathscr{E}_{n}$ so that to each Pauli error $E\in{\mathcal{E}_{n}}$ we associate a letter $e\in\mathscr{E}_{n}$. We require this to be a group homomorphism,
(39) $e_{a}e_{b}=\omega_{abc}e_{c}\quad\text{if}\quad E_{a}E_{b}=\omega_{abc}E_{c}\quad\quad\text{for all $a,b,c$}\,.$
In particular, this implies that $e_{a}^{*}e_{a}=1$ where the empty word $1$ corresponds to identity operator $\mathds{1}$, and for the Pauli group it holds that $\omega_{abc}\in\{\pm 1,\pm i\}$.
Let $M^{(\ell)}$ be the moment matrix from the state polynomial optimization framework [see Eq. (29)], indexed by all state monomials of degree at most $\ell$. We impose all constraints from Eq. (39) on $M^{(\ell)}$. Recall that the entries of the state moment matrix are denoted by $M^{(\ell)}_{vw}={\langle v^{*}w\rangle}$. In order to simplify the notation, we define
(40) $M^{(\ell)}({\langle v^{*}w\rangle}):=M^{(\ell)}_{vw}\,.$
We now formulate all the conditions for a quantum code in terms of constraints on state polynomials, that is, as constraints on the entries of $M^{(\ell)}$.
4.2. Quantum MacWilliams identity
The quantum weight enumerator $A_{j}(\varrho)$ in Eq. (10) is a state polynomial in $\varrho$. We can then define an approximation to the quantum weight enumerator as
(41) $A_{j}^{(\ell)}=\sum_{\begin{subarray}{c}e\in\mathscr{E}_{n}\\ \operatorname{wt}(e)=j\end{subarray}}M^{(\ell)}({\langle e^{*}\rangle}{\langle e\,\rangle})\,,$
so to correspond to $A_{j}(\varrho)=\sum_{E\in{\mathcal{E}_{n}}\,,\operatorname{wt}(E)=j}\operatorname{tr}(E^{\dagger}\varrho)\operatorname{tr}(E\varrho)$. However, the quantum weight enumerator $B_{j}(\varrho)=\sum_{E\in{\mathcal{E}_{n}}\,,\operatorname{wt}(E)=j}\operatorname{tr}(\varrho E^{{\dagger}}\varrho E)$ does not have such a straightforward interpretation as a state polynomial. But recall that the quantum MacWilliams identity [see Eq. (14)] linearly transforms the $A_{i}(\varrho)$ to the $B_{j}(\varrho)$ enumerators. We can then linearly transform $A^{(\ell)}_{j}(\varrho)$ to $B^{(\ell)}_{j}(\varrho)$ using
(42) $B^{(\ell)}_{j}$ $=2^{-n}\sum^{n}_{i=0}K_{j}(i;n)A^{(\ell)}_{i}\,,$
where $K_{j}(i;n)$ are the Krawtchouk polynomials defined in Eq. (16). Therefore, $B^{(\ell)}_{j}$ provides an approximation to the quantum weight enumerator $B_{j}(\varrho)$.
Note that the $A^{(\ell)}_{j}$ are not necessary compatible with a quantum state, i.e., in general there does not need to be a $\varrho$ such that $A^{(\ell)}_{j}=A_{j}(\varrho)$. However, the state polynomial optimization hierarchy provides a sequence $\{A_{j}^{(\ell)}\}_{\ell=1}^{\infty}$ whose limit converges to an enumerator $A_{j}(\varrho)$ and as a consequence, the sequence $\{B_{j}^{(\ell)}\}_{\ell=1}^{\infty}$ also converges to $B_{j}(\varrho)$.
4.3. Knill-Laflamme conditions
We now impose the Knill-Laflamme conditions as constraints on $M^{(\ell)}$. These conditions can be formulated in terms of quantum weight enumerators. Likewise, we demand that
(43) $KB^{(\ell)}_{j}=A^{(\ell)}_{j}\quad\text{for }\quad 0\leq j<\delta\,.$
Remark 5.
One can additionally include the remaining linear programming constraints shown in Eq. (3.4),
$A^{(\ell)}_{j},\,B^{(\ell)}_{j},\,S^{(\ell)}_{j}$ $\geq 0\,,$
(44) $KB^{(\ell)}_{j}$ $\geq A^{(\ell)}_{j}\,.$
Here $S^{(\ell)}_{j}$ is an approximation of the shadow enumerator obtained by Eq. (18). These extra constraints do not have an impact on completeness.
4.4. Projector constraints
We now impose conditions for a state $\varrho$ to be proportional to a projector of rank $K$ on $M^{(\ell)}$. The state $\varrho$ then fulfills for all $m\in\mathds{N}$,
(45) $\operatorname{tr}(\varrho^{m})=\frac{1}{K^{m-1}}\,.$
The converse also holds, namely that a state $\varrho$ satisfying Eq. (45) for all $m\in\mathds{N}$ must be proportional to a projector of rank $K$. To see this, note that the set of $\{\operatorname{tr}(A^{m})\,:\,m=1,\dots,r\}$ determines the characteristic polynomial of a square matrix $A$ of size $r$. The zeros of this characteristic polynomial then determine the eigenvalues of $A$. Fixing the moments as in Eq. (45) implies that the state $\varrho$ is proportional to a projector of rank $K$.
We now show how to write Eq. (45) in terms of state polynomials. For this we decompose the cyclic permutation $(1,\dots,m)$ into elementary transpositions,
(46) $(1,\dots,m)=(m,\,m-1)\dots(2,3)(1,2)\,.$
Second, expand the swap operator through the Pauli basis of $(\mathds{C}^{2})^{\otimes n}$ [72],
(47) $(1,2)={2^{-n}}\sum_{E\in{\mathcal{E}_{n}}}E\otimes E^{\dagger}\,.$
The generalization of the swap trick [27] then gives
$\operatorname{tr}(\varrho^{m})$ $=\operatorname{tr}\big{(}(1,\dots,m)\varrho^{\otimes m}\big{)}$
(48) $=2^{-n(m-1)}\!\!\!\!\!\!\sum_{E_{1},\dots,E_{m-1}\in{\mathcal{E}_{n}}}\operatorname{tr}(E_{1}\varrho)\operatorname{tr}(E^{\dagger}_{1}E_{2}\varrho)\dots\operatorname{tr}(E^{\dagger}_{m-2}E_{m-1}\varrho)\operatorname{tr}(E^{\dagger}_{m-1}\varrho)\,.$
Therefore, Eq. (45) is now expressed as
(49) $\sum_{E_{1},\dots,E_{m-1}\in{P}_{n}}\operatorname{tr}(E^{\dagger}_{1}\varrho)\operatorname{tr}(E_{1}E^{\dagger}_{2}\varrho)\dots\operatorname{tr}(E_{m-2}E^{\dagger}_{m-1}\varrho)\operatorname{tr}(E_{m-1}\varrho)=\frac{2^{n(m-1)}}{K^{m-1}}\,.$
Denote by $\mathbf{e}=(e_{1},\dots,e_{m})\in(\mathscr{E}_{n})^{m}$ vectors of length $m$ with entries in $\mathscr{E}_{n}$. Define the state monomial
(50) ${Q}_{(1,\dots,m)}(\mathbf{e})=\begin{cases}{\langle e_{1}\rangle}{\langle e^{*}_{1}e_{2}\rangle}\dots\langle e^{*}_{{m-2}}e_{m-1}\rangle{\langle e^{*}_{m-1}\rangle}\quad&\text{if}\quad m\geq 2\,,\\ {\langle 1\rangle}&\text{if}\quad m=1\,.\end{cases}$
Then, the projector constraints from Eq. (45) can be expressed in terms of state polynomials as the condition that for all $m\in\mathds{N}$,
(51) $\sum_{\textbf{e}\in(\mathscr{E}_{n})^{m}}{Q}_{(1,\dots,m)}(\mathbf{e})=\frac{2^{n(m-1)}}{K^{m-1}}\,.$
One can generalize the monomial of Eq. (50) to all elements of $S_{2\ell}$.
Decompose $\sigma\in S_{2\ell}$ into cycle permutations including one-cycles such that,
(52) $\sigma=(\alpha_{1},\dots,\alpha_{r})\dots(\xi_{1},\dots,\xi_{t})\,.$
Note that $r+\dots+t=2\ell$ since $\sigma\in S_{2\ell}$. Denote by $|\sigma|$ the number of cycles of $\sigma$, including cycles of length $1$. Defining
(53) $P^{(\ell)}_{\sigma}:=\sum_{\mathbf{e_{r}}\in(\mathscr{E}_{n})^{r},\dots,\mathbf{e_{t}}\in(\mathscr{E}_{n})^{t}}\!\!\!\!\!M^{(\ell)}(Q_{(\alpha_{1},\dots,\alpha_{r})}(\mathbf{e_{r}})\dots Q_{(\xi_{1},\dots,\xi_{t})}(\mathbf{e_{t}}))\,,$
we can then impose the constraints from Eq. (53) on $M^{(\ell)}$ as
(54) $P^{(\ell)}_{\sigma}=\Big{(}\frac{2^{n}}{K}\Big{)}^{(r-1)+\dots+(t-1)}=\Big{(}\frac{2^{n}}{K}\Big{)}^{(2\ell-|\sigma|)}\,.$
Consider a $n$-qubit quantum state $\varrho$ and the set of Pauli operators $\mathcal{E}_{n}$, and construct a moment matrix $M^{(\ell)}_{vw}=\operatorname{tr}(E_{v}^{\dagger}E_{w}\varrho)$ indexed by all state monomials in $\varrho$ of at most degree $\ell$. Then $P^{(\ell)}_{\sigma}=\operatorname{tr}(\varrho^{r})\cdots\operatorname{tr}(\varrho^{t})$ for all $\sigma\in S_{2\ell}$, $\ell\geq 2^{n}$ are necessary and sufficient conditions that $\varrho=\Pi/K$ with $\Pi$ a projector of rank $K$ [see Eq. (45)].
4.5. An SDP hierarchy for qubit codes
We now formulate a semidefinite programming hierarchy that contains the necessary and sufficient conditions for a state to correspond to a quantum code: the first condition is the projector constraints that guarantee that the state is proportional to a projector. The second condition is the Knill-Laflamme conditions, which ensure that this projector corresponds to a subspace with error correction capabilities.
Consider the following hierarchy indexed by $\ell\in\mathds{N}$:
find $M^{(\ell)}\succeq 0$
subject to $M^{(\ell)}_{vw}=M^{(\ell)}_{\alpha\beta}\quad\text{if}\quad{\langle v^{*}w\rangle}={\langle\alpha^{*}\beta\rangle}\,,$
$M^{(\ell)}(1)=1\,,$
$KB^{(\ell)}_{j}=A^{(\ell)}_{j}\quad\text{for}\quad 0\leq j<\delta,$
$P^{(\ell)}_{\sigma}=\Big{(}\frac{2^{n}}{K}\Big{)}^{2\ell-|\sigma|}\quad\text{for all}\,\,\,\sigma\in S_{2\ell},$
(55) $A^{(\ell)}$, $B^{(\ell)}_{j}$, $P^{(\ell)}_{\sigma}$ are given by Eqs. (41), (42), (53) respectively.
For every $\ell\in\mathds{N}$, the SDP (4.5) is a relaxation of the optimization problem in Eq. (2.1). This hierarchy allows us to state the following theorem.
Theorem 6.
A quantum code with parameters $(\!(n,K,\delta)\!)_{2}$ exists if and only if the semidefinite programming hierarchy of Eq. (4.5) is feasible at every level $\ell\in\mathds{N}$.
Proof.
($\Longrightarrow$): Let $\varrho$ be a state proportional to the projector of a quantum code with parameters $(\!(n,K,\delta)\!)_{2}$. Given the set of Pauli operators $\mathcal{E}_{n}$, we can construct a moment matrix $M^{(\ell)}_{vw}=\operatorname{tr}(E_{v}^{\dagger}E_{w}\varrho)$ indexed by all state monomials in $\varrho$ of at most degree $\ell$.
It is clear that for all $\ell\in\mathds{N}$, $M^{(\ell)}\succeq 0$ and $M^{(\ell)}(\mathds{1})=1$ [see Section 3.6]. Additionally, $M$ satisfies the projector conditions in Eq. (54) as $\varrho$ is a density matrix proportional to a projector of rank $K$. Finally, $M$ satisfies the state polynomial quantum MacWilliams identity (42) and Knill-Laflamme conditions in enumerator form (43). Therefore, if there exists a code $(\!(n,K,\delta)\!)_{2}$, every level of the hierarchy is feasible.
[…]
This state $\varrho$ has three desired properties: First, $\varrho$ is an $n$-qubit state. Recall that, in general, state polynomial optimization optimizes over all compatible Hilbert spaces, operators, and states. In our setting however, Corollary 4 restricts the set of operators to the direct sum of tensor products of Pauli matrices. To see this, consider the GNS construction from the sequence $\{M^{(\ell)}\}^{\infty}_{\ell=1}$: given the sequence, it constructs a set of operators and a state satisfying all relations of the optimization problem. In particular, the constructed operators form a quasi-Clifford algebra over $\mathds{C}$, whose elements square to the identity. Corollary 4 then implies that these operators are isomorphic to the $n$-qubit Pauli basis $\mathcal{E}_{n}$. As a consequence, $\varrho$ is a $n$-qubit state.
Second, $\varrho$ has to be proportional to a projector $\Pi$ of rank $K$. The projector conditions from Eq. (54) are included in the hierarchy via $P^{(\ell)}_{\sigma}$. Since the operators are in the limit isomorphic to the $n$-qubit Pauli basis, the sequence $\{P^{(\ell)}_{\sigma}\}^{\infty}_{\ell=1}$ converges to $\operatorname{tr}(\varrho^{r})\cdots\operatorname{tr}(\varrho^{t})$ with $\varrho$ being a $n$-qubit state. Eq. (54) are now necessary and sufficient conditions for $\varrho$ be proportional to the projector $\Pi$ of rank $K$.
Third, the projector $\Pi=\varrho K$ necessarily satisfies the Knill-Laflamme conditions [see Theorem 1]. Their approximation is included in the hierarchy as $KB_{j}^{(\ell)}=A_{j}^{(\ell)}$ for all $0\leq j<\delta$, becoming exact in the limit $\ell\to\infty$. Similarly to $P^{(\ell)}_{\sigma}$, the sequence $\{A_{j}^{(\ell)}\}_{\ell=1}^{\infty}$ from Eq. (41) converges to the enumerator $A_{j}(\tfrac{\Pi}{K})$. The quantum MacWilliams identity [see Eq. (42)] linearly relates $A^{(\ell)}_{j}$ with $B^{(\ell)}_{j}$, so that the sequence $\{B_{j}^{(\ell)}\}_{\ell=1}^{\infty}$ also converges to $B_{j}(\tfrac{\Pi}{K})$. As a consequence, the projector $\Pi=K\varrho$ fulfills the Knill-Laflamme conditions in the limit $\ell\to\infty$.
Therefore, if every level of the hierarchy is feasible, there exists a state proportional to a projector $\Pi$ on $(\mathds{C}^{2})^{\otimes n}$ corresponding to a quantum code subspace with parameters $(\!(n,K,\delta)\!)_{2}$. This ends the proof. ∎
In the following, we show how an intermediate level of this hierarchy yields a Lovász bound [Section 5], whose symmetrization results a quantum Delsarte bound [Section 7]. Using the Terwilliger algebra of the Hamming association scheme, we also provide efficiently computable symmetry-reductions [Section 8.3]. Finally, we provide an alternative nonexistence proof of the $(\!(7,1,4)\!)_{2}$ code [Section 9.1].
5. A Lovász bound
We now present a relaxation of the hierarchy that leads to a bound involving the Lovász theta number [Section 3.5]. Consider an intermediate level that is indexed by state monomials of the form $\langle e_{i}\rangle e_{i}$ where $e_{i}$ correspond to letters from $\mathscr{E}_{n}$ [see Eq. (39)]. For simplicity, we will use elements from the $n$-qubit error basis $E_{i}\in\mathcal{E}_{n}$ to index the intermediate level instead of their corresponding letters $e_{i}\in\mathscr{E}_{n}$.
5.1. An intermediate relaxation level
We now present a fixed level of the hierarchy from Eq. (4.5) with different indexing than before. Similarly to $M^{(\ell)}$, consider a moment matrix $\Gamma$ indexed by the symbols ${\langle E^{\dagger}_{a}\rangle}E_{a}$ for all $E_{a}\in\mathcal{E}_{n}$, where ${\langle E_{a}\rangle}$ inherits the properties of an expectation $\operatorname{tr}(E_{a}\varrho)$ for some state $\varrho$,
(56) $\Gamma_{ab}={\langle E_{a}^{\dagger}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E_{b}^{\dagger}\rangle}\,.$
Such matrix has the form,
[…]
with $N=4^{n}-1$ and $\Gamma_{a0}=\Gamma_{aa}$, which approximates ${\langle E^{\dagger}_{a}\rangle}{\langle E_{a}\rangle}$. Now note that if $\Gamma$ is a feasible solution of a semidefinite program with objective function and constraints depending only on the real part of $\Gamma$, then $(\Gamma+\Gamma^{T})/2$ will also be feasible with the same objective value. Without loss of generality, we can therefore consider $\Gamma$ to be real, allowing us to impose the constraint
(59) $\Gamma_{ab}=0\quad\quad\text{if}\quad\quad E^{\dagger}_{a}E_{b}+E_{b}E^{\dagger}_{a}=0\,.$
This follows from the fact that when two tensor-products of Pauli matrices $E_{a}$ and $E_{b}$ anti-commute, the expectation value ${\langle E_{a}E_{b}^{\dagger}\rangle}$ over a quantum state $\varrho$ must be imaginary. Consequently, ${\langle E_{a}^{\dagger}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E_{b}^{\dagger}\rangle}$ is imaginary, and vanishes under the operation $\big{(}\Gamma+\Gamma^{T}\big{)}/2$.
5.2. Structure constraint
When $E_{a}$ and $E_{b}$ commute, the matrix $\Gamma$ satisfies an extra restriction because it approximates a multiplication of three terms: ${\langle E^{\dagger}_{a}\rangle}$, ${\langle E_{b}\rangle}$, and ${\langle E_{a}E^{\dagger}_{b}\rangle}$. For example, by identifying $E^{\dagger}_{a^{\prime}}=\omega_{a,b}E_{a}E^{\dagger}_{b}$ and $E_{b^{\prime}}=E^{\dagger}_{a}$, we obtain $\omega_{a,b}E_{a^{\prime}}E^{\dagger}_{b^{\prime}}=E_{b}$ which leads to
(60) ${\langle E^{\dagger}_{a}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E^{\dagger}_{b}\rangle}={\langle E^{\dagger}_{a^{\prime}}\rangle}{\langle E_{b^{\prime}}\rangle}{\langle E_{a^{\prime}}E^{\dagger}_{b^{\prime}}\rangle}\,.$
Therefore, we can write such constraint as
$\Gamma_{ab}=\Gamma_{a^{\prime}b^{\prime}}\,,\quad\quad\quad$ $\text{if}\quad\quad E_{a}E^{\dagger}_{b}=E^{\dagger}_{b}E_{a}\quad\text{and}\quad\,$
(61) $(E^{\dagger}_{a},E_{b},\omega_{ab}E_{a}E^{\dagger}_{b})\quad\text{a permutation of}\quad(E^{\dagger}_{a^{\prime}},E_{b^{\prime}},\omega_{a^{\prime}b^{\prime}}E_{a^{\prime}}E^{\dagger}_{b^{\prime}})\,,$
where $(\omega_{ab}E_{a}E^{\dagger}_{b})\in\mathcal{E}_{n}$ and $\omega_{ab}\in\{\pm 1\}$, since when $E_{a}$ and $E^{\dagger}_{b}$ commute $\omega_{ab}$ cannot be imaginary. Note that this already includes the constraint $\Gamma_{a0}=\Gamma_{aa}$.
5.3. Projector constraints
As was done for the matrix $M^{(\ell)}$ [see Eq. (54)], we can relax the two projector constraints $\operatorname{tr}(\Pi)=K$ and $\operatorname{tr}(\Pi^{2})=K$ through sums over entries of $\Gamma$,
(62) $\sum^{N}_{a=0}\Gamma_{aa}=\frac{2^{n}}{K},\quad\quad\sum^{N}_{b=0}\Gamma_{ab}=\frac{2^{n}}{K}\Gamma_{aa}\,.$
Summing over both $a$ and $b$ leads to $\sum^{N}_{a,\,b=0}\Gamma_{ab}=\frac{4^{n}}{K^{2}}$.
5.4. Knill-Laflamme conditions
Similar to Eq (43), we can approximate the Knill-Laflamme conditions $KB_{j}(\varrho)=A_{j}(\varrho)$ if $0\leq j<\delta$ using elements of $\Gamma$ together with $KB_{j}(\varrho)\geq A_{j}(\varrho)$ for any other $j$ [see Remark 5],
(63) $\frac{K}{2^{n}}\sum^{j}_{i=0}K_{j}(i;n)\!\!\sum_{\begin{subarray}{c}E_{a}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{a})=i\end{subarray}}\Gamma_{a0}\quad\geq\quad\sum_{\begin{subarray}{c}E_{b}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{b})=j\end{subarray}}\Gamma_{0b}\quad\text{with equality for}\quad 1<j<\delta\,.$
Note that we did not include the condition $KB_{0}(\varrho)=A_{0}(\varrho)$ because it is already approximated by $\sum^{N}_{a=0}\Gamma_{aa}=\frac{2^{n}}{K}$. This can be seen by Eq. (16) that gives $K_{0}(i,n)=1$ for all $i$. Therefore, $K2^{-n}\sum^{n}_{i=0}A_{i}(\varrho)=A_{0}(\varrho)$. Since $A_{0}(\varrho)=1$ and $\Gamma_{a0}=\Gamma_{aa}$, the trace of $\Gamma$ approximates $\sum^{n}_{j=0}A_{j}(\varrho)=2^{n}/K$.
5.5. Stabilizer codes
Let $S$ be the stabilizer group of a stabilizer code. Recall that $\Gamma$ approximates a matrix with entries $\operatorname{tr}(E_{a}^{\dagger}\varrho)\operatorname{tr}(E_{b}\varrho)\operatorname{tr}(E_{a}E^{\dagger}_{b}\varrho)$. If $\omega_{a}E_{a},\omega_{b}E_{b}\in S$ with $\omega_{a},\omega_{b}=\pm 1$, then $\Gamma_{ab}=1$ since all phases cancel. If $\omega_{a}E\notin S$ or $\omega_{b}E_{b}\notin S$, then $\Gamma_{ab}=0$. Therefore, $\Gamma_{ab}\in\{0,1\}$ for all $a,b\in{\boldsymbol{4}}^{n}$. For a positive semidefinite matrix, the determinant of every $2\times 2$ minor is non-negative, leading to $\Gamma_{ab}\leq\sqrt{\Gamma_{aa}\Gamma_{bb}}$. Due to $\Gamma_{ab}\in\{0,1\}$ this implies that $\Gamma_{ab}\leq\sqrt{\Gamma_{aa}}$. This leads to the following constraint only valid for stabilizer codes
(64) $0\leq\Gamma_{ab}\leq\Gamma_{aa}$
5.6. Pure codes
In the special case of pure codes, an extra constraint can be added. A code is said to be pure, if the Knill-Laflamme conditions stated in Eq. (8) are strengthened to
(65) $\Pi E^{\dagger}_{a}E_{b}\Pi=0\quad\quad\text{if}\quad\quad 0<\operatorname{wt}(E^{\dagger}_{a}E_{b})<\delta\,.$
As a consequence, we see that $\operatorname{tr}(E^{\dagger}_{a}E_{b}\Pi)=0$ holds for all $0<\operatorname{wt}(E^{\dagger}_{a}E_{b})<\delta$. Because $\Gamma_{ab}={\langle E^{\dagger}_{a}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E_{b}^{\dagger}\rangle}$, this yields the additional constraint,
(66) $\Gamma_{ab}=0\quad\quad\text{if}\quad\quad 0<\operatorname{wt}(E^{\dagger}_{a}E_{b})<\delta\,,\quad 0<\operatorname{wt}(E^{\dagger}_{a})<\delta\,,\quad\text{or}\quad 0<\operatorname{wt}(E_{b})<\delta\,.$
In particular, choosing $E_{0}=\mathds{1}$ implies that $\Gamma_{aa}=\Gamma_{a0}=0$ for all $E_{a}$ with $0<\operatorname{wt}(E_{a})<\delta$. Eq. (66) strengthens Eq. (63).
5.7. Main semidefinite program relaxation for qubits
Using the indexing of Eq. (56), we can write the following semidefinite programming relaxation to Eq. (2.1):
find $\Gamma\succeq 0\,,$
subject to $\Gamma_{00}=1\,,$
$\Gamma_{ab}=\Gamma_{a^{\prime}b^{\prime}}\,,\quad\text{if}\quad E_{a}E^{\dagger}_{b}=E^{\dagger}_{b}E_{a}\quad\text{and}\quad(E^{\dagger}_{a},E_{b},\omega_{ab}E_{a}E^{\dagger}_{b})\quad\,$
a permutation $\hskip 93.89418pt\text{a permutation of}\,\,(E^{\dagger}_{a^{\prime}},E_{b^{\prime}},\omega_{a^{\prime}b^{\prime}}E_{a^{\prime}}E^{\dagger}_{b^{\prime}})\,,$
$\Gamma_{ab}=0\quad\quad\quad\text{if}\quad E^{\dagger}_{a}E_{b}+E_{b}E^{\dagger}_{a}=0\,,$
$\sum^{N}_{a=0}\Gamma_{aa}=\frac{2^{n}}{K}\,,$
$\sum^{N}_{b=0}\Gamma_{ab}=\frac{2^{n}}{K}\Gamma_{aa}\,,$
(67) $\frac{K}{2^{n}}\sum^{n}_{i=0}K_{j}(i,n)\!\!\sum_{\begin{subarray}{c}E_{a}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{a})=i\end{subarray}}\Gamma_{a0}\quad\geq\quad\sum_{\begin{subarray}{c}E_{b}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{b})=j\end{subarray}}\Gamma_{0b}\quad\quad\text{with equality for}\quad 0<j<\delta\,.$
The fact that $\Gamma\succeq 0$ implies that its diagonal elements are non-negative. This recovers the non-negativity of $A_{j}(\varrho)=\sum_{E_{a}\in\mathcal{E}_{n},\,\operatorname{wt}(E_{a})=j}\langle E_{a}^{\dagger}\rangle\langle E_{a}\rangle\geq 0$ corresponding to $\sum_{E_{a}\in\mathcal{E}_{n},\,\operatorname{wt}(E_{a})=j}\Gamma_{aa}$. Eq. (5.7) incorporates the Knill-Laflamme and the projector conditions and the structure relations and therefore, it contains all constraints of the hierarchy in Eq. (4.5) for the indexing set ${\langle E^{\dagger}_{a}\rangle}E_{a}$.
We can further constrain the SDP above by including [see Section 3.4]:
  1. (1)
    The shadow inequalities $S_{j}(\rho)\geq 0$ [Eq. (18)],
    (68) $2^{-n}\sum^{n}_{i=0}(-1)^{i}K_{j}(i;n)\sum_{\begin{subarray}{c}E_{a}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{a})=j\end{subarray}}\Gamma_{aa}\quad\geq\quad 0$
    with equality for $K=1$ and $n-j$ odd [57, Theorem 12].
  2. (2)
    For stabilizer codes, we additionally add $0\leq\Gamma_{ab}\leq\Gamma_{aa}$ as shown in Eq. (64) and
    (69) $\sum_{j\geq 0}\sum_{E_{a}\in\mathcal{E}_{n},\,\operatorname{wt}(E_{a})=2j}\Gamma_{aa}=\begin{cases}2^{n-\log_{2}(K)-1}\quad\text{(Type I)}\,,\\ 2^{n-\log_{2}(K)}\quad\quad\!\text{(Type II)}\,.\end{cases}$
  3. (3)
    For pure codes, Eq. (66) shows that
    (70) $\Gamma_{ab}=0\quad\quad\text{if}\quad\quad 0<\operatorname{wt}(E^{\dagger}_{a}E_{b})<\delta\,,\quad 0<\operatorname{wt}(E^{\dagger}_{a})<\delta\,,\quad\text{or}\quad 0<\operatorname{wt}(E_{b})<\delta\,.$
A relaxation of the complete hierarchy in Theorem 6 is then:
Corollary 7.
If a $(\!(n,K,\delta)\!)_{2}$ quantum code exists, then the SDP in Eq. (5.7) is feasible.
5.8. Lovász bound for self-dual codes
A quantum code with $K=1$ is called self-dual [^3] This is an extension of the nomenclature for stabilizer codes, where $k=0$ ($K=1$) implies the stabilizer group is isomorphic to a self-dual additive classical code.. By definition, such code is pure and we can apply the conditions of the previous section.
Definition 8 (Confusability graph for self-dual codes).
Let $G$ be a graph whose vertices correspond to elements of the $n$-qubit Pauli basis $\mathcal{E}_{n}$. Let two vertices $a$ and $b$ be connected, denoted by $a\sim b$, if $E_{a},E_{b}\in\mathcal{E}_{n}$ satisfy $E_{a}E^{\dagger}_{b}+E_{b}E^{\dagger}_{a}=0$ or $0<\operatorname{wt}(E_{a}E^{\dagger}_{b})<\delta$; and let $a\sim a$ if $0<\operatorname{wt}(E_{a})<\delta$.
This simply incorporates the two constraints of Eq. (59) and (66). To see this, note that if a positive semidefinite matrix has an element in the diagonal equal to zero, this implies that all elements in the same row and column are zero, too. With this notation, we relax the SDP from Eq. (5.7) for self-dual codes to:
maximize $\sum^{N}_{a=1}\Gamma_{aa}\,,$
subject to $\Gamma_{00}=1\,,$
$\Gamma_{a0}=\Gamma_{aa}\,,$
$\Gamma\succeq 0\,,$
(71) $\Gamma_{ab}=0\quad\text{if}\quad a\sim b\,.$
Because of $a\sim a$ if and only if $a\sim 0$ the last constraint can be restricted to $a,b\neq 0$ without changing the objective value. Thus SDP (5.8) is nothing else than the semidefinite program for the Lovász theta number (3.5) of the confusability graph $G$ (Def. 8) with the vertex corresponding to the identity removed [^4] In fact, this is an instance of a Lovász theta number for a hypergraph [10].. To see this, note that $\Gamma$ is indexed by $0,\dots,N$, while the sum in Eq. (5.8) starts at $1$, corresponding to the sum over $N$ elements in Eq. (3.5). For a self-dual code to exist, the constraint $\sum^{N}_{a=0}\Gamma_{aa}=\frac{2^{n}}{K}$ in Eq. (5.7) must be met. Thus the objective value of Eq. (5.8) must be larger or equal than $12^{n}-$ 1. As a consequence, we have the following corollary:
Corollary 9.
If a pure $(\!(n,1,\delta)\!)_{2}$ code exists, then $2^{n}\leq 1+\vartheta(G^{\prime})$, where $G^{\prime}$ is the confusability graph $G$ (c.f. Def. 8) of a self-dual quantum code with the identity vertex removed.
Example 10.
For the confusability graph $G^{\prime}$ of the code $(\!(4,1,3)\!)_{2}$, the Lovász number is $\vartheta(G^{\prime})=7$. By Corollary 9, $1+\vartheta(G^{\prime})<2^{4}$ implies that a $(\!(4,1,3)\!)_{2}$ code does not exist. While this result was already known from the linear programming bounds that use the shadow inequalities [see Eq. (3.4)] 555 The weights $(1,0,0,12,3)$ are non-negative and invariant under the quantum MacWillams transformation., we think that Corollary 9 provides a conceptually simpler approach.
Note that $\Gamma$ from Eq. (5.8) has zeros in all rows and columns with $0<\operatorname{wt}(E_{a})<\delta$. The positivity of $\Gamma$ is then equivalent to the positivity of the same matrix without such rows and columns. This allows us to state the following.
Remark 11.
Let $G^{\prime\prime}$ be the subgraph of the confusability graph $G$ from Def. 8, whose vertices $\mathds{1}$ and $E_{a}\in\mathcal{E}_{n}$ with $\operatorname{wt}(E_{a})<\delta$ are removed. Then, $\vartheta(G^{\prime\prime})=\vartheta(G^{\prime})$.
5.9. Self-dual stabilizer codes
One can give the following argument on the appearance of the Lovász number in the case of self-dual stabilizer codes. A stabilizer code corresponds to an abelian stabilizer group $S$ for which $-\mathds{1}\not\in S$. Given $s_{a},s_{b}\in S$, there exist $s_{c}\in S$ such that $s_{c}=s_{a}s_{b}$. A pure stabilizer code additionally satisfies $\operatorname{wt}(s_{i})\geq\delta$ for all $s_{i}\in S\setminus\mathds{1}$.
Recall that $\vartheta(G)$ upper bounds the independence number $\alpha(G)$, i.e., the maximum size of a set of disconnected vertices of a graph $G$. Let $G^{\prime}$ be the confusability graph of a self-dual code with the identity vertex removed. Then two vertices $a,b\in V$ are disconnected if they satisfy
(72) $E_{a}^{\dagger}E_{b}+E_{b}E_{a}^{\dagger}=0\quad\text{and}\quad\operatorname{wt}(E_{a}^{\dagger}),\operatorname{wt}(E_{b}),\operatorname{wt}(E_{a}^{\dagger}E_{b})\geq\delta\,.$
By setting $E_{c}=\pm E_{a}^{\dagger}E_{b}$, the equation above shows that the vertices $a,b,c\in V$ do not share any edges. Denote by $H\subseteq V$ a subset of disconnected vertices of the graph $G^{\prime}$ with maximum size $\alpha(G^{\prime})$. Any $a,b\in H$ then satisfies
(73) $E_{a}^{\dagger}E_{b}+E_{b}E_{a}^{\dagger}=0,\quad\operatorname{wt}(E_{a}^{\dagger}),\operatorname{wt}(E_{b})\geq\delta\quad\text{and}\quad c\in H\quad\text{if}\quad E_{c}=\pm E_{a}^{\dagger}E_{b}.$
The above equation allows us to state the following: Given a self-dual stabilizer code $S$, one can construct a disconnected subset $H_{S}\subseteq V$ with size $12^{n}-$ 1 by identifying each vertex $a\in V$ with a Pauli error $E_{a}$ such that $E_{a}=\pm s_{a}\in S\setminus\mathds{1}$. Conversely, given $H$, one can construct a subset of $\mathcal{E}_{n}$ by identifying each vertex $a\in V$ with $E_{a}$. Using Eq. (73), it can be seen that all elements of this set has weight equal or larger than $\delta$ and up to real phases, it forms an abelian group if one includes $+\mathds{1}$. But this is nothing else than a stabilizer group. Note that this stabilizer group can have at most $12^{n}-$ 1 nontrivial elements since the maximum number of Pauli errors which commutes between each other is $2^{n}$. Therefore, a self-dual stabilizer code exists, if and only if, $\alpha(G^{\prime})=2^{n}-1$. When $\alpha(G^{\prime})<2^{n}-1$, it is not possible to construct a stabilizer group with $2^{n}$ elements and therefore, a self-dual stabilizer code does not exist. Everything also holds for the subgraph $G^{\prime\prime}$ from Remark 11.
6. Symmetry-reduction through the Terwilliger algebra
We use the technique developed by Schrijver [65] and Gijswijt et al. [23] for symmetry-reducing semidefinite programs that bound the parameters of binary and nonbinary codes respectively. This is done by averaging the semidefinite program in Eq. (5.7) through operations that leave the distances between triples of Pauli strings invariant. Then it can be block-diagonalized with the Terwilliger algebra of the Hamming association scheme. As we will eventually deal with an alphabet comprised of the four Pauli matrices $\mathds{1},X,Y,Z$, we are concerned with quaternary codes and thus follow the exposition of Ref. [23]. We first present the general case, before explaining the averaging and symmetry-reduction of the moment matrix $\Gamma$ from Eq. (56).
6.1. Automorphism groups
Let ${\boldsymbol{q}}=\{0,\dots,q-1\}$ be an alphabet and ${\boldsymbol{q}}^{n}$ be the set of strings of length $n$ formed from elements of $q$. Denote the Hamming distance between two elements $x,y\in{\boldsymbol{q}}^{n}$ by $d(x,y)$ the number of positions in which $x$ and $y$ differ. Then consider the set of automorphism $\operatorname{Aut}(q,n)$ of ${\boldsymbol{q}}^{n}$ that preserve the Hamming distance. It is known that $\operatorname{Aut}(q,n)$ consists of the set of permutations that permute the $n$ coordinates and that independently permute the alphabet at each coordinate [23]. Let $\operatorname{Aut}_{0}(q,n)$ be the subset of $\operatorname{Aut}(q,n)$ that leaves the all-zero string invariant. This consists of the set of permutations that permute the $n$ coordinates and independently permute the $q-1$ non-zero elements at each coordinate [23]. The group $\operatorname{Aut}_{0}(q,n)$ is isomorphic to the wreath product $S_{q-1}\wr S_{n}$.
The equivalence class of a classical code is generated by $\operatorname{Aut}(q,n)$. In turn, the equivalence class of a quantum code under local Clifford operations is generated by $\operatorname{Aut}_{0}(q,n)$. The reason is that the projector onto the code space has to remain positive-semidefinite under the equivalence operation, and consequentially the identity matrix has to be a fixed point.
Let $R$ be a real matrix indexed by $x,y\in\boldsymbol{q}^{n}$. Given an automorphism $\pi\in\operatorname{Aut}(q,n)$, let it act on $R$ by permuting its rows and columns accordingly,
(74) $\pi(R)_{xy}=R_{\pi(x)\pi(y)}\,.$
Note that if $R$ is positive semidefinite, then $\pi(R)$ is also positive semidefinite for any $\pi$. Matrices invariant under the action of $\operatorname{Aut}_{0}(q,n)$ are in the Terwilliger algebra, while the commutative subalgebra invariant under $\operatorname{Aut}(q,n)$ is the Bose-Mesner algebra [23].
6.2. Nonbinary Terwilliger algebra
Let $M_{i,j}^{t,p}$ a ${\boldsymbol{q}}^{n}\times{\boldsymbol{q}}^{n}$ matrix indexed by ${\boldsymbol{q}}^{n}$ and defined by
(75) $(M_{i,j}^{t,p})_{xy}=\begin{cases}1&\text{ if }\quad|s(x)|=i\,,\quad|s(x)|=j\,,\\ &\hskip 28.45274pt\lvert s(x)\cap s(y)\rvert=t\,,\quad\lvert\{m\,:\,x_{m}=y_{m}\neq 0\}\rvert=p\,,\vspace{0.2cm}\\ 0&\text{ otherwise}\,,\end{cases}$
where $s(x)=\{m\,|\,x_{m}\neq 0\}$ is the support of $x$; $|s(x)|$ is its size (that is the number of non-zero coordinates of $x$); and $x_{m}$ is the element at coordinate $m$ of $x$.
The range of integers, outside of which $M_{i,j}^{t,p}$ necessarily vanishes, is
(76) $\mathcal{I}(q,n)=\{(i,j,t,p)\,:\,0\leq p\leq t\leq i,j\quad\text{and}\quad i+j\leq t+n\,\}\,.$
Ref. [23, Proposition 9] shows that $M_{i,j}^{t,p}$ are pairwise orthogonal and $\left\langle M_{i,j}^{t,p},M_{i,j}^{t,p}\right\rangle=\gamma_{i,j}^{t,p}$, where
(77) $\gamma_{i,j}^{t,p}$ $=(q-1)^{i+j-t}(q-2)^{t-p}\binom{n}{p,t-p,i-t,j-t}\,,$
and
(78) $\binom{n}{a_{1},\dots,a_{r}}=\frac{n!}{a_{1}!\dots a_{r}!(n-\sum^{r}_{\ell=1}a_{\ell})!}\,$
is a multinomial coefficient. The set of $M_{i,j}^{t,p}$ forms a basis for the Terwilliger algebra $\mathcal{A}_{q,\,n}$ of the $q$-ary Hamming cube. The Terwilliger algebra can be block-diagonalized [23], allowing us to check the positivity of the blocks of an element in $\mathcal{A}_{q,\,n}$.
Remark 12.
Let us give some motivation as to how the indexing with $(i,j,t,p)$ in Eq. (75) will be used later. Consider the three terms $E^{\dagger}$, $F$, and $EF^{\dagger}$ that appear in a single entry of the matrix $\Gamma$ [Eq. (56)]. Now suppose that $\operatorname{wt}(E)=i$ and $\operatorname{wt}(F)=j$. Then, set $t=|s(E)\cap s(F)|$ and $p=|\{m\,:\,(E)_{m}=(F)_{m}\neq\mathds{1}\}$ the number of tensor factors (coordinates) where both $E$ and $F$ have identical nontrivial Pauli elements. Then the overlap between the supports of $E$ and $F$ is $t$, and the number of places a nontrivial Pauli in $E$ cancels a nontrivial Pauli in $F$ is $p$. The quantity $t-p$ is then the number of places that a nontrivial Pauli in $E$ does not cancel a nontrivial Pauli in $F$. Note that if $E$ and $F$ commutes then $t-p$ is even and if $E$ and $F$ anticommutes then $t-p$ is odd. Finally, $\operatorname{wt}(EF^{\dagger})=i+j-t-p$.
6.3. Averaging over codewords
We now consider different averaging methods, whose results lie in the Terwilliger algebra. Recall that an automorphism $\pi\in\operatorname{Aut}(q,n)$ acts on a matrix $R$ by permuting its rows and columns, $\pi(R)_{xy}=R_{\pi(x)\pi(y)}$. Now let $\sigma_{v}$ be a distance-preserving automorphism that additionally satisfies $\sigma_{v}(v)=0$. With this, define $R^{v}$ as
(79) $R^{v}=\frac{1}{\lvert\operatorname{Aut}_{0}(q,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(q,n)}\pi\left(\sigma_{v}\!\left(R\right)\right)\,.$
The effect of Eq. (79) is to average $R$ over all distance-preserving automorphisms that map codewords to the zero string.
Let $C$ be a code with elements from ${\boldsymbol{q}}^{n}$. Then one can additionally average over the set of matrices $R^{v}$ where $v$ is a codeword:
(80) $\tilde{R}(C)=\frac{1}{\left\lvert C\right\rvert}\sum\limits_{v\in C}R^{v}.$
The effect of Eq. (80) is to average $R$ over all distance-preserving automorphism that map codewords to the zero string. Note that both $R^{v}$ and $\tilde{R}$ are invariant under $\operatorname{Aut}_{0}(q,n)$.
In particular, this means we can express $\tilde{R}$ in a basis of the Terwilliger algebra $\mathcal{A}_{q,n}$,
(81) $\tilde{R}=\sum\limits_{\left(i,j,t,p\right)\in\mathcal{I}(q,n)}x_{i,j}^{t,p}M_{i,j}^{t,p}\,,\quad\quad x_{i,j}^{t,p}\in\mathds{R}\,.$
We now show how to write $x_{i,j}^{t,p}$ in terms of elements from the matrix $R$. Define $\lambda^{t,p}_{i,j}=\langle\,\tilde{R},M_{i,j}^{t,p}\,\rangle$ and note that using Eq. (81), we get that $\lambda^{t,p}_{i,j}=\gamma^{t,p}_{i,j}x^{t,p}_{i,j}$. We then develop $\lambda^{t,p}_{i,j}$ using the definition of $\tilde{R}$ in Eq. (80),
$\lambda^{t,p}_{i,j}$ $=\frac{1}{\left\lvert C\right\rvert}\frac{1}{\lvert\operatorname{Aut}_{0}(q,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(q,n)}\sum\limits_{v\in C}\Big{\langle}\pi\left(\sigma_{v}\!\left(R\right)\right),M_{i,j}^{t,p}\Big{\rangle}$
$=\frac{1}{\left\lvert C\right\rvert}\left\langle\sum\limits_{v\in C}\sigma_{v}\!\left(R\right),\frac{1}{\lvert\operatorname{Aut}_{0}(q,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(q,n)}\pi^{-1}\left(M_{i,j}^{t,p}\right)\right\rangle$
(82) $=\frac{1}{\left\lvert C\right\rvert}\left\langle\sum\limits_{v\in C}\sigma_{v}\!\left(R\right),M_{i,j}^{t,p}\right\rangle$
Here we used the fact that $M^{t,p}_{i,j}$ is invariant under the action of elements from $\operatorname{Aut}_{0}(q,n)$. From the definition of $M^{t,p}_{i,j}$ [see Eq. (75)], one obtains
(83) $\lambda^{t,p}_{i,j}=\frac{1}{\left\lvert C\right\rvert}\!\!\!\!\sum\limits_{\begin{subarray}{c}x,y\in{\boldsymbol{q}}^{n}\\ |s(x)|=i\\ |s(y)|=j\\ \lvert s(x)\cap s(y)\rvert=t\\ \lvert\{m\mid x_{m}=y_{m}\neq 0\}\rvert=p\\ \end{subarray}}\!\!\!\!\sum\limits_{v\in C}\sigma_{v}(R)_{xy}\,.$
since the only elements $\sigma_{v}(R)_{xb}(M_{i,j}^{t,p})_{xb}$ different to zero are the ones satisfying $|s(x)|=i$, $|s(x)|=j$, $\lvert s(x)\cap s(b)\rvert=t$ and |$\lvert\{m\mid x_{m}=b_{m}\neq 0\}\rvert=p$.
Remark 13.
Ref. [23, Proposition 9] defines $\langle\,\tilde{R},M_{i,j}^{t,p}\,\rangle=\frac{1}{|C|}\lambda^{t,p}_{i,j}$. However, we decide to define $\langle\,\tilde{R},M_{i,j}^{t,p}\,\rangle=\lambda^{t,p}_{i,j}$ since in the quantum case, $\tilde{\Gamma}$ is averaged over $\operatorname{Aut}_{0}(q,n)$ instead of averaging over $\operatorname{Aut}_{0}(q,n)$ and codewords (see Proposition  17).
The matrix $\tilde{R}$ can be block-diagonalized. For this define the following coefficients [23, Eq. (20) and (27)]:
$\alpha(i,j,t,p,a,k)$ $=\beta_{i-a,j-a,k-a}^{n-a,t-a}\left(q-1\right)^{\frac{1}{2}\left(i+j\right)-t}\sum\limits_{g=0}^{p}\left(-1\right)^{a-g}\binom{a}{g}\binom{t-a}{p-g}\left(q-2\right)^{t-a-p+g},$
(84) $\beta_{i,j,k}^{m,t}$ $=\sum\limits_{u=0}^{m}\left(-1\right)^{t-u}\binom{u}{t}\binom{m-2k}{m-k-u}\binom{m-k-u}{i-u}\binom{m-k-u}{j-u}\,.$
It can be shown that $\tilde{R}$ is positive semidefinite, if and only if
(85) $\bigoplus_{\begin{subarray}{c}a,k\in\mathds{N}_{0}\\ 0\leq a\leq k\leq n+a-k\end{subarray}}\left(\sum\limits_{\begin{subarray}{c}t,p\in\mathds{N}_{0}\\ 0\leq p\leq t\leq i,j\\ i+j\leq t+n\end{subarray}}\alpha(i,j,t,p,a,k)x_{i,j}^{t,p}\right)_{i,j=k}^{n+a-k}\,.$
is positive semidefinite [23, Eq. (43)].
6.4. Averaging over the complement of a code
Similarly, one can average over the elements that are not in the code and obtain
(86) $\tilde{R}^{c}(C)=\frac{1}{q^{n}-\left\lvert C\right\rvert}\sum\limits_{v\in{\boldsymbol{q}}^{n}\setminus\text{C}}R^{v}.$
The average over all $\sigma\in\operatorname{Aut}(q,n)$ can then be expressed as
(87) $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu=\frac{1}{q^{n}}\sum_{v\in{\boldsymbol{q}}^{n}}R^{v}=\frac{1}{q^{n}}\left(|C|\tilde{R}+(q^{n}-|C|)\tilde{R}^{c}\right)\,.$
Since is invariant under all permutations $\sigma\in\operatorname{Aut}(q,n)$, the matrix $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is an element of the Bose–Mesner algebra [23, Eq. (40)]. Therefore, it can be written as
(88) $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu=\sum^{n}_{k=0}y_{k}N_{k}\,,$
where $y_{k}\in\mathds{R}$ and $\{N_{k}\}_{k=0}^{n}$ is a basis for the Bose-Mesner algebra given by
(89) $(N_{k})_{xy}=\begin{cases}1\quad\text{if }\,\,d(x,y)=k\,,\\ 0\quad\text{otherwise}\,.\end{cases}$
Here $d(x,y)$ is the Hamming distance between $x$ and $y$. Note that that the Terwilliger algebra $\mathcal{A}_{q,n}$ contains the Bose-Mesner algebra as a commutative subalgebra, with
(90) $N_{k}=\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(q,n)\\ i+j-t-p=k\end{subarray}}M^{t,p}_{i,j}\,.$
Suppose that $\tilde{R}^{c}_{x0}=0$ for all $x$. This happens for example when $R=\chi\chi^{T}$ where $\chi$ is the indexing vector of a classical code. Ref. [23, Proposition 8] shows that
$\tilde{R}^{c}$ $=\frac{1}{q^{n}-|C|}(q^{n}\bar{R}-|C|\tilde{R})$
(91) $=\frac{|C|}{q^{n}-|C|}\sum_{(i,j,t,p)\in\mathcal{I}(q,n)}(x^{0,0}_{i+j-t-p,0}-x^{t,p}_{i,j})M^{t,p}_{i,j}\,.$
Then, in analogy to Eq. (85), the $\tilde{R}^{c}\succeq 0$ is equivalent to
(92) $\bigoplus_{\begin{subarray}{c}a,k\in\mathds{N}_{0}\\ 0\leq a\leq k\leq n+a-k\end{subarray}}\left(\sum\limits_{\begin{subarray}{c}t,p\in\mathds{N}_{0}\\ 0\leq p\leq t\leq i,j\\ i+j\leq t+n\end{subarray}}\alpha(i,j,t,p,a,k)(x_{i+j-t-p,0}^{0,0}-x_{i,j}^{t,p})\right)_{i,j=k}^{n+a-k}\succeq 0\,.$
When $R=\chi\chi^{T}$, it follows from the diagonalization of $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ that the $B_{i}$ enumerators are non-negative. This, together with the non-negativity of the $A_{i}$ enumerator, gives the Delsarte bound [see Eq. (3.5)]  [13, 23].
6.5. The symmetry reduction of $\Gamma$
A natural mapping identifying $\mathds{1},Z,X,Y$ with $0,1,\omega,\omega^{2}\in\operatorname{\mathrm{GF}}(4)$ respectively maps, up to a phase, the multiplication of Pauli matrices to addition over $\operatorname{\mathrm{GF}}(4)$ [9]. That is, $E_{x+y}=i^{b}E_{x}E_{y}$ for some $b=0,1,2,3$. Therefore, we map $\mathcal{E}_{n}$ to $\operatorname{\mathrm{GF}}(4)^{n}$. In particular, $E_{0}=\mathds{1}^{\otimes n}$, where we write $E_{0}$ for $E_{0^{n}}$. Indexing the matrix $\Gamma$ from Eq. (56) with all elements from $\operatorname{\mathrm{GF}}(4)^{n}$ and eventually ${{\boldsymbol{q}}}^{n}$ with $q=4$, our aim is now to follow the averaging methods from Sections 6.3 and 6.4.
For any $v\in\operatorname{\mathrm{GF}}(4)^{n}$, define the automorphism $\sigma_{v}\in\operatorname{Aut}(\mathcal{E}_{n})$ acting on $\Gamma$ as
$\sigma_{v}\!\left(\Gamma\right)_{x,y}$ $=\Gamma_{{x-v},{y-v}}$
$={\langle E_{x-v}^{\dagger}\rangle}{\langle E_{y-v}\rangle}{\langle E_{x-v}E^{\dagger}_{y-v}\rangle}$
$={\langle(E_{x}E_{v}^{\dagger})^{\dagger}\rangle}{\langle E_{y}E_{v}^{\dagger}\rangle}{\langle E_{x}E_{v}^{\dagger}E_{v}E_{y}^{\dagger}\rangle}$
(93) $={\langle E_{v}E_{x}^{\dagger}\rangle}{\langle E_{y}E_{v}^{\dagger}\rangle}{\langle E_{x}E_{y}^{\dagger}\rangle}\,.$
This maps the row and column indexed by $v$ to the one indexed by $0^{n}$ respectively. We can then average $\Gamma$ over a subset of $\operatorname{Aut}(\operatorname{\mathrm{GF}}(4),n)$, such that it can be expanded in the Terwilliger algebra. Note that averaging over distance-preserving automorphisms of $\operatorname{\mathrm{GF}}(4)^{n}$ has the same effect as over distance-preserving automorphisms of ${\boldsymbol{q}}^{n}$ with $q=4$, because the Hamming distance between code words is not affected by any group structure of the underlying alphabet. Thus with some abuse of notation, we define in analogy to Eq. (79) the matrix $\Gamma^{v}$ as
(94) $\Gamma^{v}=\frac{1}{\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi\left(\sigma_{v}\!\left(\Gamma\right)\right)\,.$
Similar to Eq. (80), define
(95) $\tilde{\Gamma}(Q)=\frac{1}{|Q|}\sum\limits_{v\in Q}\Gamma^{v}\,,$
where $Q$ is a subset from ${\boldsymbol{4}}^{n}$. The effect of Eq. (95) is to average $\Gamma$ over all distance-preserving automorphisms that map an element of $Q$ to $0^{n}$. Likewise, in analogy to Eq. (87) define the averaging over all elements from $\operatorname{Aut}(\mathcal{E}_{n})$ as
(96) $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu=$ $\frac{1}{4^{n}}\sum_{v\in{\boldsymbol{4}}^{n}}\Gamma^{v}\,.$
We now consider averaging over different type of subsets $Q$. We first take $Q={\boldsymbol{4}}^{n}$. This is equivalent to averaging $\Gamma$ over all elements from $\operatorname{Aut}(4,n)$, which will lead to the quantum Delsarte bound.
7. Quantum Delsarte bound
For classical codes, the symmetrization over $\operatorname{Aut}(q,n)$ of the semidefinite program SDP1 for the Lovász theta number [Eq. (3.5)] along with the non-negativity of the matrix entries as extra constraint leads to the Delsarte bound [4]. The matrix used for the Lovász theta bound is defined as $R=\chi\chi^{T}$, where $\chi$ is the indexing vector of a classical code. Our aim is to recover a similar result for the Lovász number for self-dual quantum codes. For simplicity in the presentation, we here use the semidefinite program SDP2 [see Eq. (3.5)]. A key difference to the classical case is that the entries of the semidefinite variable from the quantum Lovász theta bound [see Corollary 9] can be negative. Due to this fact, we require a different extra constraint to the quantum Lovász theta bound to recover the quantum Delsarte bound [see Eq. (3.4)] after the symmetrization.
Suppose $\Gamma$ arises from a state. Our aim is to show that the $A_{j}(\rho)$ enumerators emerge as elements from $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ expressed in the Bose-Mesner algebra basis $\{N_{k}\}_{k=0}^{n}$ [see Eq. (90)]. Diagonalizing $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$, one then obtain by means of the quantum MacWilliams identity, the dual enumerator $B_{j}(\rho)$. As in Eq. (88), the matrix $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ is part of the Bose-Mesner algebra and can be expanded in terms of the basis $\{M^{t,p}_{i,j}\}_{(i,j,t,p)\in\mathcal{I}(4,n)}$ of the Terwilliger algebra,
(97) $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu=\sum^{n}_{k=0}y_{k}N_{k}=\quad\sum^{n}_{k=0}y_{k}\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ k=i+j-t-p\end{subarray}}M^{t,p}_{i,j}\quad=\sum_{(i,j,t,p)\in\mathcal{I}(4,n)}y_{i+j-t-p}M^{t,p}_{i,j}\,,$
where we have used the decomposition in Eq. (90). Given $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$, the goal is now to obtain $y_{k}$ in terms of elements from $\Gamma$.
Recall that $M^{t,p}_{i,j}$ is an orthogonal basis, that is $\left\langle M_{i,j}^{t,p},M_{i,j}^{t,p}\right\rangle=\gamma_{i,j}^{t,p}$ with $\gamma_{i,j}^{t,p}$ given by Eq. (77). Since all $M_{i,j}^{t,p}$ with $k=i+j-t-p$ have the same coefficient $y_{k}$, it is enough to compute the inner product with a representative, e.g., $M_{k,0}^{0,0}$,
[…]
Deriving ${\langle\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu,M^{0,0}_{k,0}\rangle}$ is thus enough to determine $y_{k}$. Following the procedure of Eq. (6.3) to (83) and using the definition of $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ in Eq. (96), one obtains
${\langle\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu,M^{0,0}_{k,0}\rangle}$ $=\frac{1}{4^{n}\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(4,n)}\sum\limits_{v\in{\boldsymbol{4}}^{n}}\left\langle\pi\left(\sigma_{v}\!\left(\Gamma\right)\right),M_{k,0}^{0,0}\right\rangle$
$=\left\langle\frac{1}{4^{n}}\sum\limits_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}\!\left(\Gamma\right),\frac{1}{\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi^{-1}\left(M_{k,0}^{0,0}\right)\right\rangle$
$=\left\langle\frac{1}{4^{n}}\sum\limits_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}\!\left(\Gamma\right),M_{k,0}^{0,0}\right\rangle$
(99) $=\frac{1}{4^{n}}\sum\limits_{\begin{subarray}{c}x\in{\boldsymbol{4}}^{n}\\ \operatorname{wt}(x)=k\end{subarray}}\sum\limits_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}(\Gamma)_{x0}\,.$
We now show that $y_{k}$ is proportional to the $A_{k}(\rho)$ enumerator. Recall the action of $\sigma_{v}$ on $\Gamma$ in Eq. (6.5). To this end, let us expand for a quantum code $\varrho=\Pi/K$ the last sum in Eq. (7),
$\sum_{{v}\in{\boldsymbol{4}}^{n}}\sigma_{v}\!\left(\Gamma\right)_{x0}$ $=\sum_{E_{v}\in\mathcal{E}_{n}}\Big{(}\operatorname{tr}(E_{v}E^{\dagger}_{x}\varrho)\operatorname{tr}(E^{\dagger}_{v}\varrho)\Big{)}\operatorname{tr}(E_{x}\varrho)$
$=\operatorname{tr}\Big{(}\sum_{E_{v}\in\mathcal{E}_{n}}(E_{v}E^{\dagger}_{x}\otimes E^{\dagger}_{v})(\varrho\otimes\varrho)\Big{)}\operatorname{tr}(E_{x}\varrho)$
$=2^{n}\operatorname{tr}\Big{(}(1,2)(E^{\dagger}_{x}\varrho\otimes\varrho)\Big{)}\operatorname{tr}(E_{x}\varrho)$
$=2^{n}\operatorname{tr}(E^{\dagger}_{x}\varrho^{2})\operatorname{tr}(E_{x}\varrho)$
(100) $=\frac{2^{n}}{K}\operatorname{tr}(E^{\dagger}_{x}\varrho)\operatorname{tr}(E_{x}\varrho)=\frac{2^{n}}{K}\Gamma_{x0}\,.$
Here we have used that $\varrho^{2}=\varrho/K$ and $\operatorname{tr}((1,2)\varrho^{\otimes 2})=\varrho^{2}$ where the swap operator is $(1,2)=2^{-n}\sum_{E_{v}\in\mathcal{E}_{n}}E_{v}\otimes E^{\dagger}_{v}$. Note that Eq. (7) can also be seen from combining the constraints Eq. (5.2) and Eq. (62) in the main semidefinite programming relaxation.
With this we can write Eq. (98) as
$y_{k}$ $=\frac{1}{4^{n}\gamma^{0,0}_{k,0}}{\langle\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu,M^{0,0}_{k,0}\rangle}$
$=\frac{1}{4^{n}\gamma^{0,0}_{k,0}}\sum\limits_{\begin{subarray}{c}x\in{\boldsymbol{4}}^{n}\\ \operatorname{wt}(x)=k\end{subarray}}\sum\limits_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}(\Gamma)_{x0}$
$=\frac{1}{4^{n}\gamma^{0,0}_{k,0}}\frac{2^{n}}{K}\sum_{\begin{subarray}{c}E_{x}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{x})=k\end{subarray}}\Gamma_{x0}$
(101) $=(2^{n}K\gamma^{0,0}_{k,0})^{-1}A_{k}(\varrho)\,.$
since $\sum_{x\in{\boldsymbol{4}}^{n},\,|s(x)|=k}\Gamma_{x0}=A_{k}(\varrho)$ and we used Eq. (7).
We now diagonalize $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$. Ref [13] shows that the eigenvalues from $N_{k}$ are the Krawtchouk polynomials $K_{k}(i)$ [see Eq. (16)]. Since $\{N_{k}\}^{n}_{k=0}$ is a commuting basis it can be simultaneous diagonalized, thus the eigenvalues of $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ are $\sum^{n}_{k=0}y_{k}K_{k}(i)$. The quaternary Krawtchouk polynomials satisfy the following relation [44, Theorem 17],
(102) $3^{i}\binom{n}{i}K_{k}(i)=3^{k}\binom{n}{k}K_{i}(k)\,.$
With Eq. (77), one can express this as $\gamma^{0,0}_{i,0}K_{k}(i)=\gamma^{0,0}_{k,0}K_{i}(k)$. We then develop the eigenvalues of $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ as follows
$\sum^{n}_{k=0}y_{k}K_{k}(i)$ $=\frac{1}{\gamma^{0,0}_{i,0}}\sum^{n}_{k=0}y_{k}\gamma^{0,0}_{k,0}K_{i}(k)$
(103) $=(2^{n}K\gamma^{0,0}_{k,0})^{-1}\sum^{n}_{k=0}K_{i}(k)A_{k}(\varrho)\,,$
where we have used Eq. (7).
Recall that $\sum_{x\in{\boldsymbol{4}}^{n},\,|s(x)|=k}\Gamma_{x0}=A_{k}(\varrho)$ for a state $\varrho=\pi/K$. By the quantum MacWilliams identity in Eq. (15), it can be seen that Eq. (7) is proportional to the dual enumerator $B_{i}(\varrho)$. Thus both the $A_{j}$ and $B_{j}$ enumerators from averaging $\Gamma$ over $\operatorname{Aut}(q,n)$, and diagonalizing the resulting $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ in the basis of the Bose-Mesner algebra $\{N_{k}\}_{k=0}^{n}$.
However, we want to point out that Eq. (7) assumes that $\Gamma$ has been constructed from a state $\varrho=\Pi/K$. Thus this is an extra constraint that has to be imposed on the Lovász bound in order to obtain the quantum Delsarte bound:
Theorem 14.
For self-dual quantum codes, the averaging of the Lovász bound from Corollary 9 over $\operatorname{Aut}(q,n)$ with the condition
(104) $\sum_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}(\Gamma)_{x0}=2^{n}\Gamma_{x0}\,,$
yields the quantum Delsarte bound [Eq. (3.4)].
Proof.
Note that if $\Gamma$ is constructed from a code $\varrho$, then Eq. (7) shows that $\sum_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}(\Gamma)_{x0}=2^{n}\Gamma_{x0}$ for all $x\in{\boldsymbol{4}}^{n}$. By including $\sum_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}(\Gamma)_{x0}=2^{n}\Gamma_{x0}$ in the program for the Lovász theta number for self-dual quantum codes from Corollary 9, we obtain,
max $\sum_{x\in\mathcal{E}_{n}\backslash\mathds{1}}\Gamma_{xx}$
subject to $\Gamma\succeq 0,\quad\Gamma_{00}=1\,,\quad\Gamma_{x0}=\Gamma_{xx}\,,$
$\Gamma_{xy}=0\quad\text{if}\quad x\sim y\,,$
(105) $\sum_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}(\Gamma)_{x0}=2^{n}\Gamma_{x0}\,.$
Note that $\Gamma_{xx}=\Gamma_{x0}\geq 0$ since $\Gamma\succeq 0$. Now average $\Gamma$ from Eq. (7) over $\operatorname{Aut}(4,n)$ to obtain $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu=\sum^{n}_{k=0}y_{k}N_{k}$. Using Eq. (98), Eq. (7), and the constraint Eq. (104) write $y_{k}$ as
$y_{k}$ $=\frac{1}{\gamma^{0,0}_{k,0}}{\langle\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu,M^{0,0}_{k,0}\rangle}$
$=\frac{1}{4^{n}\gamma^{0,0}_{k,0}}\sum\limits_{\begin{subarray}{c}x\in{\boldsymbol{4}}^{n}\\ \operatorname{wt}(x)=k\end{subarray}}\sum\limits_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}(\Gamma)_{x0}$
(106) $=(2^{n}\gamma^{0,0}_{k,0})^{-1}\sum\limits_{\begin{subarray}{c}x\in{\boldsymbol{4}}^{n}\\ \operatorname{wt}(x)=k\end{subarray}}\Gamma_{x0}\,.$
Define $a_{j}=\sum_{x\in{\boldsymbol{4}}^{n}\,,\operatorname{wt}(x)=j}\Gamma_{x0}$, which corresponds to the $A_{j}$ enumerator. Note that $\bar{\Gamma}$ only depends on $a_{j}$. The $\{a_{j}\}^{n}_{j=0}$ satisfy:
  1. (1)
    $a_{0}=\Gamma_{00}=1$,
  2. (2)
    $a_{j}=\sum_{x\in{\boldsymbol{4}}^{n}\,,\operatorname{wt}(x)=j}\Gamma_{x0}\geq 0$, since $\Gamma_{x0}\geq 0$  ,
  3. (3)
    $a_{j}=0$ if $0\leq j<\delta$, since $\Gamma_{x0}=0$ if $0<\operatorname{wt}(x)<\delta$.
  4. (4)
    $\sum^{n}_{j=0}a_{j}=\sum_{x\in{\boldsymbol{4}}^{n}}\Gamma_{xx}=\sum_{x\in{\boldsymbol{4}}^{n}}\Gamma_{x0}$.
It remains to use the fact that $\Gamma\succeq 0$. Recall that $\Gamma\succeq 0$ implies $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\succeq 0$. By Eq. (7), $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ has as eigenvalues,
(107) $(2^{n}\gamma^{0,0}_{j,0})^{-1}\sum^{n}_{j=0}K_{i}(j)A_{j}\,.$
Therefore, $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\succeq 0$ if and only if $\sum^{n}_{j=0}K_{i}(j)A_{j}\geq 0$ for all $0\leq i\leq n$. The combination with (1) - (4) above yields the quantum Delsarte bound for self-dual codes [Eq. (3.4)],
$\eta=\text{max}\quad$ $\sum^{n}_{j=0}a_{j}$
subject to $a_{0}=1\,,$
$a_{j}\geq 0\quad\text{with equality for }\,\,1<j<\delta\,,$
(108) $\sum^{n}_{i=0}K_{j}(i)a_{i}\geq 0\quad\text{for}\quad 0\leq j\leq n\,.$
If $\eta<2^{n}$ then a code with parameters $(\!(n,1,\delta)\!)_{2}$ does not exist. This ends the proof. ∎
Remark 15.
Naturally, this bound can be made stronger by formulating it as a feasibility program that includes the constraint $\sum_{j=0}^{n}a_{j}=2^{n}$. We use this for excluding the $(\!(7,1,4)\!)_{2}$ code in Section 9.1.
8. Symmetry-reduced semidefinite programming bounds
We now aim to symmetry-reduce $\Gamma$ while retaining as many constraints from the original program [see Eq. (5.7)] as possible. This is achieved by averaging $\Gamma$ over $\operatorname{Aut}_{0}(4,n)$, which corresponds to Eq. (95) with $Q=\{0^{n}\}$. We therefore define
(109) $\tilde{\Gamma}=\Gamma(0^{n})=\frac{1}{\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi\left(\Gamma\right)\,.$
In analogy to Eq. (87), we define $\tilde{\Gamma}^{c}$ as
(110) $\tilde{\Gamma}^{c}=\frac{1}{4^{n}-2^{n}/K}\left(4^{n}\bar{\Gamma}-\frac{2^{n}}{K}\tilde{\Gamma}\right)\,.$
where $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ is stated in Eq. (96). Note that we choose $2^{n}/K$ in the normalization of $\tilde{\Gamma}$, which in case of stabilizer codes this corresponds to the number of stabilizers. In general however, the matrix $\tilde{\Gamma}^{c}$ does not necessarily come from averaging $\Gamma$ over the codewords of a code, i.e., $\tilde{\Gamma}^{c}\neq\frac{1}{4^{n}-\left\lvert Q\right\rvert}\sum_{v\notin Q}\Gamma^{v}$ for some $Q\subseteq{\boldsymbol{4}}^{n}$. It is thus not immediate that $\tilde{\Gamma}^{c}$ is positive semidefinite. We now show that $\tilde{\Gamma}^{c}\succeq 0$ for all quantum codes.
Proposition 16.
For every quantum code $\tilde{\Gamma}\succeq 0$ and $\tilde{\Gamma}^{c}\succeq 0$.
Proof.
The matrix $\tilde{\Gamma}$ is positive semidefinite by construction in Eq. (109), being a sum over positive semidefinite matrices.
We now consider $\tilde{\Gamma}^{c}$. We first show that a matrix with the entries
(111) $L_{xy}=\big{[}\operatorname{tr}(E_{y}E^{\dagger}_{x}\varrho)-\operatorname{tr}(E^{\dagger}_{x}\varrho)\operatorname{tr}(E_{y}\varrho)\big{]}\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho)\,,$
is positive semidefinite. To see this, write $L=M\circ G$ where $\circ$ is the Schur product, that is $L_{xy}=M_{xy}G_{xy}$, and where
$M_{xy}$ $=\operatorname{tr}(E_{y}E^{\dagger}_{x}\varrho)-\operatorname{tr}(E^{\dagger}_{x}\varrho)\operatorname{tr}(E_{y}\varrho)\,,$
(112) $G_{xy}$ $=\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho)\,.$
The Schur product theorem states that if $A\succeq 0$ and $B\succeq 0$, then $A\circ B\succeq 0$.
As it has the form of a moment matrix it is clear that $G\succeq 0$. Let us show that also $M\succeq 0$. Define two matrices $M^{\pm}$ with entries
(113) $M^{\pm}_{xy}=2\operatorname{tr}\Big{(}R_{x}R^{\dagger}_{y}\big{(}[\operatorname{id}\pm(1,2)]\varrho^{\otimes 2}\big{)}\Big{)}\quad\text{with}\quad R_{x}:=\frac{1}{2}[\operatorname{id}\mp(1,2)](E_{x}\otimes\operatorname{id})\,.$
Now we use the fact that $\frac{1}{2}[\operatorname{id}\pm(1,2)]$ is a projector and that $(1,2)A\otimes B=B\otimes A(1,2)$. Then $M^{\pm}\succeq 0$ when $\varrho\succeq 0$, because 666 This is inspired by Ref. [59, Lemma 7.2], where for any operator $A$, the inequality $(A\otimes A^{\dagger}(1,2))^{T_{2}}\succeq 0$ is derived. Here $T_{2}$ is the partial transpose in the second subsystem.
(114) $\frac{1}{2}[\operatorname{id}\pm(1,2)]\varrho^{\otimes 2}=\frac{1}{2}[\operatorname{id}\pm(1,2)]\,\,\varrho^{\otimes 2}\,\,\frac{1}{2}[\operatorname{id}\pm(1,2)]\succeq 0\,.$
Expand the matrix $M^{\pm}$ to obtain
$M^{\pm}_{xy}$ $=\operatorname{tr}\Big{(}(E_{x}\otimes\operatorname{id})(1\mp(1,2))(E^{\dagger}_{y}\otimes\operatorname{id})\big{(}[\operatorname{id}\pm(1,2)]\varrho^{\otimes 2}\big{)}\Big{)}$
$=\operatorname{tr}([E_{x}E^{\dagger}_{y}\otimes\operatorname{id}]\varrho^{\otimes 2})\mp\operatorname{tr}([E_{x}\otimes E^{\dagger}_{y}](1,2)\varrho^{\otimes 2})$
±$\quad\pm\operatorname{tr}([E_{x}E^{\dagger}_{y}\otimes\operatorname{id}](1,2)\varrho^{\otimes 2})-\operatorname{tr}([E_{x}\otimes E^{\dagger}_{y}]\varrho^{\otimes 2})$
(115) $=\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho)\operatorname{tr}(\varrho)\mp\operatorname{tr}(E_{x}\varrho E^{\dagger}_{y}\varrho)\pm\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho^{2})-\operatorname{tr}(E_{x}\varrho)\operatorname{tr}(E^{\dagger}_{y}\varrho)\,,$
The average of both matrices reads
(116) $\frac{M^{+}_{xy}+M^{-}_{xy}}{2}=\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho)-\operatorname{tr}(E_{x}\varrho)\operatorname{tr}(E^{\dagger}_{y}\varrho)\succeq 0\,,$
which are the entries of $M^{T}$. Therefore, $M\succeq 0$ and also $L=M\circ G\succeq 0$.
We now show that
(117) $(2^{n}K)\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu-\tilde{\Gamma}=\frac{1}{\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi(L)\,.$
The expression in Eq. (117), being a sum over positive semidefinite matrices, is positive semidefinite and thus proves that $\tilde{\Gamma}^{c}\succeq 0$. To see that Eq. (117) holds, write $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ using Eq. (96) together with Eq. (94) as
$\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ $=\frac{1}{4^{n}\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum_{v\in{\boldsymbol{4}}^{n}}\sum\limits_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi\left(\sigma_{v}\!\left(\Gamma\right)\right)\,$
(118) $=\frac{1}{4^{n}\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi\left(\sum_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}\!\left(\Gamma\right)\right)\,.$
Because $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu$ arises from a quantum code $\varrho=\Pi/K$, we can develop the inner sum of Eq. (8) as in Eq. (7),
$\sum_{v\in{\boldsymbol{4}}^{n}}\sigma_{v}\!\left(\Gamma\right)_{xy}$ $=\sum_{E_{v}\in\mathcal{E}_{n}}\Big{(}\operatorname{tr}(E_{v}E^{\dagger}_{x}\varrho)\operatorname{tr}(E_{y}E^{\dagger}_{v}\varrho)\Big{)}\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho)$
(119) $=\frac{2^{n}}{K}\operatorname{tr}(E_{y}E^{\dagger}_{x}\varrho)\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho)\,.$
Using Eq. (8) together with Eq. (8), the matrix $\bar{\Gamma}$ has entries
(120) $\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu_{xy}$ $=\frac{1}{(2^{n}K)\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi\left(S\right)\,,$
where $S_{xy}=\operatorname{tr}(E_{y}E^{\dagger}_{x}\varrho)\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho)$. With this notation, the matrix $L$ can be expressed as $L=S-\Gamma$. Using the definition Eq. (109) of $\tilde{\Gamma}$ and the identity Eq. (120) for $\bar{\Gamma}$, proves that Eq. (117) holds:
$\frac{1}{\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi(L)$
$=\frac{1}{\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi(S)\quad-\quad\frac{1}{\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi(\Gamma)$
(121) $=(2^{n}K)\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu-\tilde{\Gamma}.$
Recall that since $L$ is positive semidefinite, then $\pi(L)$ is also positive semidefinite for any $\pi\in\operatorname{Aut}(4,n)$. Eq. (8) is a sum over positive semidefinite matrices and thus $(2^{n}K)\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu-\tilde{\Gamma}\succeq 0$. A multiplication by $\big{(}K(2^{n}-1/K)\big{)}^{-1}$ yields Eq. (110). This proves the claim. ∎
8.1. Positive semidefinite constraints in the Terwilliger algebra
To make use of Proposition 16 in a symmetry-reduced semidefinite program, we need to express these constraints in terms of the Terwilliger algebra.
Proposition 17.
The matrices $\tilde{\Gamma}$ and $\tilde{\Gamma}^{c}$ can in the Terwilliger algebra be expanded as
(122) $\tilde{\Gamma}$ $=\sum_{(i,j,t,p)\in\mathcal{I}(q,n)}x^{t,p}_{i,j}M^{t,p}_{i,j}\,,$
(123) $\tilde{\Gamma}^{c}$ $=\frac{2^{n}}{K}\sum_{(i,j,t,p)\in\mathcal{I}(q,n)}(x^{0,0}_{i+j-t-p}-x^{t,p}_{i,j})M^{t,p}_{i,j}\,.$
Here $x^{t,p}_{i,j}=\frac{\lambda^{t,p}_{i,j}}{\gamma^{t,p}_{i,j}}$ with $\gamma^{t,p}_{i,j}$ defined in Eq. (77) and
(124) $\lambda^{t,p}_{i,j}=\!\!\!\!\sum\limits_{\begin{subarray}{c}x,y\in{\boldsymbol{4}}^{n}\\ |s(x)|=i\\ |s(y)|=j\\ \lvert s(x)\cap s(x)\rvert=t\\ \lvert\{m\,\mid\,x_{m}=y_{m}\neq 0\}\rvert=p\\ \end{subarray}}\!\!\!\!\!\!\Gamma_{xy}\,.$
Proof.
The matrix $\tilde{\Gamma}$ is invariant under $\operatorname{Aut}_{0}(q,n)$, and consequentially it can be expanded in the basis $M^{t,p}_{i,j}$ of the Terwilliger Algebra. With the definition from Eq. (83), the parameters $\lambda^{t,p}_{i,j}$ are then given by Eq. (124) and $x^{t,p}_{i,j}=\frac{\lambda^{t,p}_{i,j}}{\gamma^{t,p}_{i,j}}$.
From Eq. (110), we can express $\tilde{\Gamma}^{c}$ in the Terwilliger Algebra basis using Eq. (97),
(125) $\tilde{\Gamma}^{c}=\sum_{(i,j,t,p)\in\mathcal{I}(4,n)}(y_{i+j-t-p}-\frac{2^{n}}{K}x^{t,p}_{i,j})M^{t,p}_{i,j}\,.$
With this, Eq. (7) with $\Gamma_{xx}=\Gamma_{x0}$ shows that
(126) $y_{k}=(2^{n}K\gamma^{0,0}_{k,0})^{-1}\sum_{\begin{subarray}{c}x\in{\boldsymbol{4}}^{n}\\ |s(x)|=k\end{subarray}}\Gamma_{x0}=\frac{2^{n}}{K}x^{0,0}_{k,0}\,,$
The combination of Eq.  (125) with Eq.  (126) results in Eq.  (123). This ends the proof. ∎
8.2. Stabilizer codes
Consider the special case of a stabilizer code with stabilizer group $S$. Then we can recover the results in Propositions 16 and 17 in an alternative fashion for the matrices
(127) $\tilde{\Gamma}(\mathcal{S})$ $=\frac{1}{|\mathcal{S}|}\sum\limits_{v\in\mathcal{S}}\Gamma^{v}\,,$
(128) $\tilde{\Gamma}^{c}(\mathcal{S})$ $=\frac{1}{4^{n}-|\mathcal{S}|}\sum\limits_{v\in\mathcal{S}\setminus{\boldsymbol{4}}^{n}}\Gamma^{v}=\frac{1}{4^{n}-|\mathcal{S}|}\Big{(}4^{n}\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu-|\mathcal{S}|\tilde{\Gamma}(\mathcal{S})\Big{)}\,,$
that is, by replacing $Q=\{0^{n}\}$ with $Q=\mathcal{S}$ where $\mathcal{S}$ contains the unsigned Pauli strings from $S$ mapped to $4^{n}$.
Then as in the classical case discussed in Sections 6.3 and 6.4, these matrices are both positive semidefinite and invariant under $\operatorname{Aut}_{0}(4,n)$ and can thus be expressed in a basis of the Terwilliger algebra $\mathcal{A}_{4,n}$, recovering the same constraints on $x^{t,p}_{i,j}$ as in the general case. This follows from the fact that by the definition of $\Gamma^{v}$ in Eq. (94), it holds that $\Gamma^{v}=\Gamma^{0}$ for any $v\in\mathcal{S}$ leading to $\tilde{\Gamma}=\Gamma^{0}$, and that for a stabilizer code $|S|=|\mathcal{S}|=2^{n}/K$.
Let us explain this in more detail: Consider Eq. (94),
(129) $\tilde{\Gamma}(\mathcal{S})$ $=\frac{1}{|\mathcal{S}|}\sum\limits_{v\in\mathcal{S}}\Gamma^{v}=\frac{1}{|\mathcal{S}|}\frac{1}{\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum\limits_{v\in\mathcal{S}}\sum\limits_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi\left(\sigma_{v}\!\left(\Gamma\right)\right)\,.$
Because $s\Pi=\Pi s=\Pi$ for all $s\in S$, we have for any $v\in\mathcal{S}$ that
$\sigma_{v}(\Gamma)_{xy}$ $=\operatorname{tr}(E^{\dagger}_{x}\varrho E_{v})\operatorname{tr}(E_{y}E^{\dagger}_{v}\varrho)\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho)$
(130) $=\operatorname{tr}(E^{\dagger}_{x}\varrho)\operatorname{tr}(E_{y}\varrho)\operatorname{tr}(E_{x}E^{\dagger}_{y}\varrho)=\Gamma_{xy}\,.$
This leads to
(131) $\tilde{\Gamma}(\mathcal{S})=\frac{1}{\lvert\operatorname{Aut}_{0}(4,n)\rvert}\sum\limits_{\pi\in\operatorname{Aut}_{0}(4,n)}\pi\left(\Gamma\right)=\tilde{\Gamma}(0^{n})\,,$
which recovers the general case [see Eq. (109)]. In analogy to Proposition (17), the expansion in the Terwilliger algebra reads
(132) $\tilde{\Gamma}(\mathcal{S})$ $=\sum_{(i,j,t,p)\in\mathcal{I}(4,n)}x^{t,p}_{i,j}M^{t,p}_{i,j}\,,$
(133) $\tilde{\Gamma}^{c}(\mathcal{S})$ $=\frac{|\mathcal{S}|}{4^{n}-|\mathcal{S}|}\sum_{(i,j,t,p)\in\mathcal{I}(q,n)}(x^{0,0}_{i+j-t-p}-x^{t,p}_{i,j})M^{t,p}_{i,j}\,,$
since for a stabilizer code $|S|=|\mathcal{S}|=2^{n}/K$.
8.3. Symmetry-reduced constraints
We now translate all constraints on $\Gamma$ into constraints on $x^{t,p}_{i,j}$. Similar to the classical case in Ref. [23, Eq. (49)], one has:
Proposition 18.
With $\tilde{\Gamma}$ and $\tilde{\Gamma}^{c}$ expanded as in Proposition 17, the $x^{t,p}_{i,j}$ satisfy:
$\operatorname{(i)}$ $\quad x^{00}_{00}=1\,,$
$\operatorname{(ii)}$ $\quad x^{t,p}_{i,j}=0\quad\quad\,\,\text{if}\quad t-p\quad\text{is odd}\,,$
$\operatorname{(iii)}$ $\quad x^{t,p}_{i,j}=x^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}\quad\text{if}\quad t-p=t^{\prime}-p^{\prime}\quad\text{is even}\quad\text{and}$
$\quad\quad\quad\quad\quad\quad\,\,(i,j,i+j-t-p)\quad\text{a permutation of}\quad(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,,$
$\operatorname{(iv)}$ $\quad\sum^{n}_{i=0}\gamma^{0,0}_{i,0}x^{0,0}_{i,0}=\frac{2^{n}}{K}\,,$
(134) $\operatorname{(v)}$ $\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ k=i+j-t-p\end{subarray}}\!\!\!\gamma^{t,p}_{i,j}x^{t,p}_{i,j}=\frac{2^{n}}{K}\gamma^{0,0}_{k,0}x^{0,0}_{k,0}.$
Proof.
Before starting with the proof, it is useful to recall from Eq. (124) that
(135) $\gamma^{t,p}_{i,j}x^{t,p}_{i,j}=\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\begin{subarray}{c}x,y\in{\boldsymbol{4}}^{n}\\ |s(x)|=i\\ |s(y)|=j\\ \lvert s(x)\cap s(x)\rvert=t\\ \lvert\{m\,\mid\,x_{m}=y_{m}\neq 0\}\rvert=p\\ \end{subarray}}\!\!\!\!\!\!\Gamma_{xy}\,.$
(i) From $\Gamma_{00}=1$, $(M_{0,0}^{0,0})_{00}=1$ and zero otherwise, and $\gamma^{0,0}_{0,0}=1$ it follows that $x^{0,0}_{0,0}=1$.
(ii) Remark 12 shows that, given $E_{x},E^{\dagger}_{y}\in\mathcal{E}_{n}$, the value $t-p$ counts the number of places in which a nontrivial Pauli element in $E_{x}$ does not cancel a nontrivial Pauli element in $E^{\dagger}_{y}$. As a consequence, if $t-p$ is even, then $E_{x}$ and $E^{\dagger}_{y}$ commute; if $t-p$ is odd, then $E_{x}$ and $E^{\dagger}_{y}$ anticommute. Since $\Gamma_{xy}=0$ if $E_{x}E^{\dagger}_{y}+E^{\dagger}_{y}E_{x}=0$, Eq. (135) then leads to $x^{t,p}_{i,j}=0$ if $t-p$ is odd. This proves condition (ii).
(iii) Eq. (5.2) shows that
$\Gamma_{xy}=\Gamma_{x^{\prime}y^{\prime}}\,,\quad\quad\quad$ $\text{if}\quad\quad E_{x}E^{\dagger}_{y}-E^{\dagger}_{y}E_{x}=0\quad\text{and}$
(136) $(E^{\dagger}_{x},E_{y},\omega_{xy}E_{x}E^{\dagger}_{y})\quad\text{a permutation of}\quad(E^{\dagger}_{x^{\prime}},E_{y^{\prime}},\omega_{x^{\prime}y^{\prime}}E_{x^{\prime}}E^{\dagger}_{y^{\prime}})\,,$
Recall from Remark 12 that $\operatorname{wt}(E_{x}E^{\dagger}_{y})=i+j-t-p$ and note that $t-p=t^{\prime}-p^{\prime}$ for any permutation from Eq. (8.3). Therefore if $t-p=t^{\prime}-p^{\prime}$ is even and $(i,j,i+j-t-p)$ is a permutation of $(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})$, then
(137) $\gamma^{t,p}_{i,j}x^{t,p}_{i,j}\quad=\sum\limits_{\begin{subarray}{c}x,y\in{\boldsymbol{4}}^{n}\\ |s(x)|=i\\ |s(y)|=j\\ \lvert s(x)\cap s(x)\rvert=t\\ \lvert\{m\,\mid\,x_{m}=y_{m}\neq 0\}\rvert=p\\ \end{subarray}}\!\!\!\!\!\!\Gamma_{xy}=\sum\limits_{\begin{subarray}{c}x^{\prime},y^{\prime}\in{\boldsymbol{4}}^{n}\\ |s(x^{\prime})|=i^{\prime}\\ |s(y^{\prime})|=j^{\prime}\\ \lvert s(x^{\prime})\cap s(x^{\prime})\rvert=t^{\prime}\\ \lvert\{m\,\mid\,x^{\prime}_{m}=y^{\prime}_{m}\neq 0\}\rvert=p^{\prime}\\ \end{subarray}}\!\!\!\!\!\!\Gamma_{x^{\prime}y^{\prime}}\quad=\quad\gamma^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}x^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}\,.$
Furthermore, one can check [Eq. (77)] that $\gamma^{t,p}_{i,j}=\gamma^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}$ which leads to condition (iii).
For the last two conditions we derive the projector constraints in Eq. (62) in terms of $x^{t,p}_{i,j}$ using that $\Gamma_{x0}=\Gamma_{xx}$:
(iv) Recall that $\sum_{x\in{\boldsymbol{4}}^{n}}\Gamma_{x0}=\frac{2^{n}}{K}$ as shown in Eq. (62). With this, Eq. (135) leads to condition (iv).
(v) Here we use the constraint $\sum_{x\in{\boldsymbol{4}}^{n}}\Gamma_{xy}=\frac{2^{n}}{K}\Gamma_{x0}$ [Eq. (62)] in combination of conditions (ii)-(iii). First, we write such constraint in terms of $x^{t,p}_{i,j}$ by using Eq. (8.3). This leads to
(138) $\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i=k\end{subarray}}\gamma^{t,p}_{i,j}x^{t,p}_{i,j}=\frac{2^{n}}{K}\gamma^{0,0}_{k,0}x^{0,0}_{k,0}\,.$
Condition (ii)-(iii), show that $x^{t,p}_{i,j}=x^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}$ if $t-p=t^{\prime}-p^{\prime}$, $i=i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime}$, and $j=j^{\prime}$. For this case also $\gamma^{t,p}_{i,j}=\gamma^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}$ by Eq. (77) and thus, $\gamma^{t,p}_{i,j}x^{t,p}_{i,j}=\gamma^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}x^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}$. This allows us to write Eq. (138) as
(139) $\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i=k\end{subarray}}\gamma^{t,p}_{i,j}x^{t,p}_{i,j}=\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},t^{\prime},p^{\prime})\in\mathcal{I}(4,n)\\ i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime}=k\end{subarray}}\gamma^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}x^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}=\frac{2^{n}}{K}\gamma^{0,0}_{k,0}x^{0,0}_{k,0}\,,$
leading to condition (v). This ends the proof. ∎
The $x^{t,p}_{i,j}$ is further constrained by the Knill-Laflamme conditions:.
Proposition 19.
The Knill-Laflamme conditions as stated in Eq. (63) imply that
(140) $\frac{K}{2^{n}}\sum^{n}_{i=0}K_{j}(i;n)\gamma^{0,0}_{i,0}x^{0,0}_{i,0}\quad\geq\quad\gamma^{0,0}_{j,0}x^{0,0}_{j,0}\quad\text{with equality for}\quad 0<j<\delta\,,$
According to Eq. (66), pure codes additionally satisfy
(141) $x^{t,p}_{i,j}=0\quad\text{if}\quad\{i,j,i+j-t-p\}\cap\{1,\cdots,\delta-1\}\neq\emptyset\,,$
which is the condition from Ref. [23, Eq. (49)(iv)].
Proof.
Eq. (135) shows that $\gamma^{0,0}_{j,0}x^{0,0}_{j,0}=\sum_{x\in{\boldsymbol{4}}^{n},\,\operatorname{wt}(E_{x})=j}\Gamma_{x0}$. This allows us to write Eq. (63) as Eq. (140). By Eq. (135), the condition for pure codes stated in Eq. (66) can be written as $x^{t,p}_{i,j}=0$ for $0<i<\delta$ or $0<j<\delta$ or $0<i+j-t-p<\delta$. But this is nothing else than Eq. (141). This ends the proof. ∎
Note that, the constraint $\sum^{n}_{i=0}\gamma^{0,0}_{i,0}x^{0,0}_{i,0}=\frac{2^{n}}{K}$ is equivalent to $\frac{K}{2^{n}}\sum^{n}_{i=0}K_{0}(i;n)\gamma^{0,0}_{i,0}x^{0,0}_{i,0}=\gamma^{0,0}_{0,0}x^{0,0}_{0,0}$, since $K_{0}(i;n)=1$ for all $i$ and $\gamma^{0,0}_{0,0}x^{0,0}_{0,0}=1$. It is thus possible to include either one or the other, according to convenience.
Remark 20.
For stabilizer codes, one also has the condition in Ref. [23, Eq. (49)(ii)] which holds for classical codes,
(142) $0\leq x_{i,j}^{t,p}\leq x_{i,0}^{0,0}\,.$
Proof.
Eq. (64) shows that stabilizer codes satify $0\leq\Gamma_{xy}\leq\Gamma_{xx}$ Recall that $\tilde{\Gamma}\propto\sum_{\pi\in\operatorname{Aut}(4,n)}\pi(\Gamma)$ [see Eq. (109)]. Thus the same relation holds after averaging and $0\leq\tilde{\Gamma}_{xy}\leq\tilde{\Gamma}_{xx}$ since $\pi$ permutes rows and columns of $\Gamma$.
Because $\tilde{\Gamma}$ is spanned by $M_{i,j}^{t,p}$, which have entries in $\{0,1\}$ is follows that $0\leq x^{t,p}_{i,j}\leq{x^{i,i}_{i,i}}$. According to condition (iii) from Proposition 18, $x^{i,i}_{i,i}=x^{0,0}_{i,0}$ and therefore, $0\leq x^{t,p}_{i,j}\leq x^{0,0}_{i,0}$. ∎
8.4. Symmetry-reduced SDP bound
We now state the symmetry reduced version of the main semidefinite programming bound in Eq. (5.7). It includes all constraints on $x^{t,p}_{i,j}$ derived in the previous sections [Propositions 16 to 19].
Theorem 21.
If a $(\!(n,K,d)\!)_{2}$ quantum code exists, then the following semidefinite program is feasible:
find $x^{t,p}_{i,j}$
subject to $x^{0,0}_{0,0}=1,$
$x^{t,p}_{i,j}=0\quad\quad\quad\text{if}\quad t-p\quad\text{is odd}\,,$
$x^{t,p}_{i,j}=x^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}\,\,\,\,\quad\text{if}\quad t-p=t^{\prime}-p^{\prime}\quad\text{is even}\quad\text{and}$
$\quad\quad\quad\quad\quad\quad\,\,(i,j,i+j-t-p)\quad\text{a permutation of}\quad(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,,$
$\sum^{n}_{i=0}\gamma^{0,0}_{i,0}x^{0,0}_{i,0}=\frac{2^{n}}{K},\quad\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ k=i+j-t-p\end{subarray}}\gamma^{t,p}_{i,j}x^{t,p}_{i,j}=\frac{2^{n}}{K}\gamma^{0,0}_{k,0}x^{0,0}_{k,0}\,,$
$K2^{-n}\sum^{n}_{i=0}K_{j}(i;n)\gamma^{0,0}_{i,0}x^{0,0}_{i,0}\quad\geq\quad\gamma^{0,0}_{j,0}x^{0,0}_{j,0}\quad\text{with equality for}\quad 0<j<\delta\,,$
$\bigoplus_{\begin{subarray}{c}a,k\in\mathds{N}_{0}\\ 0\leq a\leq k\leq n+a-k\end{subarray}}\left(\sum\limits_{\begin{subarray}{c}t,p\in\mathds{N}_{0}\\ 0\leq p\leq t\leq i,j\\ i+j\leq t+n\end{subarray}}\alpha(i,j,t,p,a,k)x_{i,j}^{t,p}\right)_{i,j=k}^{n+a-k}\succeq 0\,,$
(143) $\bigoplus_{\begin{subarray}{c}a,k\in\mathds{N}_{0}\\ 0\leq a\leq k\leq n+a-k\end{subarray}}\left(\sum\limits_{\begin{subarray}{c}t,p\in\mathds{N}_{0}\\ 0\leq p\leq t\leq i,j\\ i+j\leq t+n\end{subarray}}\alpha(i,j,t,p,a,k)(x_{i+j-t-p,0}^{0,0}-x_{i,j}^{t,p})\right)_{i,j=k}^{n+a-k}\succeq 0\,,$
where $\alpha(i,j,t,p,a,k)$ is given by Eq. (6.3).
As with the previous results, Theorem 21 applies equally to stabilizer and non-stabilizer codes. Furthermore, it is at least as strong as the linear programming bounds from Eq. (3.4) without the shadow inequalities $S_{j}(\rho)\geq 0$. Note that $\gamma_{j,0}^{0,0}x_{j,0}^{0,0}$ approximates $A_{j}(\varrho)$. This allows us to add the rest of linear programming constraints as done in Eq. (5.7). The SDP above can then be further restricted by adding:
  1. (i)
    The shadow inequalities [see Eq. (18)] leads to
    (144) $2^{-n}\sum^{n}_{i=0}(-1)^{i}K_{j}(i;n)\gamma_{i,0}^{0,0}x_{i,0}^{0,0}\quad\geq\quad 0\,,$
    with equality for $K=1$ and $n-j$ odd [57, Theorem 12].
  2. (ii)
    For stabilizer codes, $0\leq x_{i,j}^{t,p}\leq x_{i,0}^{0,0}$ [see Remark 20] and
    (145) $\sum_{j\geq 0}\gamma_{2j,0}^{0,0}x_{2j,0}^{0,0}=\begin{cases}2^{n-\log_{2}(K)-1}\quad\text{(Type I)\,,}\\ 2^{n-\log_{2}(K)}\quad\quad\!\text{(Type II)\,.}\end{cases}$
  3. (iii)
    For pure codes [see Eq. (141)],
    (146) $x^{t,p}_{i,j}=0\quad\text{if}\quad\{i,j,i+j-t-p\}\cap\{1,\cdots,\delta-1\}\neq\emptyset\,,$
Therefore, SDP (21) is at least as strong as the linear programming bounds in Eq. (3.4).
We conclude this section by point out that the main semidefinite bound in Eq. (5.7) with scaling of $O(4^{n})$ is now symmetry reduced to Eq. (21) with scaling of $O(n^{4})$. This fact allows us to handle codes with larger $n$.
9. Non-existence of quantum codes
9.1. The Lovász bound refutes a $(\!(7,1,4)\!)_{2}$ code
Using the feasibility version of the symmetry reduced Lovász number SDP for self-dual quantum codes [see Eq. (5.8)], we prove the non-existence of a $(\!(7,1,4)\!)_{2}$ code. The feasibility version includes as an extra condition the constraint $\sum_{x\in{\boldsymbol{4}}^{n}}\Gamma_{xx}=2^{n}$:
find $\Gamma$
subject to $\Gamma_{00}=1\,,$
$\Gamma_{x0}=\Gamma_{xx}\,,$
$\Gamma\succeq 0\,,$
$\Gamma_{xy}=0\quad\text{if}\quad x\sim y\,,$
(147) $\operatorname{tr}(\Gamma)=2^{n}\,.$
Following Section 8.4 we symmetry reduce the SDP (9.1) as done for the main semidefinite programming bound [Eq. (5.7)]. First apply conditions (i) and (iii) from Proposition (18) to map $\Gamma_{00}=1$ and $\Gamma_{x0}=\Gamma_{xx}$ to $x^{0,0}_{i,0}$ and $x^{0,0}_{i,0}=x^{i,i}_{i,i}$, respectively. Then consider the block-diagonalization of $\tilde{\Gamma}\succeq 0$ as done with $\tilde{R}$ [Eq. (85)]. Recall that $\Gamma_{xy}=0$ if $x\sim y$. This implies that $x^{t,p}_{i,j}=0$ if $t-p$ is odd or $\{i,j,i+j-t-p\}\cap\{1,\cdots,\delta-1\}\neq\emptyset$. Here we have used the condition for pure codes [Eq. (141)] and anti-commutator condition (ii) of Proposition (18). Finally, $\operatorname{tr}(\Gamma)=1$ is equivalent to $\sum_{x\in{\boldsymbol{4}}^{n}}\Gamma_{x0}=2^{n}$ due to $\Gamma_{x0}=\Gamma_{xx}$. This implies $\sum^{n}_{i=0}x^{0,0}_{i,0}\gamma^{0,0}_{i,0}=\frac{2^{n}}{K}$ by condition (iv) in Proposition 18. Thus the symmetry reduced version of Eq. (9.1) reads:
find $x^{t,p}_{i,j}$
subject to $x^{0,0}_{0,0}=1\,,$
$x^{0,0}_{i,0}=x^{i,i}_{i,i}\,,$
$\bigoplus_{\begin{subarray}{c}a,k\in\mathds{N}_{0}\\ 0\leq a\leq k\leq n+a-k\end{subarray}}\left(\sum\limits_{\begin{subarray}{c}t,p\in\mathds{N}_{0}\\ 0\leq p\leq t\leq i,j\\ i+j\leq t+n\end{subarray}}\alpha(i,j,t,p,a,k)x_{i,j}^{t,p}\right)_{i,j=k}^{n+a-k}\succeq 0$
$x^{t,p}_{i,j}=0\quad\text{if}\quad t-p\quad\text{is odd}$
$\hskip 49.79231pt\quad\text{or}\quad\{i,j,i+j-t-p\}\cap\{1,\cdots,\delta-1\}\neq\emptyset\,,$
(148) $\sum^{n}_{i=0}\gamma^{0,0}_{i,0}x^{0,0}_{i,0}=2^{n}\,,$
where $\alpha(i,j,t,p,a,k)$ is given by Eq. (6.3).
Remark 22.
Here we only took $\tilde{\Gamma}\succeq 0$ into account, leaving out the constraint $\tilde{\Gamma}^{c}\succeq 0$.
The dual problem can be used to prove infeasibility of a primal problem [see Eq. (37)]. Appendix A, Proposition 25 shows that the dual program is:
$\alpha\quad=\quad\max_{Y^{(a,k)}\,,\,{w}}\quad$ $(2^{n}-1)w-y_{0,0}^{0,0}$
subject to $Y^{(a,k)}\succeq 0\,,$
$y^{t,p}_{i,j}=\frac{1}{\gamma^{t,p}_{i,j}}\sum^{\min(i,j)}_{k=0}\sum^{k}_{a=\max(i,j)+k-n}\alpha(i,j,t,p,a,k)Y^{(a,k)}_{i-k,j-k}\,,$
$y_{i,i}^{i,i}+w+2y_{i,0}^{0,0}=0\quad\quad\quad\text{if}\quad\delta\leq i\leq n\,,$
$y^{t,p}_{i,j}=0\hskip 80.00012pt\text{if}\quad i,j\neq 0\,,\quad t-p\quad\text{is even }\quad$
(149) $\hskip 116.00023pt\quad\text{and}\quad i+j-t-p\geq\delta\,.$
where the dual variables are real matrices $Y^{(a,k)}$ of size $n+a-2k$ with $0\leq a\leq k\leq n+a-k$.
If $\alpha>0$, then there do not exist a set of $x^{t,p}_{i,j}$ satisfying the constraints from Eq. (9.1). This is the case for the $(\!(7,1,4)\!)_{2}$ code:
Observation 23.
The matrices $Y^{(a,k)}$ written in Appendix B are a solution for Eq. (9.1) for $n=7$ and $\delta=4$ with $\alpha=0.58>0$, maximum constraint violation of $\mathcal{O}(10^{-16})$, and minimum eigenvalue $\lambda_{\min}\approx 0.013$. This provides an infeasibility certificate for the primal program and shows that a $(\!(7,1,4)\!)_{2}$ quantum code does not exist.
Note that the dual variable $w$ is determined by the matrices $Y^{(a,k)}$ since $w=-y^{i,i}_{i,i}-y^{0,0}_{i,0}$ if $\delta\leq i\leq n$. In particular, $w=-y^{n,n}_{n,n}-y^{0,0}_{n,0}$.
9.2. The symmetry-reduced SDP refutes $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes
We first remark the following: in principle, adding constraints in the primal program makes a semidefinite program stronger. From a numerical perspective however this can be disadvantageous: The more constraints appear in the primal program, the larger is the span of the dual feasible region. As a consequence, the dual program can numerically be more costly to solve. For obtaining a nice infeasibility certificate it can thus help to drop a number of constraints.
For refuting the $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ code, the following relaxation of the SDP (21) of Theorem 21 is enough:
find $x^{t,p}_{i,j}$
subject to $x^{0,0}_{0,0}=1,$
$x^{t,p}_{i,j}=0\quad\quad\quad\text{if}\quad t-p\quad\text{is odd}\,,$
$x^{t,p}_{i,j}=x^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}\,,\quad\text{if}\quad t-p=t^{\prime}-p^{\prime}\quad\text{is even}\quad\text{and}$
$\quad\quad\quad\quad\quad\quad\,\,(i,j,i+j-t-p)\quad\text{a permutation of}\quad(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,,$
$\sum^{n}_{i=0}\gamma^{0,0}_{i,0}x^{0,0}_{i,0}=\frac{2^{n}}{K},\quad\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ k=i+j-t-p\end{subarray}}\gamma^{t,p}_{i,j}x^{t,p}_{i,j}=\frac{2^{n}}{K}\gamma^{0,0}_{k,0}x^{0,0}_{k,0}\,,$
$K2^{-n}\sum^{n}_{i=0}K_{j}(i,n)\gamma^{0,0}_{i,0}x^{0,0}_{i,0}\quad=\quad\gamma^{0,0}_{j,0}x^{0,0}_{j,0}\quad\quad\text{if}\quad 0<j<\delta\,,$
(150) $\bigoplus_{\begin{subarray}{c}a,k\in\mathds{N}_{0}\\ 0\leq a\leq k\leq n+a-k\end{subarray}}\left(\sum\limits_{\begin{subarray}{c}t,p\in\mathds{N}_{0}\\ 0\leq p\leq t\leq i,j\\ i+j\leq t+n\end{subarray}}\alpha(i,j,t,p,a,k)x_{i,j}^{t,p}\right)_{i,j=k}^{n+a-k}\succeq 0\,,$
where $\alpha(i,j,t,p,a,k)$ is given by Eq. (6.3). The SDP (9.2) is obtained from SDP 21 by removing the constraint corresponding to $\tilde{\Gamma}^{c}\succeq 0$ and to only take the equality conditions in $K2^{-n}\sum^{n}_{i=0}K_{j}(i,n)\gamma^{0,0}_{i,0}x^{0,0}_{i,0}\,\geq\,\gamma^{0,0}_{j,0}x^{0,0}_{j,0}$.
Appendix A, Proposition 26 shows that the dual program of Eq. (9.2) is:
$\alpha\quad=\quad\max_{Y^{(a,k)},Q_{i},C_{i}}\quad$ $(KD_{0}-C_{0})+\left(\frac{2^{n}}{K}-1\right)Q_{0}-{y^{0,0}_{0,0}}\quad$
subject to $Y^{(a,k)}\succeq 0,$
$y^{t,p}_{i,j}\ =\frac{1}{\gamma^{t,p}_{i,j}}\sum^{\min(i,j)}_{k=0}\sum^{k}_{a=\max(i,j)+k-n}\alpha(i,j,t,p,a,k)Y^{(a,k)}_{i-k,j-k}\,,$
$D_{i}=2^{-n}\sum^{\delta-1}_{j=0}K_{j}(i,n)C_{j}\,,$
$(KD_{i}-C_{i})+\left(\frac{2^{n}}{K}-2\right)Q_{i}-2y^{0,0}_{i,0}-Q_{0}-y^{i,i}_{i,i}=0\quad\text{if}\quad 0<i<\delta\,,$
$KD_{i}+\left(\frac{2^{n}}{K}-2\right)Q_{i}-2y^{0,0}_{i,0}-Q_{0}-y^{i,i}_{i,i}=0\quad\quad\quad\quad\text{if}\quad\delta\leq i\leq n\,,$
$\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},t^{\prime},p^{\prime})\in\mathcal{I}(4,n)\\ (t-p)=(t^{\prime}-p^{\prime})\\ (i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,\text{a}\\ \,\text{permutation of}\,(i,j,i+j-t-p)\\ \end{subarray}}\!\!\!\!\!\!\!\!\!\!\left(y^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}+Q_{i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime}}\right)=0\quad\hskip 33.00008pt\text{if}\quad(t-p)\quad\text{even}$
(151) $\hskip 170.71652pt\quad\text{and}\quad i,j,i+j-t-p\neq 0\,.$
Here the dual variables are real matrices $Y^{(a,k)}$ of size $n+a-2k$ with $0\leq a\leq k\leq n+a-k$ and $Q_{i},C_{i}\in\mathds{R}$ with $0\leq i\leq n$.
Observation 24.
The SDP from Eq. (9.2) has solutions with $\alpha>0$ for the parameters $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$. In particular, for the $(\!(8,9,3)\!)_{2}$, there is a dual solution with $\alpha\approx 0.989$, maximum constraint violation of $\mathcal{O}(10^{-13})$, and minimum eigenvalue $\lambda_{\min}\approx 0.0010$. For $(\!(10,5,4)\!)_{2}$, there is a dual solution with $\alpha\approx 1$, maximum constraint violation of $\mathcal{O}(10^{-13})$ and minimum eigenvalue $\lambda_{\min}\approx 0.050$. As a consequence, codes with parameters $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ do not exist.
9.3. Code availability
To solve the SDPs from Eq. (9.1) and Eq. (9.2), we used the Python API PICOS [62] toghether with the SDP solver CVXOPT [43]. The program can be found online in https://github.com/ganglesmunne/SDP_bounds_for_quantum_codes. To obtain solutions with some rational values we run the SDP iteratively, where in each iteration an an entry of $Y^{(a,k)}$ was forced to be equal to a rounded previous entry. This procedure leads to the rational solution in Appendix B.
10. Discussion
We provided a complete semidefinite programming hierarchy for the existence of quantum codes, and relaxations that recover a Lovász bound for self-dual quantum codes, a quantum Delsarte bound, and the quantum version of Schrijver and Gijswijt’s SDP bounds that use the Terwilliger algebra. With this we show that $(\!(7,1,4)\!)_{2}$, $(\!(8,9,3)\!)_{2}$, and $(\!(10,5,4)\!)_{2}$ codes do not exist. While the non-existence of $(\!(7,1,4)\!)_{2}$ was known previously through analytical methods [30], to our knowledge the latter two bounds seem to be new [61, Table III].
An interesting aspect are the changes that appear in the quantum versions of these semidefinite programming bounds. These are perhaps best illustrated in the Lovász bound [Eqs. (3.5) and (3.5)]: in the classical case the Hamming distance indicates which entries in $M$ are set to zero. For quantum codes also the anti-commutation relations between pairs of observables yield constraints: $\Gamma_{ab}$ is additionally set to zero if the Pauli strings $E_{a}$, $E_{b}$ anticommute. This can be captured by a graph with loops, and makes the quantum Lovász bound naturally fit into a recently developed framework for a hypergraph Lovász $\vartheta$ number [10]. In particular, this leads to both different quantum Lovász and quantum Delsarte bounds compared to classical ${\bf{4}}^{n}$ codes.
A further interesting aspect is the use of the quantum MacWilliams identity to recover expressions of the form $\operatorname{tr}(E\rho E^{\dagger}\varrho)$ that, in principle, are outside of the state polynomial optimization framework. One can hope that this opens the door to the application of state polynomial optimization to optimize over arbitrary expressions in observables and states. A possible approach is to use higher-order generalizations of quantum MacWilliams identities such as presented in Ref. [69].
We note that our hierarchy also applies to the existence of absolutely maximally entangled and $m$-uniform states, these corresponding to self-dual (pure) codes [60, 28]. With small modifications, the hierarchy also lower bounds the average purities of $n$-qubit states, such as found in the Meyer-Wallach [45, 8, 66], potential of multipartite entanglement [19], and concentratable entangled measures [63].
We end with some open questions:
  • Can one extend the Lovász and Delsarte bounds for self-dual quantum codes to codes with $K>1$? This can possibly achieved by purification of the code with a reference system.
  • The shadow bounds provide powerful constraints, e.g. these alone refute the existence of a $(\!(4,1,3)\!)_{2}$ code. Is there a formulation of the generalized shadow bounds that can be useful as part of an SDP hierarchy [58]?
  • Can our SDP bounds be strengthened using the techniques developed by Laurent [39], the split Terwilliger algebra [70], or by 4-point bounds [24, 40]?
  • Can similar methods be established for other types of codes [1], e.g. subsystem codes [2], hybrid quantum-classical codes [25, 46], quantum MDS codes [29] codes over arbitrary quantum channels [49], or entanglement-assisted codes [37]? Likewise, can LP and SDP bounds be derived for quantum spherical codes [31] by analyzing the Johnson scheme as in the classical case [14, 5]?
  • The Hamming and Singleton bounds can be recovered from the both the classical Delsarte bound [12]. In the quantum case, the quantum Singleton bound can be obtained, as well as a quantum Hamming bound for non-degenerate codes [3]. It is an open question whether a quantum Hamming bound applies also to degenerate codes, with recent results suggesting that it does [11, 48]. Can the SDP techniques introduced in this paper be used to prove that the quantum Hamming bound must apply to any quantum code?
  • In the case of pure codes the Knill-Laflamme constraints on the variables $x_{i,j}^{t,p}$ in Proposition 19, Eq. (141) appear to be stronger than those of the linear programming bounds that make use of the quantum MacWilliams identity. Does there exist a quantum MacWilliams-type identity also for the $x_{i,j}^{t,p}$ (e.g. similar to Ref. [69]), allowing for finer-grained constraints on the minimum distance of impure codes?
  • How can this be generalized to qudit codes; can a characterization of quantum algebraic tori be used in place of the result by Gastineau-Hills [50]?
Finally, our approach is formally dimension-free. That is, on can target scenarios where the Hilbert space is not known. This could be interesting for the study of quantum version of the quantum zero-error communication capacity of a quantum graph [17].
Appendix A Dual programs
Here we derive the dual of the primal programs from Section 9 to obtain infeasibility certificates for codes with parameters $(\!(7,1,4)\!)_{2}$, $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$.
Following Boyd and Vandenberghe [7, Section 4.6.2, Example 5.12], we first explain such procedure for the case when the primal in its standard form [see Eq. (3.8)] which reads
minimize $\quad\operatorname{tr}(CX)$
subject to $\quad\operatorname{tr}(A_{i}X)=b_{i}\quad\text{for}\quad i=1,\dots,p\,,$
(152) $\quad X\succeq 0\,.$
The transition between a primal and the dual problems is done via their Lagrangian: associate with each equality constraint $\operatorname{tr}(A_{i}X)=b_{i}$ a dual variable $y_{i}\in\mathds{R}$ and with the semidefinite constraint $X\succeq 0$ a dual semidefinite variable $Y\succeq 0$. The Lagrangian is then defined as
$\mathcal{L}(X,y_{i},Y)$ $=\operatorname{tr}(CX)-\operatorname{tr}(YX)+\sum^{p}_{i=1}y_{i}\big{(}\operatorname{tr}(A_{i}X)-b_{i}\big{)}$
(153) $=\operatorname{tr}\Big{(}\big{(}C-Y+\sum^{p}_{i=1}y_{i}A_{i}\big{)}X\Big{)}-\sum_{i=1}^{p}y_{i}b_{i}\,,$
so the dual objective function is
(154) $\inf_{X}\mathcal{L}(X,y_{i},Y)=\begin{cases}-\sum^{p}_{i=1}y_{i}b_{i}\quad\text{if}\quad\sum^{p}_{i=1}y_{i}A_{i}-Y+C=0\,,\\ -\infty\quad\quad\quad\quad\text{otherwise}\,.\end{cases}$
The dual problem can then be expressed as
maximize $-\sum^{p}_{i=1}y_{i}b_{i}$
subject to $C+\sum^{p}_{i=1}y_{i}A_{i}=Y\quad\text{for}\quad i=1,\dots,p\,,$
(155) $Y\succeq 0\,.$
By eliminiating the slack variable $Y$ and redefining $y_{i}$ with $-y_{i}$, one obtains
maximize $\sum^{p}_{i=1}y_{i}b_{i}$
(156) subject to $C-\sum^{p}_{i=1}y_{i}A_{i}\succeq 0\quad\text{for}\quad i=1,\dots,p\,.$
A.1. Dual of the Lovász bound
We use this to write the dual program of the feasibility version of the symmetry reduced quantum Lovász number for self-dual codes,
Proposition 25.
The dual program of Eq. (9.1) is
$\alpha\quad=\quad\max_{Y^{(a,k)}\,,\,{w}}\quad$ $(2^{n}-1)w-y_{0,0}^{0,0}$
subject to $Y^{(a,k)}\succeq 0\,,$
$y^{t,p}_{i,j}=\frac{1}{\gamma^{t,p}_{i,j}}\sum^{\min(i,j)}_{k=0}\sum^{k}_{a=\max(i,j)+k-n}\alpha(i,j,t,p,a,k)Y^{(a,k)}_{i-k,j-k}\,,$
$y_{i,i}^{i,i}+w+2y_{i,0}^{0,0}=0\quad\quad\quad\text{if}\quad\delta\leq i\leq n\,,$
$y^{t,p}_{i,j}=0\hskip 80.00012pt\text{if}\quad i,j\neq 0\,,\quad t-p\quad\text{is even }\quad$
(157) $\hskip 116.00023pt\quad\text{and}\quad i+j-t-p\geq\delta\,.$
where $Y^{(a,k)}$ with $0\leq a\leq k\leq n+a-k$ are real square matrices of size $(n+a-2k)$.
Proof.
For the derivation of the dual program, we slightly modify the conditions from Eq. (9.1) when $x^{t,p}_{i,j}=0$ to obtain the equivalent formulation:
find $x^{t,p}_{i,j}$
subject to $x^{0,0}_{0,0}=1,\quad x^{0,0}_{i,0}=x^{i,i}_{i,i},\quad\sum^{n}_{i=0}\gamma^{0,0}_{i,0}x^{0,0}_{i,0}=2^{n},$
$x^{t,p}_{i,j}=0\quad\quad\text{if}\quad t-p\quad\text{is odd}\quad\text{or}\quad 0<i+j-t-p<\delta\,,$
(158) $\bigoplus_{\begin{subarray}{c}a,k\in\mathds{N}_{0}\\ 0\leq a\leq k\leq n+a-k\end{subarray}}\left(\sum\limits_{\begin{subarray}{c}t,p\in\mathds{N}_{0}\\ 0\leq p\leq t\leq i,j\\ i+j\leq t+n\end{subarray}}\alpha(i,j,t,p,a,k)x_{i,j}^{t,p}\right)_{i,j=k}^{n+a-k}\succeq 0\,.$
where $\alpha(i,j,t,p,a,k)$ is given by Eq. (6.3). Note that the condition $x^{t,p}_{i,j}=0$ if $0<i+j-t-p<\delta$ is equivalent to the pure code condition from Eq. (141). This can be seen from that fact that when $j=0$, it simplifies to $x^{0,0}_{i,0}=x^{i,i}_{i,i}=0$ if $0<i<\delta$. But $x_{i,i}^{i,i}=0$ implies that $x_{i,j}^{t,p}=0$ for any $j,t,p$: Consider $x\in{\boldsymbol{4}}^{n}$ with $s(x)=i$. Then $\tilde{\Gamma}_{xx}=x^{i,i}_{i,i}=0$. The constraint $\tilde{\Gamma}\succeq 0$ implies that $\tilde{\Gamma}_{xy}=x^{t,p}_{i,j}=0$ for any $y$ with $|s(y)|=j$, $\lvert s(x)\cap s(y)\rvert=t$ and |$\lvert\{m\,\mid\,x_{m}=y_{m}\neq 0\}\rvert=p$. This is due to the fact that the $2\times 2$ minor of $\tilde{\Gamma}$ involving $x$ and $y$ must be positive-semidefinite, and thus if $\tilde{\Gamma}_{xx}=0$ then also $\tilde{\Gamma}_{xy}=0$.
Consider now the Lagrangian of SDP (A.1). Associate with each equality constraint a dual variable ${u},{v_{i}},{w},r^{t,p}_{i,j}\in\mathds{R}$ and with each semidefinite constraint a dual semidefinite variable $Y^{(a,k)}\succeq 0$. Eq. (A.1) is a feasibility problem and so the objective function equals zero. Then
$\mathcal{L}$ $=\left(1-x^{0,0}_{0,0}\right){u}+\sum^{n}_{i=1}\left(x^{0,0}_{i,0}-x^{i,i}_{i,i}\right){v_{i}}$
$\quad+\left(2^{n}-\sum^{n}_{i=0}\gamma^{0,0}_{i,0}x^{0,0}_{i,0}\right){w}\quad+\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ (t-p)\,\text{odd}\text{ or}\\ 0<i+j-t-p<\delta\end{subarray}}x^{t,p}_{i,j}r^{{t,p}}_{i,j}$
(159) $\quad-\sum_{(i,j,t,p)\in\mathcal{I}(4,n)}\left(\sum^{\min(i,j)}_{k=0}\sum^{k}_{\begin{subarray}{c}a=\max(i,j)\\ +k-n\end{subarray}}\!\!\!\!\!\!\alpha(i,j,t,p,a,k)Y^{(a,k)}_{i-k,j-k}\right)x^{t,p}_{i,j}\,.$
For the last term, we have developed the trace of each block of the direct sum in Eq. (A.1) multiplied by the dual variable $Y^{(a,k)}$ to obtain
(160) $\sum_{\begin{subarray}{c}a,k\in\mathds{N}_{0}\\ 0\leq a\leq k\leq n+a-k\end{subarray}}\left(\sum^{n+a-k}_{i,j=k}\sum\limits_{\begin{subarray}{c}t,p\in\mathds{N}_{0}\\ 0\leq p\leq t\leq i,j\\ i+j\leq t+n\end{subarray}}\alpha(i,j,t,p,a,k)x_{i,j}^{t,p}Y^{(a,k)}_{i-k,j-k}\right)\,,$
where we have used the fact that $Y^{(a,k)}$ is symmetric. Define the following set corresponding to the sums appearing in Eq. (160),
=$\mathcal{R}=\big{\{}(i,j,t,p,a,k)\,\big{|}\,\,$ $k\leq i,j\leq n+a-k\,,\quad 0\leq a\leq k\leq n+a-k\,,$
(161) $i+j\leq t+n\quad\text{and}\quad 0\leq p\leq t\leq i,j\big{\}}\,.$
Eq. (160) then simplifies to $\sum_{(i,j,t,p,a,k)\in\mathcal{R}}\alpha(i,j,t,p,a,k)x^{t,p}_{i,j}Y^{(a,k)}_{i-k,j-k}$. Note that $\mathcal{R}$ is equivalently defined by Eq. (160) is then simplify to $\sum_{(i,j,t,p,a,k)\in\mathcal{R}}\alpha(i,j,t,p,a,k)x^{t,p}_{i,j}Y^{(a,k)}_{i-k,j-k}$. Note that the set $\mathcal{R}$ is equivalently defined by
=$\mathcal{R}=\big{\{}(i,j,t,p,a,k)\,\big{|}\,\,$ $(i,j,t,p)\in\mathcal{I}(4,n)\,,\quad 0\leq k\leq\min(i,j)\,,$
(162) $\max(i,j)\leq n+a-k\,,\quad\text{and}\quad a\leq k\big{\}}\,,$
with $\mathcal{I}(4,n)$ given by Eq. (76). To see this note that the first two inequalities in Eq. (A.1) combine to $0\leq a\leq k\leq i,j\leq n+a-k$ from which $i,j\leq n$ follows. Then the remaining condition $0\leq p\leq t\leq i,j$ is captured by $\mathcal{I}(4,n)$. Therefore, $\sum_{(i,j,t,p,a,k)\in\mathcal{R}}\alpha(i,j,t,p,a,k)x^{t,p}_{i,j}Y^{(a,k)}_{i-k,j-k}$ can also be written as
(163) $\sum_{(i,j,t,p)\in\mathcal{I}(4,n)}\left(\sum^{\min(i,j)}_{k=0}\sum^{k}_{\begin{subarray}{c}a=\max(i,j)+k-n\end{subarray}}\!\!\!\!\!\!\alpha(i,j,t,p,a,k)Y^{(a,k)}_{i-k,j-k}\right)x^{t,p}_{i,j}\,.$
Define now $y^{t,p}_{i,j}\ =\sum^{\min(i,j)}_{k=0}\sum^{k}_{a=\max(i,j)+k-n}\alpha(i,j,t,p,a,k)Y^{(a,k)}_{i-k,j-k}$ which is the inner term of the last expression in Eq. (A.1). Since $Y^{(a,k)}_{i-k,j-k}$ and $\alpha(i,j,t,p,a,k)$ are symmetric functions with respect to $i$ and $j$, then also $y^{t,p}_{i,j}$ is symmetric. Factorize the primal variables in the Lagrangian of Eq. (A.1),
$\mathcal{L}$ $=({u}+2^{n}{w})+x^{0,0}_{0,0}\Big{(}-{u}-{w}-y^{0,0}_{0,0}\Big{)}+\sum^{\delta-1}_{i=1}x^{0,0}_{i,0}\Big{(}{v_{i}}-\gamma^{0,0}_{i,0}{w}+r^{0,0}_{i,0}+r^{0,0}_{0,i}-2y^{0,0}_{i,0}\Big{)}$
$\quad+\sum^{n}_{i=\delta}x^{0,0}_{i,0}\Big{(}{v_{i}}-\gamma^{0,0}_{i,0}{w}-2y^{0,0}_{i,0}\Big{)}+\sum^{n}_{i=1}x^{i,i}_{i,i}\Big{(}-{v_{i}}-y^{i,i}_{i,i}\Big{)}$
(164) $\quad+\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i,j\neq 0\\ (t-p)\,\text{odd}\text{ or}\\ 0<i+j-t-p<\delta\end{subarray}}x^{t,p}_{i,j}\Big{(}r^{t,p}_{i,j}-y^{t,p}_{i,j}\Big{)}-\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i,j\neq 0\\ (t-p)\text{ even and}\\ i+j-t-p\geq\delta\end{subarray}}x^{t,p}_{i,j}y^{t,p}_{i,j}\,.$
Here we used that $\gamma^{0,0}_{0,0}=1$, $x^{00}_{i0}=x^{00}_{0i}$, and $y^{00}_{i0}=y^{00}_{0i}$, to decompose
(165) $\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ (t-p)\,\text{odd}\text{ or}\\ 0<i+j-t-p<\delta\end{subarray}}x^{t,p}_{i,j}r^{{t,p}}_{i,j}$ $=\sum^{\delta-1}_{i=1}x^{0,0}_{i,0}(r^{{0,0}}_{i,0}+r^{{0,0}}_{0,i})+\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i,j\neq 0\\ (t-p)\,\text{odd}\text{ or}\\ 0<i+j-t-p<\delta\end{subarray}}x^{t,p}_{i,j}r^{{t,p}}_{i,j}\,,$
(166) $\sum_{(i,j,t,p)\in\mathcal{I}(4,n)}x^{t,p}_{i,j}y^{t,p}_{i,j}$ $=x^{0,0}_{0,0}y^{0,0}_{0,0}+\sum^{n}_{i=1}2x^{0,0}_{i,0}y^{0,0}_{i,0}+\!\!\!\!\!\!\!\!\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i,j\neq 0\\ (t-p)\,\text{odd}\text{ or}\\ 0<i+j-t-p<\delta\end{subarray}}x^{t,p}_{i,j}y^{{t,p}}_{i,j}+\!\!\!\!\!\!\!\!\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i,j\neq 0,\,\\ (t-p)\text{ even and}\\ i+j-t-p\geq\delta,\\ \end{subarray}}x^{t,p}_{i,j}y^{t,p}_{i,j}\,.$
The dual objective function is then
(167) $\inf_{x^{t,p}_{i,j}}\mathcal{L}$ $=({u}+2^{n}w)\quad\text{if}\quad\begin{cases}(1)\quad 0=-{u}-{w}-y^{0,0}_{0,0}\,,\\ (2)\quad 0={v_{i}}-\gamma^{0,0}_{i,0}{w}+r^{0,0}_{i,0}+r^{0,0}_{0,i}-2y^{0,0}_{i,0}&\text{if}\quad 0<i<\delta\,,\\ (3)\quad 0={v_{i}}-\gamma^{0,0}_{i,0}{w}-2y^{0,0}_{i,0}&\text{if}\quad\delta\leq i\leq n\,,\\ (4)\quad 0=-{v_{i}}-y^{i,i}_{i,i}&\text{if}\quad 0<i\leq n\,,\\ (5)\quad 0=r^{t,p}_{i,j}-y^{t,p}_{i,j}&\text{if}\quad(t-p)\,\,\text{odd}\quad\text{or}\\ \quad&\quad 0<i+j-t-p<\delta\\ &\quad\quad\text{with}\quad i,j\neq 0\,,\\ (6)\quad 0=y^{t,p}_{i,j}&\text{if}\quad i,j\neq 0,\quad(t-p)\,\,\text{even }\\ &\quad\text{and}\quad\delta\leq i+j-t-p\,.\end{cases}$
Otherwise, it tends to $\infty$- ∞. Note that conditions (2) and (5) appearing in Eq. (167),
${v_{i}}-\gamma^{0,0}_{i,0}{w}+r^{0,0}_{i,0}+r^{0,0}_{0,i}-2y^{0,0}_{i,0}=0\quad$ $\text{if}\quad 0<i<\delta\,,$
$r^{t,p}_{i,j}-y^{t,p}_{i,j}=0\quad$ $\text{if}\quad(t-p)\,\,\text{odd}\quad$
(168) $\quad\quad\text{or}\quad 0<i+j-t-p<\delta\quad\text{with}\quad i,j\neq 0\,.$
do not constrain the objective function. This can be seen from the fact that the variables $r^{t,p}_{i,j}$ do not appear in the objective function and neither in the rest of constraints. Thus, they act as slack variables which can always take values so that the above two conditions are satisfied.
All dual variables can be expressed in terms of the semidefinite dual constraints $Y^{(a,k)}$ and $w$ using the remaining constraints in Eq. (167): use condition (1) to eliminate $u$ and condition (4) to eliminate $v_{i}$.
The dual problem can then be expressed as Eq. (25). In particular, the objective function reads $u+2^{n}w=(2^{n}-1)w-y_{0,0}^{0,0}$, and the remaining constraints (3) and (6) read:
00 0 $=y_{i,i}^{i,i}+\gamma_{i,0}^{0,0}w+2y_{i,0}^{0,0}\quad$ if $\delta\leq i\leq n$
(169) 00 0 $=y^{t,p}_{i,j}$ if $i,j\neq 0,\quad(t-p)\,\,\text{even}\,,\quad i+j-t-p\geq\delta\,.$
The dual program then reads
$\alpha\quad=\quad\max_{Y^{(a,k)}\,,\,{w}}\quad$ $(2^{n}-1)w-y_{0,0}^{0,0}$
subject to $Y^{(a,k)}\succeq 0\,,$
$y^{t,p}_{i,j}=\sum^{\min(i,j)}_{k=0}\sum^{k}_{a=\max(i,j)+k-n}\alpha(i,j,t,p,a,k)Y^{(a,k)}_{i-k,j-k}\,,$
$y_{i,i}^{i,i}+\gamma_{i,0}^{0,0}w+2y_{i,0}^{0,0}=0\quad\quad\text{if}\quad\delta\leq i\leq n\,,$
$y^{t,p}_{i,j}=0\hskip 85.00012pt\text{if}\quad i,j\neq 0\,,\quad t-p\quad\text{is even }\quad$
(170) $\hskip 116.00023pt\quad\text{and}\quad i+j-t-p\geq\delta\,.$
Finally, one can check from Eq. (77) that $\gamma^{0,0}_{i,0}=\gamma^{i,i}_{i,i}$. Thus, by mapping the variable $y^{t,p}_{ij}\mapsto\gamma^{t,p}_{i,j}y^{t,p}_{ij}$, the constraints in the SDP above simplify to Eq. (25) since $\gamma^{t,p}_{ij}$ factorizes. This ends the proof. ∎
A.2. Dual of the symmetry-reduced SDP
Proposition 26.
Given the relaxation of symmetry reduced SDP in Eq. (9.2), the dual program reads
$\alpha\quad=\quad\max_{Y^{(a,k)},Q_{i},C_{i}}\quad$ $(KD_{0}-C_{0})+\left(\frac{2^{n}}{K}-1\right)Q_{0}-{y^{0,0}_{0,0}}\quad$
subject to $Y^{(a,k)}\succeq 0,$
$y^{t,p}_{i,j}\ =\frac{1}{\gamma^{t,p}_{i,j}}\sum^{\min(i,j)}_{k=0}\sum^{k}_{a=\max(i,j)+k-n}\alpha(i,j,t,p,a,k)Y^{(a,k)}_{i-k,j-k}\,,$
$D_{i}=2^{-n}\sum^{\delta-1}_{j=0}K_{j}(i,n)C_{j}\,,$
$(KD_{i}-C_{i})+\left(\frac{2^{n}}{K}-2\right)Q_{i}-2y^{0,0}_{i,0}-Q_{0}-y^{i,i}_{i,i}=0\quad\text{if}\quad 0<i<\delta\,,$
$KD_{i}+\left(\frac{2^{n}}{K}-2\right)Q_{i}-2y^{0,0}_{i,0}-Q_{0}-y^{i,i}_{i,i}=0\quad\quad\quad\quad\text{if}\quad\delta\leq i\leq n\,,$
$\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},t^{\prime},p^{\prime})\in\mathcal{I}(4,n)\\ (t-p)=(t^{\prime}-p^{\prime})\\ (i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,\text{a}\\ \,\text{permutation of}\,(i,j,i+j-t-p)\\ \end{subarray}}\!\!\!\!\!\!\!\!\!\!\left(y^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}+Q_{i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime}}\right)=0\quad\hskip 33.00008pt\text{if}\quad(t-p)\quad\text{even}$
(171) $\hskip 170.71652pt\quad\text{and}\quad i,j,i+j-t-p\neq 0\,.$
Here the dual variables are real matrices $Y^{(a,k)}$ of size $n+a-2k$ with $0\leq a\leq k\leq n+a-k$ and $Q_{i},C_{i}\in\mathds{R}$ with $0\leq i\leq n$.
Proof.
For convenience, we perform the following variations to the SDP (9.2):
  1. (1)
    Substitute $\sum^{n}_{i=0}\gamma^{0,0}_{i,0}x^{0,0}_{i,0}=\frac{2^{n}}{K}$ for $\frac{K}{2^{n}}\sum^{n}_{i=0}K_{0}(i;n)\gamma^{0,0}_{i,0}x^{0,0}_{i,0}=x^{0,0}_{0,0}\gamma^{0,0}_{0,0}$ since they are equivalent (see Proposition 19 ff.).
  2. (2)
    Consider $x^{0,0}_{i,0}=x^{i,i}_{i,i}$ independent from the remaining conditions in (iii) from Proposition 18.
As result the SDP from Eq. (9.2) is now written as
find $x^{t,p}_{i,j}$
subject to $x^{0,0}_{0,0}=1,\quad x^{0,0}_{i,0}=x^{i,i}_{i,i}\,,$
$x^{t,p}_{i,j}=0\quad\quad\quad\text{if}\quad t-p\quad\text{is odd}\,,$
$x^{t,p}_{i,j}=x^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}\,,\quad\text{if}\quad t-p=t^{\prime}-p^{\prime}\quad\text{is even}\quad\text{and}\quad(i,j,i+j-t-p)\quad\text{a}$
permutation $\hskip 73.97733pt\text{permutation of}\,\,(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\quad\text{with}\quad i,j,i+j-t-p\neq 0\,,$
$\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ k=i+j-t-p\end{subarray}}\gamma^{t,p}_{i,j}x^{t,p}_{i,j}=\frac{2^{n}}{K}\gamma^{0,0}_{k,0}x^{0,0}_{k,0}\,,$
$K2^{-n}\sum^{n}_{i=0}K_{j}(i,n)\gamma^{0,0}_{i,0}x^{0,0}_{i,0}=\gamma^{0,0}_{j,0}x^{0,0}_{j,0}\quad\text{if}\quad 0\leq j<\delta\,,$
(172) $\bigoplus_{\begin{subarray}{c}a,k\in\mathds{N}_{0}\\ 0\leq a\leq k\leq n+a-k\end{subarray}}\left(\sum\limits_{\begin{subarray}{c}t,p\in\mathds{N}_{0}\\ 0\leq p\leq t\leq i,j\\ i+j\leq t+n\end{subarray}}\alpha(i,j,t,p,a,k)x_{i,j}^{t,p}\right)_{i,j=k}^{n+a-k}\succeq 0\,,$
where $\alpha(i,j,t,p,a,k)$ is given by Eq. (6.3). In all the development, we will impose that $\gamma^{0,0}_{0,0}=1$.
Similarly to the previous section, we now construct the Lagrangian. Associate with each equality constraint a dual variable ${u},{v_{i}},r^{t,p}_{i,j},s^{t,t^{\prime},p,p^{\prime}}_{i,i^{\prime},j,j^{\prime}},Q_{i},C_{i}\in\mathds{R}$ and with each semidefinite constraint a dual semidefinite variable $Y^{(a,k)}\succeq 0$ and the objective function equals to 00. Then
$\mathcal{L}$ $=\left(1-x^{0,0}_{0,0}\right){u}+\sum^{n}_{i=1}\left(x^{0,0}_{i,0}-x^{i,i}_{i,i}\right){v_{i}}$
$\quad+\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ (t-p)\,\text{odd}\end{subarray}}x^{t,p}_{i,j}r^{{t,p}}_{i,j}\quad+\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i,j,i+j-t-p\neq 0\\ (t-p)\,\text{even}\end{subarray}}\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},t^{\prime},p^{\prime})\in\mathcal{I}(4,n)\\ (t-p)=(t^{\prime}-p^{\prime})\\ (i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,\text{a}\\ \,\text{permutation of}\,(i,j,i+j-t-p)\\ \end{subarray}}\!\!\!\!\!\!\!\!\!\!\!\!(x^{t,p}_{i,j}-x^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}})s^{{t,t^{\prime},p,p^{\prime}}}_{i,i^{\prime},j,j^{\prime}}$
$\quad+\sum^{n}_{k=0}\frac{2^{n}}{K}\gamma^{0,0}_{k,0}x^{0,0}_{k,0}Q_{k}\quad-\quad\sum^{n}_{k=0}\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ k=i+j-t-p\end{subarray}}\gamma^{t,p}_{i,j}x^{t,p}_{i,j}Q_{i+j-t-p}$
$\quad+\sum^{n}_{i=0}K\left(2^{-n}\sum^{\delta-1}_{j=0}K_{j}(i,n)C_{j}\right)\gamma^{0,0}_{i,0}x^{0,0}_{i,0}-\sum^{\delta-1}_{j=0}\gamma^{0,0}_{j,0}x^{0,0}_{j,0}C_{j}$
(173) $\quad-\sum_{(i,j,t,p)\in\mathcal{I}(4,n)}\left(\sum^{\min(i,j)}_{k=0}\sum^{k}_{\begin{subarray}{c}a=\max(i,j)\\ +k-n\end{subarray}}\!\!\!\!\!\!\alpha(i,j,t,p,a,k)Y^{(a,k)}_{i-k,j-k}\right)x^{t,p}_{i,j}\,.$
As shown in the development from Eq. (160) to Eq. (163), the last term arises as the the trace of the product of the primal and dual semidefinite variables. Before proceeding with the Lagrangian, we simplify the last term in the second line of Eq. (A.2) coming from the structure constraint by reorganizing its elements as
$\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i,j,i+j-t-p\neq 0\\ (t-p)\,\text{even}\end{subarray}}\!\!\!\!\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},t^{\prime},p^{\prime})\in\mathcal{I}(4,n)\\ (t-p)=(t^{\prime}-p^{\prime})\\ (i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,\text{a}\\ \,\text{permutation of}\,(i,j,i+j-t-p)\\ \end{subarray}}\!\!\!\!\!\!\!\!\!\!(x^{t,p}_{i,j}-x^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}})s^{{t,t^{\prime},p,p^{\prime}}}_{i,i^{\prime},j,j^{\prime}}$
$=\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i,j,i+j-t-p\neq 0\\ (t-p)\,\text{even}\end{subarray}}x^{t,p}_{i,j}\Bigg{(}\!\!\!\!\!\!\!\!\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},t^{\prime},p^{\prime})\in\mathcal{I}(4,n)\\ (t-p)=(t^{\prime}-p^{\prime})\\ (i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,\text{a}\\ \,\text{permutation of}\,(i,j,i+j-t-p)\\ \end{subarray}}\!\!\!\!\!\!\!\!\!\!(s^{{t,t^{\prime},p,p^{\prime}}}_{i,i^{\prime},j,j^{\prime}}-s^{{t^{\prime},t,p^{\prime},p}}_{i^{\prime},i,j^{\prime},j})\Bigg{)}$
(174) $=\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i,j,i+j-t-p\neq 0\\ (t-p)\,\text{even}\end{subarray}}x^{{t,p}}_{i,j}f^{{t,p}}_{i,j}\,,$
where
(175) $f^{{t,p}}_{i,j}=\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},t^{\prime},p^{\prime})\in\mathcal{I}(4,n)\\ (t-p)=(t^{\prime}-p^{\prime})\\ (i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,\text{a}\\ \,\text{permutation of}\,(i,j,i+j-t-p)\\ \end{subarray}}\!\!\!\!\!\!\!\!\!\!(s^{{t,t^{\prime},p,p^{\prime}}}_{i,i^{\prime},j,j^{\prime}}-s^{{t^{\prime},t,p^{\prime},p}}_{i^{\prime},i,j^{\prime},j})\,.$
From the definition of $f^{t,p}_{i,j}$ follows a first constraint: for $(a,b,c,d)\in\mathcal{I}(4,n)$ with $(c-d)$ even and $a,b,a+b-c-d~{}\neq~{}0$,
(176) $\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ (c-d)=(t-p)\\ (i,j,i+j-t-p)\,\text{a}\\ \,\text{permutation of}\,(a,b,a+b-c-d)\\ \end{subarray}}\!\!\!\!\!\!\!\!\!\!f^{t,p}_{i,j}=0\,.$
This can be seen by taking Eq. (175) and developing the left hand side of Eq. (176) as
(177) $\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ (c-d)=(t-p)\\ (i,j,i+j-t-p)\,\text{a}\\ \,\text{permutation of}\,(a,b,a+b-c-d)\\ \end{subarray}}\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},t^{\prime},p^{\prime})\in\mathcal{I}(4,n)\\ (t-p)=(t^{\prime}-p^{\prime})\\ (i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,\text{a}\\ \,\text{permutation of}\,(i,j,i+j-t-p)\\ \end{subarray}}\!\!\!\!\!\!\!\!\!\!s^{{t,t^{\prime},p,p^{\prime}}}_{i,i^{\prime},j,j^{\prime}}-\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ (c-d)=(t-p)\\ (i,j,i+j-t-p)\,\text{a}\\ \,\text{permutation of}\,(a,b,a+b-c-d)\\ \end{subarray}}\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},t^{\prime},p^{\prime})\in\mathcal{I}(4,n)\\ (t-p)=(t^{\prime}-p^{\prime})\\ (i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,\text{a}\\ \,\text{permutation of}\,(i,j,i+j-t-p)\\ \end{subarray}}\!\!\!\!\!\!\!\!\!\!s^{{t^{\prime},t,p^{\prime},p}}_{i^{\prime},i,j^{\prime},j}$
Note that the right and left hand side of the above equation sum over the same elements and thus, they are equal.
Define also $y^{t,p}_{i,j}\ =\sum^{\min(i,j)}_{k=0}\sum^{k}_{a=\max(i,j)+k-n}\alpha(i,j,t,p,a,k)Y^{(a,k)}_{i-k,j-k}$ and $D_{i}=2^{-n}\sum^{\delta-1}_{j=0}K_{j}(i,n)C_{j}$. We then factorize the primal variables so that
$\mathcal{L}$ $={u}+x^{0,0}_{0,0}\left(-{u}+(KD_{0}-C_{0})+\left(\frac{2^{n}}{K}-1\right)Q_{0}-y^{0,0}_{0,0}\right)$
$\quad+\sum^{\delta-1}_{i=1}x^{0,0}_{i,0}\Big{(}{v_{i}}+\gamma^{0,0}_{i,0}(KD_{i}-C_{i})+\frac{2^{n}}{K}\gamma^{0,0}_{i,0}Q_{i}-2\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}\Big{)}$
$\quad+\sum^{n}_{i=\delta}x^{0,0}_{i,0}\Big{(}{v_{i}}+\gamma^{0,0}_{i,0}KD_{i}+\frac{2^{n}}{K}\gamma^{0,0}_{i,0}Q_{i}-2\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}\Big{)}$
$\quad+\sum^{n}_{i=1}x^{i,i}_{i,i}\Big{(}-{v_{i}}-\gamma^{i,i}_{i,i}Q_{0}-y^{i,i}_{i,i}\Big{)}$
$\quad+\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ (t-p)\,\text{odd}\end{subarray}}x^{t,p}_{i,j}\Big{(}r^{t,p}_{i,j}-\gamma^{t,p}_{i,j}Q_{i+j-t-p}-y^{t,p}_{i,j}\Big{)}$
(178) $\quad+\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i,j,i+j-t-p\neq 0\\ (t-p)\,\text{even}\end{subarray}}x^{t,p}_{i,j}(f^{t,p}_{i,j}-\gamma^{t,p}_{i,j}Q_{i+j-t-p}-y^{t,p}_{i,j})\,.$
As done in Eq. (165), we have decomposed
(179) $\sum_{(i,j,t,p)\in\mathcal{I}(4,n)}x^{t,p}_{i,j}y^{t,p}_{i,j}$ $=x^{0,0}_{0,0}y^{0,0}_{0,0}+\sum^{n}_{i=1}2x^{0,0}_{i,0}y^{0,0}_{i,0}+\!\!\!\!\!\!\!\!\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ (t-p)\,\text{odd}\end{subarray}}x^{t,p}_{i,j}r^{{t,p}}_{i,j}+\!\!\!\!\!\!\!\!\sum_{\begin{subarray}{c}(i,j,t,p)\in\mathcal{I}(4,n)\\ i,j,i+j-t-p\neq 0\\ (t-p)\,\text{even}\end{subarray}}x^{t,p}_{i,j}y^{t,p}_{i,j}\,.$
By Eq. (A.2), the dual objective function is then
(180) $\inf_{x^{t,p}_{i,j}}\mathcal{L}$ $={u}\quad\text{if}\quad\begin{cases}(1)\quad 0=-{u}+(KD_{0}-C_{0})+\left(\frac{2^{n}}{K}-1\right)Q_{0}-{y^{0,0}_{0,0}}\,,\\ (2)\quad 0={v_{i}}+\gamma^{0,0}_{i,0}(KD_{i}-C_{i})+\frac{2^{n}}{K}\gamma^{0,0}_{i,0}Q_{i}-2\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}&\text{if}\quad 0<i<\delta\,,\\ (3)\quad 0={v_{i}}+\gamma^{0,0}_{i,0}KD_{i}+\frac{2^{n}}{K}\gamma^{0,0}_{i,0}Q_{i}-2\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}&\text{if}\quad\delta\leq i\leq n\,,\\ (4)\quad 0=-{v_{i}}-\gamma^{i,i}_{i,i}Q_{0}-y^{i,i}_{i,i}\quad&\text{if}\quad 0<i\leq n\,,\\ (5)\quad 0=r^{t,p}_{i,j}-\gamma^{t,p}_{i,j}Q_{i+j-t-p}-y^{t,p}_{i,j}&\text{if}\quad(t-p)\,\,\text{odd}\,,\\ (6)\quad 0=f^{t,p}_{i,j}-\gamma^{t,p}_{i,j}Q_{i+j-t-p}-y^{t,p}_{i,j}&\text{if}\quad(t-p)\,\,\text{even}\,,\\ &\quad i,j\neq 0\quad\text{and}\\ &\quad\quad i+j-t-p\neq 0\,.\end{cases}$
Otherwise, it tends to $\infty$- ∞. Similarly to Eq. (A.1), condition (5) above,
(181) $r^{t,p}_{i,j}-\gamma^{t,p}_{i,j}Q_{i+j-t-p}-y^{t,p}_{i,j}=0\quad\text{if}\quad(t-p)\,\,\text{odd}$
does not constrain the objective function. Again, $r^{t,p}_{i,j}$ acts as a slack variable which can always take values such that Eq. (181) is satisfied.
All dual variables can be expressed in terms of $Y^{(a,k)},Q_{i},C_{i}$ using the remaining of constraints in Eq. (180). First, use condition (1) to derive the objective function of the dual which reads ${u}=(KD_{0}-C_{0})+\left(\frac{2^{n}}{K}-1\right)Q_{0}-{y^{0,0}_{0,0}}$. Then apply condition (4) to (2) and (3) to obtain respectively,
(182) 00$\quad$ 0 $=\gamma^{0,0}_{i,0}(KD_{i}-C_{i})+\left(\frac{2^{n}}{K}-2\right)\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}-\gamma^{i,i}_{i,i}Q_{0}-y^{i,i}_{i,i}\quad\text{if}\quad 0<i<\delta\,,$
(183) 00$\quad$ 0 $=\gamma^{0,0}_{i,0}KD_{i}+\left(\frac{2^{n}}{K}-2\right)\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}-\gamma^{i,i}_{i,i}Q_{0}-y^{i,i}_{i,i}\quad\quad\quad\quad\text{if}\quad\delta\leq i\leq n\,.$
Finally, we use the property of $f^{t,p}_{i,j}$ described in Eq. (176) together with condition (6) and thus,
$\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},t^{\prime},p^{\prime})\in\mathcal{I}(4,n)\\ (t-p)=(t^{\prime}-p^{\prime})\\ (i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,\text{a}\\ \,\text{permutation of}\,(i,j,i+j-t-p)\\ \end{subarray}}\!\!\!\!\!\!\!\!\!\!\left(y^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}+\gamma^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}Q_{i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime}}\right)=0\quad\hskip 3.00003pt\text{if}\quad(t-p)\quad\text{even}$
(184) $\quad\text{and}\quad i,j,i+j-t-p\neq 0.$
The dual program then reads as
$\alpha\quad=\quad\max_{Y^{(a,k)},Q_{i},C_{i}}\quad$ $(KD_{0}-C_{0})+\left(\frac{2^{n}}{K}-1\right)Q_{0}-{y^{0,0}_{0,0}}\quad$
subject to $Y^{(a,k)}\succeq 0,$
$y^{t,p}_{i,j}\ =\sum^{\min(i,j)}_{k=0}\sum^{k}_{a=\max(i,j)+k-n}\alpha(i,j,t,p,a,k)Y^{(a,k)}_{i-k,j-k}\,,$
$D_{i}=2^{-n}\sum^{\delta-1}_{j=0}K_{j}(i,n)C_{j}\,,$
$\gamma^{0,0}_{i,0}(KD_{i}-C_{i})+\left(\frac{2^{n}}{K}-2\right)\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}-\gamma^{i,i}_{i,i}Q_{0}-y^{i,i}_{i,i}=0\quad\text{if}\quad 0<i<\delta\,,$
$\gamma^{0,0}_{i,0}KD_{i}+\left(\frac{2^{n}}{K}-2\right)\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}-\gamma^{i,i}_{i,i}Q_{0}-y^{i,i}_{i,i}=0\quad\quad\quad\quad\text{if}\quad\delta\leq i\leq n\,,$
$\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},t^{\prime},p^{\prime})\in\mathcal{I}(4,n)\\ (t-p)=(t^{\prime}-p^{\prime})\\ (i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\,\text{a}\\ \,\text{permutation of}\,(i,j,i+j-t-p)\\ \end{subarray}}\!\!\!\!\!\!\!\!\!\!\left(y^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}+\gamma^{t^{\prime},p^{\prime}}_{i^{\prime},j^{\prime}}Q_{i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime}}\right)=0\quad\hskip 56.00014pt\text{if}\quad(t-p)$
(185) $\hskip 170.71652pt\text{even}\quad\text{and}\quad i,j,i+j-t-p\neq 0.$
Finally, one can check from Eq. (77) that $\gamma^{t,p}_{ij}=\gamma^{t^{\prime},p^{\prime}}_{i^{\prime}j^{\prime}}$ if $t-p=t^{\prime}-p^{\prime}$ is even and $(i,j,i+j-t-p)$ is a permutation of $(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})$. In particular, $\gamma^{0,0}_{i,0}=\gamma^{i,i}_{i,i}$. Thus, by mapping the variable $y^{t,p}_{ij}\mapsto\gamma^{t,p}_{i,j}y^{t,p}_{ij}$, the constraints in the SDP simplify to Eq. (26) since $\gamma^{t,p}_{ij}$ factorizes. This ends the proof. ∎
Appendix B Infeasibility certificate $(\!(7,1,4)\!)_{2}$ code
The following matrices $Y^{(a,k)}$ are a solution of the dual program (9.1) for $n=7$ and $\delta=4$ with $\alpha=0.58>0$ and a maximum constraint violation of $\mathcal{O}(10^{-16})$. Note that the first constraint from Eq. (25) allows to determine the value from $w$ using $Y^{(a,k)}$ since $w=-y^{i,i}_{i,i}-y^{0,0}_{i,0}$ if $\delta\leq i\leq n$. In particular, one can choose $w=-y^{n,n}_{n,n}-y^{0,0}_{n,0}$. The martices thus provide an infeasibility certificate and show that a $(\!(7,1,4)\!)_{2}$ quantum code does not exist.
$Y^{(0,0)}$ $=\begin{bmatrix}124&0&0&0&-9&-15&-26&-43\\ 0&115&0&0&0&0&0&0\\ 0&0&22&0&0&0&0&0\\ 0&0&0&3&0&0&0&0\\ -9&0&0&0&1&1&2&1\\ -15&0&0&0&1&3&2&2\\ -26&0&0&0&2&2&7&11\\ -43&0&0&0&1&2&11&43\\ \end{bmatrix}$
$Y^{(0,1)}$ $=\begin{bmatrix}120&0&0&0&0&0\\ 0&238&0&0&0&0\\ 0&0&66&0&0&0\\ 0&0&0&9&10&-10\\ 0&0&0&10&15&3\\ 0&0&0&-10&3&92\\ \end{bmatrix}$
$Y^{(0,2)}$ $=\begin{bmatrix}140&0&0&0\\ 0&274&0&0\\ 0&0&91&24\\ 0&0&24&330\\ \end{bmatrix}$
$Y^{(0,3)}$ $=\begin{bmatrix}84&0\\ 0&12\\ \end{bmatrix}$
$Y^{(1,1)}$ $=\begin{bmatrix}120&0&0&0&0&0&0\\ 0&120&0&0&0&0&0\\ 0&0&63&0&0&0&0\\ 0&0&0&11&4&-5&5\\ 0&0&0&4&7&4&-8\\ 0&0&0&-5&4&29&56\\ 0&0&0&5&-8&56&269\\ \end{bmatrix}$
$Y^{(1,2)}$ $=\begin{bmatrix}117&0&0&0&0\\ 0&291&0&0&0\\ 0&0&130&56&285\\ 0&0&56&419&325\\ 0&0&285&325&831\\ \end{bmatrix}$
$Y^{(1,3)}$ $=\begin{bmatrix}126&0&0\\ 0&\frac{405}{4}&207\\ 0&207&470\\ \end{bmatrix}$
$Y^{(1,4)}$ $=\begin{bmatrix}\frac{93}{2}\\ \end{bmatrix}$
$Y^{(2,2)}$ $=\begin{bmatrix}114&0&0&0&0&0\\ 0&109&0&0&0&0\\ 0&0&53&26&85&20\\ 0&0&26&108&-69&123\\ 0&0&85&-69&352&147\\ 0&0&20&123&147&937\\ \end{bmatrix}$
$Y^{(2,3)}$ $=\begin{bmatrix}140&0&0&0\\ 0&127&65&\frac{125}{3}\\ 0&65&57&46\\ 0&\frac{125}{3}&46&81\\ \end{bmatrix}$
$Y^{(2,4)}$ $=\begin{bmatrix}18&-29\\ -29&56\\ \end{bmatrix}$
$Y^{(3,3)}$ $=\begin{bmatrix}107&0&0&0&0\\ 0&24&-3&\frac{125}{4}&70\\ 0&-3&71&98&13\\ 0&\frac{125}{4}&98&224&149\\ 0&70&13&149&246\\ \end{bmatrix}$
$Y^{(3,4)}$ $=\begin{bmatrix}99&\frac{45}{2}&70\\ \frac{45}{2}&\frac{263}{4}&-60\\ 70&-60&158\\ \end{bmatrix}$
$Y^{(3,5)}$ $=\begin{bmatrix}\frac{39}{2}\\ \end{bmatrix}$
$Y^{(4,4)}$ $=\begin{bmatrix}106&10&\frac{-290}{3}&0\\ 10&\frac{16187991}{644740}&-44&38\\ \frac{-290}{3}&-44&\frac{6623}{36}&-82\\ 0&38&-82&88\\ \end{bmatrix}$
$Y^{(4,5)}$ $=\begin{bmatrix}\frac{25998073}{322370}&44\\ 44&\frac{379}{12}\\ \end{bmatrix}$
$Y^{(5,5)}$ $=\begin{bmatrix}\frac{21984965}{392883}&\frac{-78}{5}&\frac{-186}{5}\\ \frac{-78}{5}&\frac{1061}{80}&\frac{93}{5}\\ \frac{-186}{5}&\frac{93}{5}&\frac{34838427}{857647}\\ \end{bmatrix}$
$Y^{(5,6)}$ $=\begin{bmatrix}\frac{907}{8}\\ \end{bmatrix}$
$Y^{(6,6)}$ $=\begin{bmatrix}\frac{289}{5}&\frac{-548}{5}\\ \frac{-548}{5}&\frac{164534807}{670060}\\ \end{bmatrix}$
(186) $Y^{(7,7)}$ $=\begin{bmatrix}\frac{3813979}{522046}\\ \end{bmatrix}$